### v1

# Integral canonical models of exceptional Shimura varieties

Abstract
We prove that Shimura varieties admit integral canonical models for sufficiently large primes. In the case of abelian-type Shimura varieties, this recovers work of Kisin-Kottwitz for sufficiently large primes. We also prove the existence of integral canonical models for images of period maps corresponding to geometric families. We deduce several consequences from this, including an unramified rigid analogue of Borel’s extension theorem, a version of Tate semisimplicity, CM lifting theorems, and a version of Tate’s isogeny theorem for ordinary points.
1 Introduction
This paper is motivated by the following question, raised by Langlands and made precise by Milne: Given a Shimura variety $S_{K}(G,X)$ over its reflex field $E$, does $S_{K}(G,X)$ admit a “good model” $\mathscr{S}_{K}(G,X)$, over completions $\mathcal{O}_{E_{v}}$?
For the moduli space of principally polarized abelian varieties $\mathcal{A}_{g}$, for instance, the moduli problem provides a natural model. As not all Shimura varieties are known to have an associated geometric moduli problem (notably the exceptional Shimura varieties), it is preferable to have a model which is uniquely determined by certain intrinsic properties, possibly only for sufficiently large primes. At the very least, we would like our model to be smooth over $\mathcal{O}_{E_{v}}$, which already may not exist at certain primes. Indeed, even in the case of the modular curve, the level structure at $p$ needs to be maximal in order for there to exist a smooth model at $p$.
Smoothness alone however does not a priori guarantee a unique model. Milne defined an “extension property” (and this notion was refined by Moonen in [Moonen]) which formally guarantees uniqueness. Following Milne and Moonen, an integral model is said to be an integral canonical model if it is smooth and if it satisfies their extension property. The moduli space of principally polarized abelian varieties $\mathcal{A}_{g}$, for instance, is known to admit integral canonical models over $\mathbb{Z}_{p}$ whenever the level structure is prime to $p$, i.e. when the level structure is hyperspecial at $p$. Indeed, smoothness follows from the fact that the deformation theory of polarized abelian varieties with a prime-to-$p$-polarization is unobstructed (proved by Grothendieck), and Chai and Faltings [faltingschai] prove an extension result for abelian schemes over certain bases (see [Moonen, Section 3] for the precise results they prove) from which the Milne–Moonen extension property for $\mathcal{A}_{g}$ follows directly.
Langlands suggested more generally that integral canonical models should exist whenever the compact open $K_{p}$ at $p$ is hyperspecial. In the case of PEL Shimura varieties (i.e. Shimura varieties parameterizing polarized abelian varieties with extra endomorphisms), Kottwitz in [Ko] constructs integral models when the level structure at $p$ is hyperspecial. Kisin generalizes Kottwitz’ work to the setting of Shimura varieties of abelian type ([Kisinintegral]). In [PappasRapoport], Pappas and Rapoport lay out a framework for integral models for all Shimura varieties, and amongst various other things, they show that the integral models constructed by Kisin satisfy their framework. In [Lovering], Lovering constructs integral canonical models for automorphic data associated to Shimura varieties of abelian type. The extension property proven in most of these works ultimately relies on the work of Chai and Faltings, which uses the moduli intepretation of $\mathcal{A}_{g}$, and therefore does not carry over to exceptional Shimura varieties.
1.1 The extension property
We now define our extension property. We note that our extension property is the equivalent to the one formulated by Milne-Moonen, and we include a discussion about this fact below. In the proper case, the extension property is very simple:
Definition 1.1.
Let $F$ be a local field. We say that an integral model $\mathscr{S}_{K}(G,X)/\mathcal{O}_{F}$ of a proper Shimura variety $S_{K}(G,X)/F$ has the extension property if for every smooth scheme $\mathcal{T}/\mathcal{O}_{F}$, every map over the generic fiber $\phi\in\operatorname{Hom}(\mathcal{T}_{F},S_{K}(G,X))$ extends to a map from $\mathcal{T}$ to $\mathscr{S}_{K}(G,X)$.
In the non-proper case, one must take care to ensure that the special fiber doesn’t map to the “boundary” of $\mathscr{S}_{K}(G,X)$. For primes $\ell\neq p$, we fix an $\ell$-adic automorphic local system ${}_{{\mathrm{et}}}\mathbb{V}_{\ell}$ on $\mathscr{S}_{K}(G,X)$ corresponding to a faithful representation $V$ of $G^{\operatorname{ad}}$ (see Section 5.1 for precise definitions of ${}_{{\mathrm{et}}}\mathbb{V}_{\ell}$).
Definition 1.2.
We say that a model $\mathscr{S}_{K}(G,X)/\mathcal{O}_{F}$ of $S_{K}(G,X)$ satisfies the extension property if for every smooth scheme $\mathcal{T}/\mathcal{O}_{F}$, and every map over the generic fiber $\phi\in\operatorname{Hom}(\mathcal{T}_{F},S_{K}(G,X))$ such that $\phi^{*}\mathbb{V}_{\ell}$ extends to $\mathcal{T}$ for some $\ell\neq p$, the map $\phi$ extends to a map from $\mathcal{T}$ to $\mathscr{S}_{K}(G,X)$.
1.2 Main Results for Shimura Varieties
Our main theorem is to prove the existence of integral canonical models for arbitrary Shimura varieties $S$, for sufficiently large primes. Concretely, we prove the following:
Theorem 1.3.
Let $S$ be a Shimura variety over its reflex field $E$. For almost all finite places $v$, there exists a model $\mathcal{S}$ over $\mathcal{O}_{v}$ which has the extension property, and which admits a log-smooth compactification over $\mathcal{O}_{v}$.
Moreover, for the places $v$ as above, $\mathcal{S}$ is the unique smooth model of $S$ over $\mathcal{O}_{v}$ which admits a log-smooth compactification over $\mathcal{O}_{v}$ and to which the local system ${}_{{\mathrm{et}}}\mathbb{V}_{\ell}$ extends.
Note that the uniqueness claim implies that any global model of $S_{K}(G,X)$ will specialize to the integral canonical model at almost all places.
While our theorem applies to all Shimura varieties, the fact that it only applies for almost all finite places means that much stronger results ([Kisinintegral]) are known in the abelian setting. Therefore, the case of greatest interest is when $S$ is a Shimura variety of exceptional type.
1.2.1 Applications
We deduce several applications for mod $v$ (and $v$-adic) Shimura varieties.
Firstly, we prove the following unramified $p$-adic analogue of Borel’s theorem, which we will state in the proper case for ease of exposition (see Section 6 for a more general statement)
Theorem 1.4.
Let $S_{K}(G,X)$ be an exceptional proper Shimura variety. Let $\mathsf{D}$ denote the closed unit rigid-analytic disc, and let $\mathsf{D}^{\times}=\mathsf{D}\setminus\{0\}$. Let $f:\mathsf{D}^{\times}\rightarrow S_{K}(G,X)^{{\mathrm{an}}}$ denote a rigid-analytic map defined over $W(\overline{\mathbb{F}}_{p})[1/p]$. Then, $f$ extends to a map $\mathsf{D}\rightarrow S_{K}(G,X)^{{\mathrm{an}}}$.
In fact we also prove a version of Theorem 1.4 in the context of integral models of images of period maps (see Theorem 1.9).
We also prove results pertaining to the mod $v$ arithmetic and geometry of $\mathscr{S}_{K}(G,X)$. The main theorem implies that Hecke correspondences away from $p$ extend to $\mathscr{S}_{K}(G,X)/\mathcal{O}_{E_{v}}$. We use these Hecke correspondences to prove the following theorem:
Theorem 1.5.
Let $x\in\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})$ be any point. Then, the Frobenius endomorphism on ${}_{{\mathrm{et}}}\mathbb{V}_{\ell,\bar{x}}$ is semi-simple.
Work of Esnault-Groechenig shows that $\mathbb{V}_{p}/S_{K}(G,X)$ is a crystalline local system, which yields a relative Faltings-Fontaine-Laffaille on the $p$-adic completion of $\mathscr{S}_{K}(G,X)$. Analogous to [Kisinintegral, Section 1.5], we give an explicit description of the this Faltings-Fontaine-Laffaille module restricted to complete local rings of $\mathscr{S}_{K}(G,X)$ at closed points (see Theorem 8.9 for the precise statement). We then use this in conjunction with the semi-simplicity of $\ell$-adic Frobenii, to prove the following theorem (which generalizes the Serre-Tate canonical lift).
Theorem 1.6.
The $\mu$-ordinary locus in $\mathscr{S}_{K}(G,X)_{\mathbb{F}_{v}}$ is open dense. Let $x\in\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})$ denote a $\mu$-ordinary point. Then, there exists a canonical special point $\tilde{x}\in\mathscr{S}_{K}(G,X)(W(\mathbb{F}_{q}))$ that lifts $x$.
The above theorem then immediately has the following consequence.
Theorem 1.7.
Let $x\in\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})$ denote a $\mu$-ordinary point, and let ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$ denote the $F$-crystal at $x$ (associated to the Fontaine-Laffaille module defined by ${}_{{\mathrm{et}}}\mathbb{V}_{p}/S_{K}(G,X)$). Then, the crystalline $q$-Frobenius endomorphism on ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$ is semisimple.
Finally, we also have an analogue of Tate’s isogeny theorem for ordinary points.
Theorem 1.8.
Let $x$ and $y$ be ordinary mod $p$ points whose $\ell$-adic Galois representations are isomorphic. Then a rational Hodge structures underlying the canonical lifts $\tilde{x}$ and $\tilde{y}$ are isomorphic.
This result should be thought of as an exceptional analogue (in the ordinary case) of Tate’s isogeny theorem, which states that two abelian varieties over finite fields are isogenous if the $\ell$-adic Galois representations associated to the abelian varieties are isomorphic.
1.3 Comparison with the extension property of Milne-Moonen
The usual (Milne-Moonen) phrasing of the extension property is to let the level $K$ be of the form $K=K^{p}K_{p}$ where $K_{p}\subset G(\mathbb{Q}_{p})$ and $K^{p}\subset G(\mathbf{A}^{p}_{f})$, and consider the tower of varieties $\mathscr{S}_{K}(G,X)$ where $K_{p}$ is fixed and $K^{p}$ varies. One then asks for a tower of integral models with finite etale transition maps, and that a map from the generic fiber of any scheme $\mathcal{T}$ to this tower extends. This however entails defining a category of admissible test schemes to be able to map to this entire tower, which is something like “pro-smooth schemes over $\mathbf{Z}_{p}$”, and this requires some care to do precisely. Geometrically these are more complicated (pro-) objects, which is why we choose to work with our definition.
[…]
1.4 Integral canonical models for images of geometric period maps
Our methods in fact provide integral canonical models at sufficiently large primes for many varieties which are interpretable as moduli spaces of Hodge structures. We describe the case of variations coming from geometry, but first give some general background from Hodge theory (see Section 2.7 for more details).
Suppose an irreducible normal complex variety $P$ (which we can without loss of generality assume smooth) is equipped with a polarizable integral variation of Hodge structures $(\mathbb{W}_{\mathbb{Z}},F^{\bullet})$. The associated period map $P^{\mathrm{an}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$ will factor as $P^{\mathrm{an}}\to Y^{\mathrm{an}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$ where $Y$ is a generically inertia-free Deligne–Mumford stack with quasiprojective coarse moduli space, $P\to Y$, has connected generic geometric fiber, and $Y^{\mathrm{an}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$ is finite [bbt]. This map is obtained by taking a resolution $\tilde{P}\to P$ and then a partial compactification $\tilde{P}\subset P^{\prime}$ for which the period map extends to a proper period map and forming the Stein factorization. In particular, the variation $(\mathbb{W}_{\mathbb{Z}},F^{\bullet})$ is pulled back from $(\mathbb{V}_{\mathbb{Z}},F^{\bullet})$ on $Y$. In general $\tilde{P}$ may be a stack as well, but the inertia of $Y$ acts faithfully on $\mathbb{V}_{\mathbb{Z}}$, so $Y$ is a global quotient of a normal quasiprojective variety by the action of a finite group, and after passing to a finite étale cover of $P$ (for instance by adjoining level structure) we may assume $Y$ (and therefore $P$) is a quasiprojective variety. We call such a $Y$ the Stein factorization of the period map.
Now let $P$ be an irreducible normal variety defined over a number field $E$, $f:Z\to P$ a smooth projective $F$-morphism, and for any fixed $k$ let ${}_{et}\mathbb{W}_{\ell}:=R^{k}f_{*}\mathbb{Z}_{\ell}$ be the arithmetic local system—or more generally a subquotient obtained from maps between such local systems induced by $P$-morphisms of families [^1] We lay out the precise conditions for a variation of Hodge structures to “come from geometry” in the sense we need it in Section 4.. We let $Y$ be the Stein factorization of the period map of the variation with underlying geometric local system $\mathbb{V}_{\mathbb{Z}}=R^{k}f^{\mathrm{an}}_{*}\mathbb{Z}$ for some embedding $E\subset\mathbb{C}$. Then $Y$ is defined over $E$, does not depend on the embedding, and the local system ${}_{et}\mathbb{W}_{\ell}$ descends to $Y$.
For example, for any moduli space $P$ admitting an infinitesimal Torelli theorem (including moduli spaces of Calabi–Yau varieties and most moduli spaces of complete intersections in $\mathbb{P}^{n}$, to name a couple), $Y$ is a canonical partial compactification (namely, the natural “Hodge–theoretic” compactification) of $P$.
Theorem 1.9.
Let $E$ be a number field. Let $Y/F$ be the Stein factorization of the period map of a geometric variation of Hodge structures which we assume to be inertia-free (and therefore a quasiprojective variety). Let ${}_{et}\mathbb{V}_{\ell}$ be the descent of the natural $\ell$-adic arithmetic local system to $Y$.
  1. 1.
    There is a finite set of primes $\Sigma_{Y}$ of $\mathcal{O}_{E}$ such that, for all finite places $v$ of $\mathcal{O}_{E,\Sigma_{Y}}$, there is a model $\mathcal{Y}_{v}$ over $\mathcal{O}_{v}$ such that:
    1. (a)
      Each ${}_{et}\mathbb{V}_{\ell}$ extends to $\mathcal{Y}_{v}$ for $\ell\neq p$, where $p$ is the integral prime below $v$.
    2. (b)
      For any irreducible smooth $\mathcal{X}/\mathcal{O}_{v}$ with generic point $\eta_{\mathcal{X}}$, any morphism $\phi_{0}:\eta_{\mathcal{X}}\to Y_{E_{v}}$ extends to $\phi:\mathcal{X}\to\mathcal{Y}_{v}$ if and only if $\phi_{0}^{*}(_{et}\mathbb{V}_{\ell})$ extends to $\mathcal{X}$ for some $\ell\neq p$.
    3. (c)
      There exists a global model $\mathcal{Y}$ of $Y$ which restricts to $\mathcal{Y}_{v}$ for all places $v$ of $\mathcal{O}_{E,\Sigma_{Y}}$.
  2. 2.
    For any $\mathcal{O}_{v}$-model $\mathcal{Z}_{v}$ of $Y$ that admits a uniform log smooth resolution (see Definition 2.12) and to which ${}_{et}\mathbb{V}_{\ell}$ extends for some $\ell\neq p$, the identity map over the generic fiber extends to a $\mathcal{O}_{v}$-morphism $\mathcal{Z}_{v}\to\mathcal{Y}_{v}$.
  3. 3.
    Let $Y,Y^{\prime}/E$ be the Stein factorizations of two possibly different period maps arising as above, and for a finite place $v$ of $\mathcal{O}_{E,\Sigma_{Y}\cup\Sigma_{Y^{\prime}}}$ let $\mathcal{Y}_{v},\mathcal{Y}^{\prime}_{v}$ be models as in (1). Then any morphism $\phi_{0}:\eta_{\mathcal{Y}}\to\mathcal{Y}^{\prime}_{v}$ extends to $\phi:\mathcal{Y}_{v}\to\mathcal{Y}^{\prime}_{v}$ if and only if $\phi_{0}^{*}(_{et}\mathbb{V}_{\ell})$ extends to $\mathcal{Y}_{\mathcal{O}_{v}}$ for some $\ell\neq p$. In particular, the local model $\mathcal{Y}_{v}$ of $Y$ is canonically determined.
In general, the image of a period map has no reason to be smooth (already over $\mathbb{C}$), so it is important to allow for singularities. Even starting with a smooth moduli space $S$ parametrizing varieties with an infinitesimal Torelli theorem so that the period map $S^{\mathrm{an}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$ is immersive, it will often not be proper; the “Hodge-theoretic” compactification is essentially the closure of the image of $S^{\mathrm{an}}\to\mathbf{G}(\mathbb{Z})\backslash D$ (up to a finite map) and will therefore usually be singular anyway. This is already true for the image of the Torelli map for the moduli space of curves (see 2.11).
1.5 A purity property for Fontaine–Laffaille modules
The extension property in the settings of Theorem 1.3 and Theorem 1.9 boils down to a purity property of a $p$-torsion Fontaine–Laffaille module pulled back from a variety on which it has immersive Kodaira–Spencer map. It is motivated by classical Hodge theory and the arguments of [dejongoort, moretbailly, faltingschai]. Briefly, let $F$ be a mixed characteristic local field with ring of integers $\mathcal{O}$ and residue field $k$, $\mathcal{X}/\mathcal{O}$ a smooth scheme, and $f_{F}:X\to Y$ a morphism over the generic point to a model of an image of a period map $\mathcal{Y}$. There are three steps to proving the extension of $f_{F}$ to a morphism $f:\mathcal{X}\to\mathcal{Y}$:
  1. 1.
    $\mathcal{Y}$ will support an $\ell$-adic local system ${}_{et}\mathbb{V}_{\ell}$ with large local monodromy at the boundary. It then follows (assuming the existence of a reasonable resolution and log smooth compactification) that the map extends in codimension one as soon as $f_{F}^{*}(_{et}\mathbb{V}_{\ell})$ extends to $\mathcal{X}$. This in particular establishes the claim if $\dim\mathcal{X}=1$.
  2. 2.
    If $\dim\mathcal{X}>2$, then assuming the result for $\dim\mathcal{X}=2$, the conclusion follows by induction. An extension is unique if it exists, and by assumption exists on a general hyperplane section $\mathcal{H}$. Deformation theory can then be used to locally lift the extension to the completion of $\mathcal{X}$ along $\mathcal{H}$, which is enough to deduce the existence of the extension to $\mathcal{X}$. This is the same argument used by [dejongoort, moretbailly, faltingschai].
  3. 3.
    The main step is then to handle the case $\dim\mathcal{X}=2$, and here we must use that $\mathcal{Y}$ supports a crystalline local system whose associated Fontaine–Laffaille module has immersive Kodaira–Spencer map mod $p$. In this case, $\mathcal{X}/\mathcal{O}$ is a family of smooth curves, and can be extended on a blow-up $\mathcal{X}^{\prime}$ at points in the special fiber. It is proven in [OV] that the strong divisibility property of $p$-torsion Fontaine–Laffaille modules implies a stability property: they cannot contain positive degree subsheaves. Fontaine–Laffaille modules cannot be pulled back along maps from nodal curves if the lift to $W_{2}$ smooths out some nodes, since there are no longer local lifts of Frobenius and therefore the inverse Cartier transform no longer makes sense. We show however it exists in a weaker sense up to twisting by the nodes, and this is enough to rule out nodal curves in $\mathcal{Y}$ that lift to $W_{2}$ and smooth in a controlled way. In particular it follows that the exceptional curves must be contracted—otherwise, the positive degree tangent bundle of a rational curve would contradict the stability property we prove. For a precise statement, see Proposition 3.9.
1.6 Comparison with past work
In the Abelian case which Kisin handles, proving the existence of smooth models is difficult, whereas the extension property follows directly from the extension property for $\mathcal{A}_{g}$. Unlike [Kisinintegral], we don’t directly address the question of for which places $v$ the Shimura variety $S_{K}(G,X)$ admits smooth models over $\mathcal{O}_{v}$. Instead, we “spread out” the global model $S_{K}(G,X)$ over $\mathcal{O}_{E}[1/N]$ for some large $N$, which in particular guarantees smoothness at primes $v\nmid N$ (amongst other good properties). However, the lack of an interpretation of our Shimura varieties as a moduli space means we need an entirely different argument for establishing the extension property.
Regarding the applications to the mod $p$ and $p$-adic geometry of Shimura varieties, Kisin ([Kisinmodppoints]) proves that every $\overline{\operatorname{\mathbb{F}}}_{p}$-point of abelian Shimura varieties admit CM lifts up to isogeny. Note that the up-to-isogeny requirement is crucial as work of Oort [oortJAG] shows (for a more general statement, see also [KLSS]). Noot in [Noot] proves the existence of CM lifts (on the nose) for points of Hodge-type Shimura varieties that intersect the ordinary locus of $\mathcal{A}_{g}$. Noot’s theorem was generalized to $\mu$-ordinary points by [MoonenSerreTate] in the PEL-type case to the and [ShankarZhou] in the Hodge type case. In [padicBorel], the authors prove the $p$-adic extension and algebraicity theorems for maps to Shimura varieties (and Rapoport-Zink spaces) of abelian type defined over discretely valued $p$-adic fields and where the source of the map can have any dimension.
1.7 Acknowledgements
B. B. benefited from many conversations with Yohan Brunebarbe throughout, and first learned of the work of Ogus–Vologodsky from Raju Krishnamoorthy. We are also grateful to Michael Groechenig, Kiran Kedlaya, Mark Kisin, Arthur Ogus, Abhishek Oswal, and Vadim Vologodsky for helpful conversations. B. B. was partially supported by NSF DMS-2131688. A. S. was partially supported by the NSF grant DMS-2100436, and would like thank MSRI for their hospitality during the spring of 2023.
1.8 Outline
In Section 2 we formulate the definition of an integral canonical model in terms of the extension property, and show that extension in codimension 1 can be controlled in many cases (including those of interest) by a local system. We also discuss the descent and extension of arithmetic local systems to integral models of period images, and the relation of the extension property to purity results in classical Hodge theory. In Section 3 we prove some weak semistability results for the pull-backs of Fontaine–Laffaille modules to nodal curves. In Section 4 we formulate precise conditions under which the presence of a Fontaine–Laffaille module on a local model implies the extension property, and in Section 5 we apply it to the case of Shimura varieties and period images to prove the existence of integral canonical models as in Theorem 1.3 and Theorem 1.9. In Section 6 we use the same semistability results to prove Theorem 1.4. In Section 7 we prove Hecke correspondences extend and use this to prove Theorem 1.5. In Sections 8 and 9, we analyze the complete local rings of the integral canonical models of Shimura varieties at closed points and prove Theorems 1.6, 1.7, and 1.8.
1.9 Notation
With the exception of Section 3, we use unadorned Roman letters $S,Y$ to denote algebraic varieties over a field $F$ (typically a number field or a characteristic 0 local field), script letters $\mathcal{S},\mathcal{Y}$ to denote their integral models over a ring of integers $\mathcal{O}_{F}$ in $F$, and a subscript (either $\mathcal{S}_{v},\mathcal{Y}_{v}$ or $\mathcal{S}_{0},\mathcal{Y}_{0}$) to denote the special fiber—the base-change to the residue field $k$ of $\mathcal{O}_{F}$—at a finite place $v$ (or 00 in the local situation, when there is no ambiguity). In section Section 3, we denote by unadorned Roman letters $S,Y$ varieties over a perfect characteristic $p$ field $k$, and by $\tilde{S},\tilde{Y}$ a lift to $W_{2}(k)$.
2 The extension property and natural boundary
In this section we single out a class of integral models for which extension on the level of fundamental groups is the only obstruction to extending morphisms defined over the generic point. In particular, such models will be uniquely determined when they exist. The key property these models will have is the extension of morphisms from smooth schemes, just as in the work of Milne and Moonen. In the Shimura case (or whenever the model is smooth) this property obviously uniquely determines the model; in the general case it will be important to allow for singularities, as even for smooth moduli spaces, the extension property will often force a partial compactification which is not smooth—for instance, see 2.11. We also discuss how the extendability on the level of fundamental groups can be detected by the extendability of local systems.
2.1 The extension property
Definition 2.1.
Let $\mathcal{O}$ be a complete DVR with fraction field $F$ and residue field $k$.
  1. 1.
    Let $\mathcal{Y},\mathcal{Y}^{\prime}/\mathcal{O}$ be finite type flat $\mathcal{O}$-schemes. An $F$-morphism $f_{F}:Y\to Y^{\prime}$ of the generic fibers is $\mathcal{O}$-admissible if the maps on fundamental groups fits into a diagram
    ${\pi_{1}^{{\mathrm{et}}}(Y,\bar{y})}$${\pi_{1}^{{\mathrm{et}}}(Y^{\prime},\bar{y}^{\prime})}$${\pi_{1}^{{\mathrm{et}}}(\mathcal{Y},\bar{y})}$${\pi_{1}^{{\mathrm{et}}}(\mathcal{Y}^{\prime},\bar{y}^{\prime})}$
    for compatibly chosen basepoints.
  2. 2.
    We say a finite type flat $\mathcal{O}$-scheme $\mathcal{Y}/\mathcal{O}$ satisfies the smooth extension property over $\mathcal{O}$ if for any smooth $\mathcal{O}$-scheme $\mathcal{X}/\mathcal{O}$, any $\mathcal{O}$-admissible $F$-morphism $f_{F}:X\to Y$ on the generic fiber extends to an $\mathcal{O}$-morphism $f:\mathcal{X}\to\mathcal{Y}$.
2.2 Natural boundary and extension in codimension 1
In the presence of a sufficiently well-behaved log smooth compactification of a resolution (one with “natural boundary” as defined below), admissible morphisms with smooth source will extend over codimension 1 points, and this will be the first step of the proof of the extension property in the setting of section 4.
The following definition is made with a view towards the general case where $\mathcal{Y}$ is not smooth, but for Shimura varieties we will have $\mathcal{S}=\mathcal{Y}$.
Definition 2.2.
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$.
  1. 1.
    Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme. A resolution of $\mathcal{Y}/\mathcal{O}$ is a smooth $\mathcal{O}$-scheme $\mathcal{S}/\mathcal{O}$ with a proper dominant map $\pi:\mathcal{S}\to\mathcal{Y}$ such that $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}\pi_{*}\mathcal{O}_{\mathcal{S}}$ via the natural map.
  2. 2.
    We say $(\bar{\mathcal{S}},\mathcal{D})/\mathcal{O}$ is log smooth if it is étale locally isomorphic over $\mathcal{O}$ to $(\mathbb{A}^{n}_{\mathcal{O}},\mathcal{D})$, where $\mathcal{D}$ is a union of coordinate hyperplanes.
  3. 3.
    Let $\mathcal{S}/\mathcal{O}$ be a smooth $\mathcal{O}$-scheme. By a log smooth compactification we mean a log smooth proper $\mathcal{O}$-scheme $(\bar{\mathcal{S}},\mathcal{D})/\mathcal{O}$ with $\bar{\mathcal{S}}\setminus\mathcal{D}=\mathcal{S}$ (as an $\mathcal{O}$-scheme).
  4. 4.
    Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme, $\pi:\mathcal{S}\to\mathcal{Y}$ a resolution, and $\bar{\mathcal{S}}/\mathcal{O}$ a log smooth compactification of $\mathcal{S}$. We say that the boundary of $\bar{\mathcal{S}}$ is natural (with respect to $\mathcal{Y}$) if for every generic point $\eta$ of a stratum of $\mathcal{D}$, the geometric inertia of $\eta$ injects in $\pi_{1}^{{\mathrm{et}}}(\mathcal{Y},\bar{y})$ for a chosen geometric basepoint $\bar{y}$ of $\mathcal{Y}$.
Lemma 2.3.
Let $\mathcal{O}$ be a characteristic 0 complete DVR with fraction field $F$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme, $\pi:\mathcal{S}\to\mathcal{Y}$ a resolution, and $\bar{\mathcal{S}}/\mathcal{O}$ a log smooth compactification of $\mathcal{S}$ with natural boundary. Let $\mathcal{X}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme and $f_{F}:X\to Y$ an $\mathcal{O}$-admissible morphism over the generic point. Then there is a finite type flat $\mathcal{X}^{\prime}/\mathcal{O}$, a proper dominant generically finite morphism $\mathcal{X}^{\prime}\to\mathcal{X}$, and a morphism $\mathcal{X}^{\prime}\to\mathcal{S}$ such that the following diagram on the generic fibers commutes:
[…]
Moreover, if the image of $f_{F}$ meets the locus of $Y$ where $\pi$ is an isomorphism, then $\mathcal{X}^{\prime}\to\mathcal{X}$ can be taken to be an isomorphism over a dense open set.
Proof.
The map $f_{F}:X\to Y$ lifts to $S$ over a dense open subset of $X$; let $\mathcal{W}\subset\mathcal{X}\times_{\mathcal{O}}\bar{\mathcal{S}}$ be the closure of the graph. We then have that $\mathcal{W}\to\mathcal{X}$ is proper dominant with a morphism $\mathcal{W}\to\bar{\mathcal{S}}$, and taking an appropriate complete intersection of ample hypersurfaces, we may find $\mathcal{X}^{\prime}\subset\mathcal{W}$ which is proper dominant generically finite over $\mathcal{X}$. If $X$ meets the locus where $\pi$ is an isomorphism, then we take $\mathcal{X}^{\prime}=\mathcal{W}$ and the morphism $\mathcal{X}^{\prime}\to\mathcal{X}$ is a generic isomorphism. Either way, if the boundary of $\bar{\mathcal{S}}$ intersects $\mathcal{X}^{\prime}$ nontrivially, then there is a finite extension $\mathcal{O}^{\prime}$ of $\mathcal{O}$ an $\mathcal{O}^{\prime}$-point of $\mathcal{X}^{\prime}/\mathcal{O}$ whose closed point maps to the boundary. Because the boundary is natural, by the argument of [PST, Thm 4.4] this contradicts the admissibility assumption.
Corollary 2.4.
In the setting of the lemma, assuming the image of $f_{F}$ meets the locus of $Y$ where $\pi$ is an isomorphism, there is a codimension $\geq 2$ subset $\mathcal{Z}\subset\mathcal{X}$ and an extension of $f_{F}$ to a morphism $f:\mathcal{X}\setminus\mathcal{Z}\to\mathcal{Y}$.
Proof.
$\mathcal{X}^{\prime}\to\mathcal{X}$ is proper so every codimension 1 point $\eta$ lifts to $\mathcal{X}^{\prime}$ and maps to $\mathcal{S}$, so $f_{F}$ extends over $\eta$. ∎
2.3 Natural boundary and local systems
For $\mathcal{S}/\mathcal{O}$ smooth over a Dedekind domain and $\bar{\mathcal{S}}$ a log smooth compactification, both the naturality of the boundary and the admissibility condition on $\mathcal{O}$-morphisms to $\mathcal{S}$ can be checked using an $\ell$-adic local system $\mathbb{V}_{\ell}$ on $\mathcal{S}$. The key property such a local system must have is “nonextendability” over the generic fiber, meaning it does not extend to any nontrivial partial compactification. This notion is treated carefully in [brunebarbe, §3] (where it is referred to as “maximality”); the following definition is easily seen to be equivalent to the definition therein.
Definition 2.5.
Let $Y/F$ be a normal variety with an $\ell$-adic local system $\mathbb{V}_{\ell}$. We say $\mathbb{V}_{\ell}$ has faithful geometric local monodromy if for some (hence any) resolution $\pi:S\to Y$ and log smooth compactification $S\subset\bar{S}$, the geometric inertia at the generic point of every stratum of the boundary $D$ acts faithfully on $\pi^{*}\mathbb{V}_{\ell}$. We say an $\ell$-adic local system on an integral model has faithful geometric local monodromy if the restriction to the generic fiber does.
Note that faithfulness of geometric local monodromy can be checked over the algebraic closure $\bar{F}$, or even over $\mathbb{C}$ for a choice of embedding $F\subset\mathbb{C}$. Faithfulness of geometric local monodromy restricts well:
Lemma 2.6 ([brunebarbe, §3]).
Let $Y/\mathbb{C}$ be a normal variety and $\mathbb{V}_{\ell}$ an $\ell$-adic local system on $Y$. Then the following are equivalent:
  1. 1.
    $\mathbb{V}_{\ell}$ has faithful geometric local monodromy.
  2. 2.
    For any curve $C/\mathbb{C}$ and any proper morphism $f:C\to Y$, $f^{*}\mathbb{V}_{\ell}$ has faithful geometric local monodromy.
  3. 3.
    For any normal variety $X/\mathbb{C}$ and any proper morphism $f:X\to Y$, $f^{*}\mathbb{V}_{\ell}$ has faithful geometric local monodromy.
The next lemma interpets naturality of the boundary and admissibility in terms of the local system as promised.
Lemma 2.7.
Suppose $\mathcal{Y}/\mathcal{O}$ is a normal finite type flat $\mathcal{O}$-scheme, $\pi:\mathcal{S}\to\mathcal{Y}$ a resolution, $\bar{\mathcal{S}}/\mathcal{O}$ a log smooth compactification of $\mathcal{S}$, and $\mathbb{V}_{\ell}$ an $\ell$-adic local system on $\mathcal{Y}$ such that $\mathbb{V}_{\ell}|_{Y}$ has faithful geometric local monodromy. Then
  1. 1.
    $\bar{\mathcal{S}}$ has natural boundary.
  2. 2.
    For $\mathcal{X}/\mathcal{O}$ regular, a morphism $f_{K}:X\to Y$ over the generic point which meets the locus of $Y$ where $\pi$ is an isomorphism is $\mathcal{O}$-admissible if and only if $f_{K}^{*}\mathbb{V}_{\ell}$ extends to $\mathcal{X}$.
Proof.
For 1, the geometric inertia of every stratum acts faithfully on $\pi_{K}^{*}\mathbb{V}_{\ell}$, and this action factors through $\pi_{1}^{{\mathrm{et}}}(\mathcal{S},\bar{s})\to\pi_{1}^{{\mathrm{et}}}(\mathcal{Y},\bar{y})$, so it follows that the boundary is natural.
For 2, the forward implication is obvious. For the reverse implication, the same argument as in Lemma 2.3 shows that $f_{F}$ extends to $f:\mathcal{X}\setminus\mathcal{Z}\to\mathcal{Y}$ for a codimension $\geq 2$ subset $\mathcal{Z}$, but by purity of the branch locus $\pi_{1}^{{\mathrm{et}}}(\mathcal{X}\setminus\mathcal{Z},\bar{x})\to\pi_{1}^{{\mathrm{et}}}(\mathcal{X},\bar{x})$ is an isomorphism. ∎
Corollary 2.8.
For $\mathcal{Y},\mathbb{V}_{\ell}$ as in the lemma, $\mathcal{X}/\mathcal{O}$ a normal finite type flat $\mathcal{O}$-scheme, and $f_{F}:X\to Y$ a morphism over the generic point such that $f_{F}^{*}\mathbb{V}_{\ell}$ extends to $\mathcal{X}$. Then there is a codimension $\geq 2$ subset $\mathcal{Z}\subset\mathcal{X}$ and an extension of $f_{F}$ to a morphism $f:\mathcal{X}\setminus\mathcal{Z}\to\mathcal{Y}$.
2.4 Integral canonical models of smooth varieties
For simplicity, we first define integral canonical models in the smooth case (to be generalized in the next section).
Definition 2.9.
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$ and $\mathcal{Y}/\mathcal{O}$ a quasiprojective smooth $\mathcal{O}$-scheme. We say $\mathcal{Y}/\mathcal{O}$ is an integral canonical model if for every closed point $v$ of $\operatorname{Spec}\mathcal{O}$, the base-change $\mathcal{Y}_{v}/\mathcal{O}_{v}$ satisfies the extension property over $\mathcal{O}_{v}$.
The following is immediate:
Lemma 2.10.
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$. Let $\mathcal{S},\mathcal{S}^{\prime}/\mathcal{O}$ be two smooth integral canonical models and $f_{F}:S\to S^{\prime}$ a morphism over the generic point.
  1. 1.
    $f_{F}$ extends to a morphism $f:\mathcal{S}\to\mathcal{S}^{\prime}$ if and only if there is a commutative diagram
    ${\pi_{1}^{{\mathrm{et}}}(S,\bar{s})}$${\pi_{1}^{{\mathrm{et}}}(S^{\prime},\bar{s}^{\prime})}$${\pi_{1}^{{\mathrm{et}}}(\mathcal{S},\bar{s})}$${\pi_{1}^{{\mathrm{et}}}(\mathcal{S}^{\prime},\bar{s}^{\prime})}$
    for compatibly chosen basepoints.
  2. 2.
    Suppose further that $\mathcal{S}^{\prime}$ admits a log smooth compactification and an $\ell$-adic local system $\mathbb{V}_{\ell}$ such that $\mathbb{V}_{\ell}|_{S^{\prime}}$ has faithful geometric local monodromy. Then $f_{F}$ extends to a morphism $f:\mathcal{S}\to\mathcal{S}^{\prime}$ if and only if $f_{F}^{*}(\mathbb{V}_{\ell})$ extends to $\mathcal{S}$.
2.5 Integral canonical models of normal varieties
We begin with a simple example to demonstrate that, even for a smooth variety, a model which satisfies the smooth extension property might necessarily involve a partial compactification with singularities.
Example 2.11.
Consider the moduli stacks $\mathcal{A}_{g}$ and $\mathcal{M}_{g}$ of principally polarized $g$-dimensional abelian varieties and genus $g$ curves, say over $\mathbb{Z}_{p}$ for $p\gg 0$, and let $\mathcal{M}_{g}\to\mathcal{A}_{g}$ be the Torelli map. After taking sufficient level structure, both of these stacks are schemes, and the Torelli map is generically immersive. $\mathcal{A}_{g}$ satisfies the smooth extension property, so the closure of the image of $\mathcal{M}_{g}$ in $\mathcal{A}_{g}$ will as well. This closure is quite singular however: the map $\mathcal{M}_{g}\to\mathcal{A}_{g}$ extends to the partial compactification ~$\tilde{\mathcal{M}}_{g}\to\mathcal{A}_{g}$ parametrizing compact-type nodal curves (that is, nodal curves whose dual graph is a tree), but the boundary gets contracted with positive-dimensional (non-uniruled) fibers, as the Jacobian does not record the attaching nodes.
In the normal case, to deduce the extension of arbitrary morphisms from the case of morphisms with smooth source, we will require the existence of an integral resolution with the property that the subvarieties contracted in a special fiber are limits of subvarieties contracted in the generic fiber.
Definition 2.12.
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme.
  1. 1.
    A resolution of $\mathcal{Y}/\mathcal{O}$ is a smooth $\mathcal{O}$-scheme $\mathcal{S}/\mathcal{O}$ with a proper dominant map $\pi:\mathcal{S}\to\mathcal{Y}$ such that $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}\pi_{*}\mathcal{O}_{\mathcal{S}}$ via the natural map.
  2. 2.
    A resolution $\pi:\mathcal{S}\to\mathcal{Y}$ is uniform if there is a flattening stratification of $\pi$ by locally closed subsets $\mathcal{Y}^{j}$ of $\mathcal{Y}$, each of which is flat over $\operatorname{Spec}\mathcal{O}$.
Lemma 2.13.
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$, $Y/F$ a normal variety, $\pi_{F}:S\to Y$ a resolution, and $\pi:\mathcal{S}\to\mathcal{Y}$ a model of $\pi_{F}$ defined over $\mathcal{O}$. Then $\pi$ is a uniform resolution over a (nonempty) open set of $\mathcal{O}$.
Proof.
Since $\mathcal{O}_{\mathcal{Y}}\to\pi_{*}\mathcal{O}_{\mathcal{S}}$ is an isomorphism over the generic point, it is over an open set of $\operatorname{Spec}\mathcal{O}_{F}[1/N]$. Taking a flattening stratification $\mathcal{Y}^{j}$, the strata which do not dominate map to finitely many primes. ∎
Definition 2.14.
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$ and $\mathcal{Y}/\mathcal{O}$ a normal quasiprojective flat $\mathcal{O}$-scheme. We say $\mathcal{Y}/\mathcal{O}$ is an integral canonical model if:
  1. 1.
    $\mathcal{Y}$ admits a uniform resolution $\pi:\mathcal{S}\to\mathcal{Y}$.
  2. 2.
    For every closed point $v$ of $\operatorname{Spec}\mathcal{O}$, the base-change $\mathcal{Y}_{v}/\mathcal{O}_{v}$ satisfies the extension property over $\mathcal{O}_{v}$.
Note that if $\mathcal{Y}/\mathcal{O}$ is smooth, the condition on the existence of a uniform resolution is automatic, so this definition generalizes Definition 2.9. The next proposition generalizes Lemma 2.10 to the normal case.
Proposition 2.15.
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$. Let $\mathcal{Y}/\mathcal{O}$ be a normal quasiprojective flat $\mathcal{O}$-scheme admitting a uniform resolution, $\mathcal{Y}^{\prime}/\mathcal{O}$ an integral canonical model, and $f_{F}:Y\to Y^{\prime}$ a morphism over the generic point.
  1. 1.
    $f_{F}$ extends to a morphism $f:\mathcal{Y}\to\mathcal{Y}^{\prime}$ if and only if there is a commutative diagram
    ${\pi_{1}^{{\mathrm{et}}}(Y,\bar{y})}$${\pi_{1}^{{\mathrm{et}}}(Y^{\prime},\bar{y}^{\prime})}$${\pi_{1}^{{\mathrm{et}}}(\mathcal{Y},\bar{y})}$${\pi_{1}^{{\mathrm{et}}}(\mathcal{Y}^{\prime},\bar{y}^{\prime})}$
    for compatibly chosen basepoints.
  2. 2.
    Suppose further that $\mathcal{Y}^{\prime}$ admits a uniform resolution $\pi^{\prime}:\mathcal{S}^{\prime}\to\mathcal{Y}^{\prime}$, a log smooth compactification of $\mathcal{S}^{\prime}$, and that $\mathcal{Y}^{\prime}$ has an $\ell$-adic local system $\mathbb{V}_{\ell}$ such that $\mathbb{V}_{\ell}|_{Y^{\prime}}$ has faithful geometric local monodromy. Then $f_{F}$ extends to a morphism $f:\mathcal{Y}\to\mathcal{Y}^{\prime}$ if and only if $f_{F}^{*}(\mathbb{V}_{\ell})$ extends to $\mathcal{Y}$.
For the proof we need the following:
Lemma 2.16 (Rigidity Lemma, [debarre, Lemma 1.15]).
Let $X,Y,Z$ be integral finite type schemes, $f:X\to Y$ a proper morphism with $\mathcal{O}_{Y}\xrightarrow{\cong}f_{*}\mathcal{O}_{X}$ via the natural map, and $g:X\to Z$ a morphism which is constant on $f^{-1}(y)$ for some point $y$ of $Y$. Then for an open neighborhood $y\in U\subset Y$, $g|_{f^{-1}(U)}:f^{-1}(U)\to Z$ factors through $f|_{f^{-1}(U)}:f^{-1}(U)\to Y$.
Proof.
The proof is exactly the same as in [debarre]. Consider the (proper) morphism $h=f\times g:X\to Y\times Z$, and let $W$ be the image with projection map $p:W\to Y$. Since $\mathcal{O}_{Y}\subset p_{*}\mathcal{O}_{W}\subset f_{*}\mathcal{O}_{X}=\mathcal{O}_{Y}$, the Stein factorization of $p$ is trivial. Since $h^{-1}((p^{-1}(y)))=f^{-1}(y)$ maps to a point in $Z$, $p^{-1}(y)$ is a point, hence the fibers of $p$ are 0-dimensional in a neighborhood of $y$, and therefore $p:W\to Y$ is an isomorphism in a neighborhood of $y$. ∎
Corollary 2.17.
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$ Let $\mathcal{Y},\mathcal{Y}^{\prime}/\mathcal{O}$ be normal finite type flat $\mathcal{O}$-schemes, and $\pi:\mathcal{S}\to\mathcal{Y}$ a uniform resolution. Assume $\mathcal{Y}^{\prime}/\mathcal{O}$ is quasiprojective. Then any morphism $f:\mathcal{S}\to\mathcal{Y}^{\prime}$ which factors through $\pi_{F}:S\to Y$ over the generic point factors through $\pi$.
Proof.
Let $L^{\prime}$ be a very ample bundle on $\mathcal{Y}^{\prime}$, let $\mathcal{Y}^{j}\subset\mathcal{Y}$ be an $\mathcal{O}$-flat stratum in a flattening stratification of $\mathcal{S}\to\mathcal{Y}$, let $\mathcal{S}^{j}\to\mathcal{Y}^{j}$ be the base-change to $\mathcal{Y}^{j}$, and consider the composition $\mathcal{S}^{j}\to\mathcal{S}\to\mathcal{Y}^{\prime}$. For any fiber $F$ of $\mathcal{S}^{j}\to\mathcal{Y}^{j}$, $L|_{F}$ is globally generated, hence has a section which is generically nonzero on every geometric component. On the other hand, $L^{\vee}$ is trivial on every geometric fiber of $\mathcal{S}^{j}\to\mathcal{Y}^{j}$ over the generic point, hence by semicontinuity $L^{\vee}|_{F}$ has at least one nonzero section. The product is a section of $\mathcal{O}_{F}$ which is not identically zero, so since the fibers of $\mathcal{S}^{j}\to\mathcal{Y}^{j}$ are geometrically connected it must be nowhere zero. Thus, $L|_{F}$ is trivial, and so $f$ contracts every fiber of $\mathcal{S}\to\mathcal{Y}$. By the rigidity lemma, $f$ factors through $\mathcal{S}\to\mathcal{Y}$. ∎
Proof of Proposition 2.15.
The forward implication of part 1 is trivial; for the reverse implication, by the smooth extension property we have an extension of the composition $S\to Y\to Y^{\prime}$ to $\mathcal{S}_{v}\to\mathcal{Y}^{\prime}_{v}$ for every finite place, and therefore an extension $\mathcal{S}\to\mathcal{Y}^{\prime}$. Then by Corollary 2.17, the map $\mathcal{S}\to\mathcal{Y}^{\prime}$ factors through the desired extension $f:\mathcal{Y}\to\mathcal{Y}^{\prime}$. Part 2 follows from part 1 by Lemma 2.7.
Corollary 2.18.
Assume $Y$ admits a local system $\mathbb{V}_{\ell}$ with faithful geometric local monodromy. Then:
  1. 1.
    There is at most one integral canonical model $\mathcal{Y}/\mathcal{O}$ to which $\mathbb{V}_{\ell}$ extends and which admits a uniform resolution with a log smooth compactification.
  2. 2.
    Assume such an integral canonical model $\mathcal{Y}$ exists. Then for any normal finite type flat model $\mathcal{Z}/\mathcal{O}$ to which $\mathbb{V}_{\ell}$ extends and which admits a uniform resolution, the identity map over the generic fiber extends to a morphism $\mathcal{Z}\to\mathcal{Y}$.
2.6 Descending and extending arithmetic local systems
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$, $\mathcal{P}/\mathcal{O}$ smooth with geometrically irreducible generic fiber, and $\mathcal{Y}/\mathcal{O}$ normal finite type flat with geometrically irreducible generic fiber. Let $f:\mathcal{P}\to\mathcal{Y}$ be a proper morphism with $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}f_{*}\mathcal{O}_{\mathcal{P}}$. In this section we describe how arithmetic local systems can be extended to integral models and descended along $f$.
We start with the following:
Lemma 2.19.
Let $f:X\to Y$ be a proper morphism of schemes with $Y$ integral and assume $f$ has geometrically connected fibers and geometrically reduced generic fiber. Then over a dense open $U\subset Y$ we have the following:
  1. 1.
    The base-change $X_{U}\to U$ induces an exact sequence
    $\pi_{1}^{{\mathrm{et}}}(X_{\bar{\eta}},\bar{x})\to\pi_{1}^{{\mathrm{et}}}(X_{U},\bar{x})\to\pi_{1}^{{\mathrm{et}}}(U,\bar{u})\to 1$
    for compatibly chosen basepoints, where $\eta$ is the generic point of $Y$.
  2. 2.
    For every geometric point $\bar{u}$ of $U$, the specialization map $sp:\pi_{1}^{{\mathrm{et}}}(X_{\bar{\eta}},\bar{x})\to\pi_{1}^{{\mathrm{et}}}(X_{\bar{u}},\bar{x})$ is surjective.
Proof.
According to [stacks-project, Tag 0579], since the generic fiber is geometrically reduced we may take $U$ to be the interior of the set of points with geometrically reduced fibers. We may further assume $X_{U}\to U$ is flat by generic flatness. Then the first claim follows from the homotopy exact sequence [stacks-project, Tag 0C0J] and the second from base-change to the henselization of the local ring. ∎
The following is standard but we include it for completeness. For a morphism of schemes $X\to Y$ which induces a surjection $\pi_{1}^{{\mathrm{et}}}(X,\bar{x})\to\pi_{1}^{{\mathrm{et}}}(Y,\bar{y})$ for compatibly chosen basepoints, we denote by $I_{X/Y}\subset\pi_{1}^{{\mathrm{et}}}(X,\bar{x})$ the kernel (somewhat abusively, since we suppress the basepoints from the notation).
Lemma 2.20.
The map $f_{F}:P\to Y$ induces the following commutative diagram with exact rows and columns for compatibly chosen basepoints.
[…]
Proof.
The exactness of the rows is standard and the surjectivity of the vertical maps to the bottom row is a consequence of normality and the generic fiber being geometrically connected. ∎
Corollary 2.21.
Let $\mathbb{V}_{\ell}$ be an $\mathbb{Z}_{\ell}$-local system on $P$ for which the geometric local system $\bar{\mathbb{V}}_{\ell}$ on $P_{\bar{F}}$ descends to $Y_{\bar{F}}$. Then $\mathbb{V}_{\ell}$ descends to $Y$.
Lemma 2.22.
Assume $\mathcal{O}$ has perfect residue fields [^2] Over perfect fields, a scheme is geometrically reduced if and only if it is reduced.. Then after inverting finitely many primes of $\mathcal{O}$, any $\ell$-adic local system $\mathbb{V}_{\ell}$ on $\mathcal{P}$ which geometrically descends to $Y_{\bar{F}}$ descends to $\mathcal{Y}$. Moreover, the same is true for any further localization of $\mathcal{O}$.
Proof.
We first claim that after inverting finitely many primes, the following is true: for every geometric point $\bar{y}_{0}$ of $\mathcal{Y}$ there is a geometric point $\bar{y}$ of $Y$ specializing to it such that the specialization map $\pi_{1}^{{\mathrm{et}}}(\mathcal{P}_{\bar{y}},\bar{y})\to\pi_{1}^{{\mathrm{et}}}(\mathcal{P}_{\bar{y}_{0}},\bar{y}_{0})$ is surjective. Using Lemma 2.19, there is a stratification $\mathcal{Y}^{i}$ of $\mathcal{Y}$ such that this is true for $(\mathcal{P}_{\mathcal{Y}^{i}})^{\mathrm{red}}\to\mathcal{Y}^{i}$, provided $\mathcal{Y}^{i}$ is flat over $\mathcal{O}$. After shrinking $\mathcal{O}$, we may assume this is the case for all $\mathcal{Y}^{i}$, and therefore the claim is proven.
By Corollary 2.21, we may assume $\mathbb{V}_{\ell}$ descends to $Y$. To finish it suffices to show any finite étale cover $\mathcal{P}^{\prime}\to\mathcal{P}$ which descends to $Y$ over $P$ descends to $\mathcal{Y}$. The descent clearly patches, so we may assume $\mathcal{Y}=\operatorname{Spec}R$ for a local ring $R$. Base-changing to the henselization $\bar{R}$, the cover $\mathcal{P}^{\prime}_{R}\to\mathcal{P}_{R}$ descends to $R$ by [stacks-project, Tag 0A48], since it is trivial on every geometric fiber of $\mathcal{P}_{R}\to\operatorname{Spec}R$ by the above surjection of specialization maps. The descent to $R$ comes equipped with descent data which gives a finite étale cover of $\operatorname{Spec}R$ which pulls back to $\mathcal{P}^{\prime}$.
Corollary 2.23.
Assume $\mathcal{O}$ has perfect residue fields. Then after inverting finitely primes of $\mathcal{O}$, the following is true. For any $\ell$ and any $n$, any $\mathbb{Z}_{\ell}$-local system $\mathbb{V}_{\ell}$ on $\mathcal{P}[1/n]$ which geometrically descends to $Y_{\bar{F}}$ descends to $\mathcal{Y}[1/n]$.
2.7 Relation to Hodge theory over $\mathbb{C}$
We end this section by summarizing some classical extension results in Hodge theory using the above language. This is both motivation for the using extension properties to determine integral models and is needed to prove the characteristic 0 part of the extension in Theorem 1.9.
Lemma 2.24.
Let $P$ be a smooth complex algebraic variety equipped with a polarizable integral variation of Hodge structures $(W_{\mathbb{Z}},F^{\bullet}W)$, and let $\phi:P^{\mathrm{an}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$ be the period map.
  1. 1.
    For any partial log smooth compactification $(P^{\prime},D)$, $(W_{\mathbb{Z}},F^{\bullet})$ extends to $P^{\prime}$ if and only if the monodromy of $W_{\mathbb{Z}}$ around each component of $D$ is trivial.
  2. 2.
    $\phi$ is proper if and only if and only if $V_{\mathbb{Z}}$ has faithful geometric local monodromy.
  3. 3.
    There is a finite étale cover $\pi:\tilde{P}\to P$ and a partial log smooth compactification $\tilde{P}^{\prime}$ of $\tilde{P}$ such that $\pi^{*}W_{\mathbb{Z}}$ extends to $\tilde{P}^{\prime}$ with faithful geometric local monodromy.
  4. 4.
    If the monodromy of $W_{\mathbb{Z}}$ is unipotent, then there is a partial log smooth compactification $P^{\prime}$ of $P$ such that $W_{\mathbb{Z}}$ extends to $P^{\prime}$ with failtful geometric local monodromy.
Proof.
(1) is a result of Griffiths [griffithsiii, Theorem 9.5]. For (2), (3), and (4), see [brunebarbe, §3]. ∎
For simplicity, we will assume $W_{\mathbb{Z}}$ has unipotent local monodromy. In this case, it follows from the lemma that for any period map $\phi:P\to[\mathbf{G}(\mathbb{Z})\backslash D]$, there is a partial log smooth compactification $P^{\prime}$ of $P$, a proper extension $\phi^{\prime}:P^{\prime{\mathrm{an}}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$, and an algebraic map $f:P^{\prime}\to Y$ to a normal $Y$ with $\mathcal{O}_{Y}\to f_{*}\mathcal{O}_{P^{\prime}}$ such that $\phi^{\prime}$ factors as $P^{\prime{\mathrm{an}}}\xrightarrow{f^{\mathrm{an}}}Y^{\mathrm{an}}\xrightarrow{\psi^{\mathrm{an}}}$. We call $Y$ the Stein factorization of the period map $\phi$. Observe that the local system $W_{\mathbb{Z}}$ as well as the associated filtered flat vector bundle $(W,\nabla,F^{\bullet}W)$ are pulled back from $V_{\mathbb{Z}}$ and $(V,\nabla,F^{\bullet}V)$ on $Y$.
Corollary 2.25.
Let $Y$ be the Stein factorization of a period map $(W_{\mathbb{Z}},F^{\bullet}W)$ on $P$ with unipotent local monodromy. Let $X$ be a smooth complex algebraic variety and $U\subset X$ a dense Zariski open set. Then a morphism $\phi_{U}:U\to Y$ extends to $\phi:X\to Y$ if and only if $\pi_{U}^{*}\mathbb{W}_{\mathbb{Z}}$ extends to $X$.
Lemma 2.26.
Suppose $P$ is defined over a number field $F$.
  1. 1.
    If the filtered flat vector bundle $(W,\nabla,F^{\bullet}W)$ underling $(W_{\mathbb{Z}},F^{\bullet}W)$ is defined over $F$, so is the Stein factorization of the (compactified) period map $f:P^{\prime}\to Y$. Moreover, the descent $(V,\nabla,F^{\bullet}V)$ is also defined over $F$.
  2. 2.
    If $W_{\mathbb{Z}_{\ell}}$ extends to an arithmetic local system ${}_{et}\mathbb{W}_{\ell}$ on $P$, then $V_{\mathbb{Z}_{\ell}}$ extends to an arithmetic local system $\mathbb{V}_{\ell}$ on $Y$.
3 Weak stability of Fontaine–Laffaille modules on nodal curves
In this section we prove some semistability properties of Fontaine–Laffaille modules over nodal curves (see Section 3.3). These results will be critical in the main step in the proof of the extension of morphisms from smooth sources in codimenison $\geq 2$ in Section 4.
3.1 The inverse Cartier transform
We begin by recalling the nonabelian Hodge correspondence in characteristic $p$ of Ogus–Vologodsky [OV] via exponential twisting [LSZtwist].
Let $k$ be a perfect field of characterstic $p>2$. We denote by $W_{2}=W_{2}(k)$ the length 2 Witt vectors. Throughout, by a lift of a scheme (usually from $k$ to $W_{2}$) we mean a flat lift. For a variety $Y/k$, we denote by $F_{Y}:Y\to Y$ the absolute Frobenius, dropping the subscript if there’s no chance of confusion. We have a natural lift of Frobenius $F_{W}:W_{2}\to W_{2}$, and in the sequel by a lift of Frobenius to a lift of $Y$ we always mean a lift commuting with $F_{W}$.
3.1.1 Flat sheaves and Higgs sheaves on varieties
We first recall the definitions of Higgs sheaves and flat sheaves on smooth $k$-varieties. As $p\neq 2$ we may unambiguously speak of $\Omega_{X}^{2}$. Throughout we denote by $T_{X}:={\mathcal{H}\operatorname{om}}_{\operatorname{\mathcal{O}}_{X}}(\Omega_{X},\operatorname{\mathcal{O}}_{X})$.
By a flat sheaf on $X$ we mean a locally free $\operatorname{\mathcal{O}}_{X}$-module $V$ with a $k$-linear map $\nabla:V\to V\otimes\Omega_{X}$ which satisfies the Leibniz rule and for which the $\operatorname{\mathcal{O}}_{X}$-linear map $\nabla\wedge\nabla:V\to V\otimes\Omega_{X}^{2}$ vanishes. By a Higgs sheaf on $X$ we mean a locally free $\operatorname{\mathcal{O}}_{X}$-module $V$ with an $\operatorname{\mathcal{O}}_{Y}$-linear map $\theta:V\to V\otimes\Omega_{X}$ such that $\theta\wedge\theta:V\to V\otimes\Omega^{2}_{X}$ vanishes. Observe that since differentials pull back, both flat sheaves and Higgs sheaves pull back under arbitrary morphisms $f:X\to Y$. We denote by $MIC_{m}(X)$ (resp. $HIG_{m}(X)$) the category of flat sheaves $(V,\nabla)$ (resp. Higgs sheaves $(V,\theta)$) with $p$-curvature $\psi(\nabla):=\nabla^{p}-\nabla_{(-)^{p}}$ (resp. Higgs field $\theta$) nilpotent of exponent $\leq m$, meaning $\psi(\nabla)^{m}=0$ (resp. $\theta^{m}=0$) with the natural notion of morphisms. We will always assume $m\leq p$.
Recall that for any locally free $\mathcal{O}_{X}$-module $V$ on $X$, the pull-back along absolute Frobenius $F^{*}V$ admits a canonical flat connection $\nabla_{can}$ whose flat sections are the pull-back sections $F^{-1}V$. Explicitly, if we write $F^{*}V=\operatorname{\mathcal{O}}_{X}\otimes_{\operatorname{\mathcal{O}}_{X}}V$ where the tensor is taken with respect to the $p$-power algebra structure $\operatorname{\mathcal{O}}_{X}\to\operatorname{\mathcal{O}}_{X}$, then $\nabla_{can}=d\otimes\operatorname{id}$.
3.1.2 Nonabelian Hodge correspondence à la Lan–Sheng–Zuo
Let $Y/k$ be smooth and $\tilde{Y}/W_{2}$ a flat lift. Let $Y_{i}$ be a Zariski cover of $Y$ by affines with overlaps $Y_{ij}:=Y_{i}\cap Y_{j}$ and $\tilde{F}_{i}:\tilde{Y}_{i}\to\tilde{Y}_{i}$ lifts of absolute Frobenius. By lifting to $\operatorname{\mathcal{O}}_{\tilde{Y}}$ and applying $p^{-1}(\tilde{F}_{i}-\tilde{F}_{j})$ (resp. $p^{-1}d\tilde{F}_{i}$) we obtain $F$-derivations of $\operatorname{\mathcal{O}}_{Y_{ij}}$ (resp. $\operatorname{\mathcal{O}}_{Y_{i}}$) valued in $\operatorname{\mathcal{O}}_{Y_{ij}}$ (resp. $\Omega_{Y_{i}}$), which we may think of as $\operatorname{\mathcal{O}}_{Y}$-linear morphisms
$s_{ij}$ $:=p^{-1}(\tilde{F}_{i}-\tilde{F}_{j}):F^{*}\Omega_{Y_{ij}}\to\mathcal{O}_{Y_{ij}}$
$\zeta_{i}$ $:=p^{-1}d\tilde{F}_{i}:F^{*}\Omega_{Y_{i}}\to\Omega_{Y_{i}}$
which satisfy $ds_{ij}=\zeta_{i|j}-\zeta_{j|i}$. For a Higgs sheaf $(V,\theta)\in HIG_{m}(Y)$, it is shown in [LSZtwist] that the inverse Cartier transform $C_{\tilde{Y}}^{-1}(V,\theta)$ is the flat sheaf obtained by locally forming $(F^{*}V|_{Y_{i}},\nabla_{can}+\zeta_{i}(F^{*}\theta))$ on $Y_{i}$ and gluing via
$\exp(s_{ij}(F^{*}\theta)):(F^{*}V|_{Y_{ij}},\nabla_{can}+\zeta_{i}(F^{*}\theta))\to(F^{*}V|_{Y_{ij}},\nabla_{can}+\zeta_{j}(F^{*}\theta))$
up to negating the sign of the connection. Observe that for any Higgs subsheaf $(V^{\prime},\theta^{\prime})\subset(V,\theta)$ we have a natural inclusion of flat sheaves $C_{\tilde{Y}}^{-1}(V^{\prime},\theta^{\prime})\subset C^{-1}_{\tilde{Y}}(V,\theta)$. In particular, if $(G,0)$ is a Higgs subsheaf of $(V,\theta)$, then $(F^{*}G,\nabla_{can})$ is a flat subsheaf of $C^{-1}_{\tilde{Y}}(V,\theta)$.
3.2 An inverse Cartier transform on nodal curves
If $X$ is not smooth, the construction from the previous section fails because a given lift $\tilde{X}$ may not admit local lifts of Frobenius. If $X$ is normal, one may get some mileage out of restricting to the smooth locus as in [langer], and there is a logarithmic version due to Schepler [schepler]. However, the case of interest for us will be that of nodal curves whose nodes smooth out. In this section we construct two types of inverse Cartier transforms: the first for Higgs sheaves pulled-back from a smooth variety $Y$, and the second for Higgs subsheaves of such pull-backs to nodal curves after twisting sufficiently by the ideal of the nodes.
3.2.1 The inverse Cartier transform of pullback Higgs sheaves
Let $\tilde{Y}/W_{2}$ be a smooth $W_{2}$-scheme, $X/k$ a smooth variety, and $f:X\to Y$ a morphism. Let $(V_{Y},\theta_{Y})\in HIG_{m}(Y)$ be a Higgs sheaf on $Y$. Denote $V_{X}=f^{*}V_{Y}$ and by $\rho_{Y|X}:f^{*}\Omega_{Y}\to\Omega_{X}$ the natural map. We have two sorts of Higgs fields on $V_{X}$, each satisfying the obvious integrability condition: $\theta_{Y|X}:V_{X}\to V_{X}\otimes f^{*}\Omega_{Y}$ and $\theta_{X}:V_{X}\to V_{X}\otimes\Omega_{X}$, where $\theta_{X}=(\operatorname{id}\otimes\rho_{Y|X})\circ\theta_{Y|X}$.
Lemma 3.1.
The pullback as a flat sheaf $f^{*}C_{\tilde{Y}}^{-1}(V_{Y},\theta_{Y})$ is obtained by locally forming $(F^{*}V_{X},\nabla_{can}+\rho_{Y|X}\circ(f^{*}\zeta_{i})(F^{*}\theta_{Y|X}))$ and gluing via $\exp((f^{*}s_{ij})(F^{*}\theta_{Y|X}))$.
Proof.
Frobenius commutes with $f$. ∎
In fact the pullback twisting data may be used to define an inverse Cartier transform of a more general kind of object which we call a pullback Higgs sheaf on $X$: precisely, this is a pair $(V,\theta)$ consisting of a locally free $\operatorname{\mathcal{O}}_{X}$-module $V$ on $X$ and an $\operatorname{\mathcal{O}}_{X}$-linear map $\theta:V\to V\otimes f^{*}\Omega_{Y}$ satisfying the integrability condition. Indeed, by the same computations as in [LSZtwist], we define a flat sheaf $C^{-1}_{\tilde{Y}|X}(V,\theta)$ by locally forming $(F^{*}V,\nabla_{can}+\rho_{Y\mid X}\circ(f^{*}\zeta_{i})(F^{*}\theta))$ and gluing via $\exp((f^{*}s_{ij})(F^{*}\theta))$. If we define morphisms of pullback Higgs sheaves in the natural way, and $HIG_{m}(Y|X)$ the category of pullback Higgs sheaves of exponent $\leq m$, we obtain the following result, whose details we leave to the reader:
Lemma 3.2.
The above definition yields a functor $C^{-1}_{\tilde{Y}|X}:HIG_{p}(Y|X)\to MIC_{p}(X)$ and there is a natural isomorphism $C^{-1}_{\tilde{Y}|X}(V_{X},\theta_{Y|X})\cong f^{*}C^{-1}_{\tilde{Y}}(V_{Y},\theta_{Y})$.
Corollary 3.3.
For $(V_{Y},\theta_{Y})\in HIG_{m}(Y)$ as above, if $(G,0)\to(V_{X},\theta_{Y|X})$ is a morphism of restricted Higgs sheaves, then
$(F^{*}G,\nabla_{can})\cong C^{-1}_{\tilde{Y}|X}(G,0)\to f^{*}C_{\tilde{Y}}^{-1}(V_{Y},\theta_{Y})$
is a morphism of flat sheaves.
3.2.2 The case of nodal curves
Now take $\tilde{X}/W_{2}$ to be a flat nodal curve and $\tilde{\iota}:\tilde{X}\to\tilde{Y}$ a locally closed embedding over $W_{2}$, with base-change $\iota:X\to Y$ to $k$. The case of an embedding will be enough to deduce the case of a general $W_{2}$-liftable morphism (see the proof of Proposition 3.7). Let $\nu:X^{\prime}\to X$ be the normalization of the fiber over $k$, $\iota^{\prime}=\nu\circ\iota$, and let $D$ be the divisor of points of $X^{\prime}$ mapping to nodes of $X$, which we call prenodes. Let $Y_{i}$ be an affine open covering of $Y$. Set $X_{i}=X\cap Y_{i}$ and $X_{ij}=X_{i}\cap X_{j}$; likewise let $X^{\prime}_{i}=\nu^{-1}(X_{i})$ and $X^{\prime}_{ij}=\nu^{-1}(X_{ij})$. We can and will assume that no overlap $X_{ij}$ contains a node.
If $I$ is the ideal of $X$ in $Y$, $N_{X/Y}^{\vee}=I/I^{2}$ is the conormal bundle of $X$, which is locally free. We have a natural exact sequence
[…]
Let $\tilde{I}$ be the ideal of $\tilde{X}$ in $\tilde{Y}$. By lifting to $\tilde{I}$ we have morphisms
$t_{i}:=p^{-1}\tilde{\iota}^{*}\tilde{F}_{i}:F^{*}N_{X_{i}/Y_{i}}^{\vee}\to\mathcal{O}_{X_{i}}.$
since the map $p^{-1}\tilde{F}_{i}:\tilde{I}\to F_{*}\mathcal{O}_{X_{i}}$ kills $pI+\tilde{I}^{2}$. Note that $\rho_{Y|X}\circ\iota^{*}\zeta_{i}\circ F^{*}\delta=dt_{i}$ on $X_{i}$ and $\iota^{*}s_{ij}\circ F^{*}\delta=t_{i|j}-t_{j|i}$ on the overlap $X_{ij}$.
The pullback of (3.2.1) to $X^{\prime}$ fits into a commutative diagram with exact rows and columns
00{0}00{0}00{0}${Q}$${\nu^{*}\Omega_{X}}$${\Omega_{X^{\prime}}}$00{0}00{0}${N_{X^{\prime}/Y}^{\vee}}$${\iota^{\prime*}\Omega_{Y}}$${\Omega_{X^{\prime}}}$00{0}${\nu^{*}N_{X/Y}^{\vee}}$${\nu^{*}N_{X/Y}^{\vee}}$00{0}00{0}$\scriptstyle{\rho_{Y|X^{\prime}}}$$\scriptstyle{\nu^{*}\rho_{Y|X}}$$\scriptstyle{\nu^{*}\delta}$
The locally free sheaf $N_{X^{\prime}/Y}^{\vee}$ is the conormal bundle of the individual branches of $X^{\prime}$ in $Y$. Note that for any locally free sheaf $V$ on a curve and any subspace $U\subset k(x)\otimes V$ of the fiber at $x$, we may define the subsheaf $F$ of $V(x)$ of sections with simple poles in $U$ at $x$ by taking $F(-x)$ to be the kernel of the composition $V\to k(x)\otimes V\to(k(x)\otimes V)/U$. For any prenode $x_{1}$, there is a canonical subspace of the fiber $k(x_{1})\otimes N_{X^{\prime}/Y}$ given by the tangent direction of the curve at the other prenode $x_{2}$ which is attached to $x_{1}$ in $X$. We define $M_{X^{\prime}/Y}\subset N_{X^{\prime}/Y}(D)$ to be the subsheaf of sections of $N_{X^{\prime}/Y}$ with simple poles at each prenode $x_{1}$ in this subspace.
Lemma 3.4.
  1. 1.
    $Q$ has length one at each prenode. In particular, there is a natural inclusion $\nu^{*}N_{X/Y}\subset N_{X^{\prime}/Y}(D)$.
  2. 2.
    Via this inclusion, $M_{X^{\prime}/Y}=\nu^{*}N_{X/Y}$.
Proof.
We may prove the claim étale locally, and therefore assume we have coordinates $y_{1},\ldots,y_{n}$ of $Y$ and that $X$ has ideal $(y_{1}y_{2},y_{3},\ldots,y_{n})$. On the branch cut out by $(y_{2},y_{3},\ldots,y_{n})$, $N^{\vee}_{X^{\prime}/Y}$ is generated by $dy_{2},\ldots,dy_{n}$, while $\nu^{*}N^{\vee}_{X/Y}$ is generated by $y_{1}dy_{2},dy_{3},\ldots,dy_{n}$. This gives the first claim. Since $M_{X^{\prime}/Y}$ is generated by $y_{1}^{-1}\frac{\partial}{\partial y_{2}},\frac{\partial}{\partial y_{3}},\ldots,\frac{\partial}{\partial y_{n}}$, the second claim is established as well. ∎
Let $W_{X^{\prime}/Y}$ be the subsheaf of $(\iota^{\prime*}T_{Y})(D)$ of sections with simple poles at each prenode in the tangent direction of the attached branch. By the lemma, there is then a natural exact sequence
$0\to\nu^{*}N_{X/Y}^{\vee}\to W^{\vee}_{X^{\prime}/Y}\to\Omega_{X^{\prime}}\to 0$ (3.2.2)
Choose a splitting of (3.2.2) on each $X_{i}^{\prime}$ and use it to obtain $\hat{t}_{i}$ such that the following diagram commutes.
[…]
The lifts $\hat{t}_{i}$ yield $\eta_{i}:F^{*}\Omega_{X_{i}^{\prime}}\to\Omega_{X_{i}^{\prime}}$ such that $\eta_{i}\circ F^{*}\rho_{Y_{i}|X^{\prime}_{i}}=\rho_{Y_{i}|X^{\prime}_{i}}\circ\iota^{\prime*}\zeta_{i}-d\hat{t}_{i}$ as maps $F^{*}W_{X^{\prime}_{i}/Y_{i}}^{\vee}\to\Omega_{X_{i}^{\prime}}$, and moreover on overlaps we may write $u_{ij}\circ F^{*}\rho_{Y_{ij}|X^{\prime}_{ij}}=\iota^{\prime*}s_{ij}-(\hat{t}_{i|j}-\hat{t}_{j|i})$ for $u_{ij}:F^{*}\Omega_{X_{ij}}\to\mathcal{O}_{X_{ij}}$. We have $du_{ij}=\eta_{i|j}-\eta_{j|i}$.
Let $(V,\theta)\in HIG_{m}(Y|X^{\prime})$ be a pullback Higgs sheaf. Let $J$ be the ideal of the (reduced) set of prenodes of $X^{\prime}$ at which $\theta$ is nonzero in the tangent direction of the attached branch. We write $\hat{t}_{i}(F^{*}\theta)$ for the resulting composition
$F^{*}(J\otimes V|_{X_{i}^{\prime}})\xrightarrow{F^{*}\theta}F^{*}(V|_{X_{i}^{\prime}}\otimes W_{X_{i}^{\prime}/Y_{i}}^{\vee})\xrightarrow{\operatorname{id}\otimes\hat{t}_{i}}F^{*}V|_{X_{i}^{\prime}}$
and likewise for $\iota^{\prime*}s_{ij}(F^{*}\theta)$ and $u_{ij}\circ F^{*}\rho_{Y|X^{\prime}}(F^{*}\theta)$. We therefore have local morphisms
$\exp(-\hat{t}_{i}(F^{*}\theta)):F^{*}(J^{m-1}\otimes V)|_{X_{i}^{\prime}}\to V|_{X_{i}^{\prime}}$
and, because the overlaps contain no nodes, commutative diagrams on the overlaps
[…]
Denote by $(V,\theta_{X})$ the Higgs sheaf associated to $(V,\theta)$ and let $(G,0)\subset(V,\theta_{X})$ be a Higgs subsheaf with vanishing Higgs field. By restricting the above local endomorphisms, we obtain local maps
$\exp(-\hat{t}_{i}(F^{*}\theta)):F^{*}(J^{m-1}\otimes G)|_{X_{i}^{\prime}}\to F^{*}V|_{X_{i}^{\prime}}$ (3.2.3)
and commutative diagrams on the overlaps
[…]
again because there are no nodes in the overlaps.
Lemma 3.5.
The local maps (3.2.3) give morphisms of flat sheaves
$\exp(-\hat{t}_{i}(F^{*}\theta)):(F^{*}(J^{m-1}\otimes G)|_{X_{i}^{\prime}},\nabla_{can})\to(F^{*}V|_{X_{i}^{\prime}},\nabla_{can}+\rho_{Y_{i}|X^{\prime}_{i}}\circ\iota^{\prime*}\zeta_{i}(F^{*}\theta))$
and glue via (3.2.4) to a global morphism of flat sheaves $(F^{*}(J^{m-1}\otimes G),\nabla_{can})\to C_{\tilde{Y}|X^{\prime}}^{-1}(V,\theta)$. The morphism does not depend on the choice of splittings of (3.2.2).
Proof.
We first show that the local map (3.2.3) respects the connections. Let $s$ be a local section of $J^{m-1}\otimes G$, so that $\nabla_{can}F^{*}s=0$. We have
$\nabla_{can}\exp(-\hat{t}_{i}(F^{*}\theta))F^{*}s$ $=\exp(-\hat{t}_{i}(F^{*}\theta))(-d\hat{t}_{i}(F^{*}\theta))F^{*}s$
$=\exp(-\hat{t}_{i}(F^{*}\theta))(-\rho_{Y_{i}|X^{\prime}_{i}}\circ\iota^{\prime*}\zeta_{i}(F^{*}\theta))F^{*}s$
$=-\rho_{Y_{i}|X_{i}^{\prime}}\circ\iota^{\prime*}\zeta_{i}(F^{*}\theta)\exp(-\hat{t}_{i}(F^{*}\theta))F^{*}s$
which implies the compatibility of the connections. According to the description preceding Lemma 3.2, these maps glue as morphisms of flat sheaves since the vertical map of (3.2.4) is the gluing map for $C_{\tilde{Y}|X^{\prime}}^{-1}(V,\theta)$.
A different choice of splittings of (3.2.2) yields $\hat{t}_{i}^{\prime}$ for which $\hat{t}_{i}-\hat{t}_{i}^{\prime}=r_{i}\circ F^{*}\rho_{Y_{i}|X^{\prime}_{i}}$ for $r_{i}:F^{*}\Omega_{X_{i}^{\prime}}\to\mathcal{O}_{X_{i}^{\prime}}$. Since $r_{i}\circ F^{*}\rho_{Y_{i}|X^{\prime}_{i}}(F^{*}\theta)$ kills $F^{*}(J^{m-1}\otimes G)$, the maps in (3.2.3) do not depend on the choice of lift. ∎
Remark 3.6.
Its reasonable to ask whether there is a version of Lemma 3.5 for a Higgs subsheaf $(G,\psi)\subset(V,\theta_{X})$ for which the Higgs field is nonzero. The maps (3.2.3) glue via the maps $\exp(u_{ij}(F^{*}\psi))$, and the natural local connection which glues is $\nabla_{can}+\eta_{i}(F^{*}\psi)$, which must be interpreted as a connection with log poles. It is no longer clear that either the resulting flat sheaf with log poles or the morphism is independent of the choice of splittings of (3.2.2) (even up to isomorphism).
In summary, we have the following:
Proposition 3.7.
Let $\tilde{Y}/W_{2}$ be smooth, $\tilde{X}/W_{2}$ a nodal curve, $\tilde{f}:\tilde{X}\to\tilde{Y}$ a $W_{2}$-morphism, $\nu:X^{\prime}\to X$ the normalization of the fiber over $k$, and $f^{\prime}=f\circ\nu$. Let $(V,\theta)\in HIG_{m}(Y)$ with $m\leq p$, and let $J$ be the ideal of the (reduced) set of prenodes in $X^{\prime}$ at which $\theta$ is nonzero in the tangent direction of the attached branch. Then for any Higgs subsheaf $(G,0)\subset f^{\prime*}(V,\theta)$ there is a flat subsheaf $(F^{*}(J^{m-1}\otimes G),\nabla_{can})\subset f^{\prime*}C^{-1}_{\tilde{Y}}(V,\theta)$. The morphism of the underlying sheaves is locally isomorphic to the natural inclusion $F^{*}(J^{m-1}\otimes G)\subset F^{*}f^{\prime*}V$.
Proof.
By [OV], the inverse Cartier transform commutes with pullback under morphisms of smooth $k$-varieties which lift to $W_{2}$. Thus we may assume $f$ is an embedding $\iota$ by replacing $Y$ with $\mathbb{P}^{N}\times Y$. The claim then follows from Lemma 3.2 and Lemma 3.5. ∎
3.3 Applications to Fontaine–Laffaille modules
In this subsection we prove some degree bounds on Higgs subsheaves of the graded object of a Fontaine–Laffaille module pulled-back to a nodal curve (Propositions 3.9 and 3.10 below). Recall [OV, §4.6] that we may think of a (strict) $p$-torsion Fontaine–Laffaille module $\mathbf{M}=(M,\nabla,F^{\bullet},\Phi)$ on a lift $\tilde{Y}$ of a smooth variety $Y/k$ as a locally free flat sheaf $(M,\nabla)$ with a locally split filtration $F^{\bullet}$ satisfying Griffiths transversality and equipped with an isomorphism of flat sheaves
$\Phi:C^{-1}_{\tilde{Y}}\operatorname{gr}_{F}\mathbf{M}\xrightarrow{\cong}(M,\nabla).$
where $\operatorname{gr}_{F}\mathbf{M}:=(\operatorname{gr}_{F}M,\operatorname{gr}_{F}\nabla)$ is the associated Higgs bundle. The category of $p$-torsion Fontaine–Laffaille modules $\mathbf{M}$ whose graded object $\operatorname{gr}_{F}M$ has nonzero pieces in degrees $[a,w+a]$ is denoted $\mathrm{MF}^{\nabla}_{[a,w+a],1}(\tilde{Y})$. Note that for $p$-torsion Fontaine–Laffaille modules, the index of the Hodge filtration may be freely shifted, as there is no integrality to force the degrees to be nonnegative.
We give the setup for these results now for convenience.
Setup 3.8.
Let $\tilde{Y}/W_{2}$ be a smooth $W_{2}$-scheme equipped with a strict $p$-torsion Fontaine–Laffaille module $\mathbf{M}=(M,\nabla,F^{\bullet}M,\Phi)\in\mathrm{MF}^{\nabla}_{[a,w+a],1}(\tilde{Y})$. Assume $p$ is an odd prime with $p-1\geq w$.
Let $\tilde{X}/W_{2}$ be a nodal curve together with a $W_{2}$-morphism $\tilde{f}:\tilde{X}\to\tilde{Y}$. Let $\nu_{B}:B\to X$ be the normalization of a proper component of the special fiber $X$ and let $g:B\to Y$ be the composition $g=f\circ\nu_{B}$. Let $D_{B,\mathbf{M}}$ the divisor of prenodes on $B$ at which the Higgs field $\operatorname{gr}_{F}\nabla$ is nonzero in the tangent direction of the attached branch, and $d_{B,\mathbf{M}}=\deg D_{B,\mathbf{M}}$ its degree.
Proposition 3.9.
Assume Setup 3.8. Suppose $(V,\theta)\subset g^{*}\operatorname{gr}_{F}\mathbf{M}$ is a Higgs subsheaf. Then $\mu(V)\leq wd_{B,\mathbf{M}}$ provided $p-1>wd_{B,\mathbf{M}}\operatorname{rk}(V)$.
Proof.
We argue as in the proof of Ogus–Vologodsky [OV]. Note that the exponent of $\operatorname{gr}_{F}\mathbf{M}$ is $m\leq w+1$. We first reduce to the case $(V,\theta)=(L,0)$ with rank one by taking the $\operatorname{rk}(V)$-th wedge power. We must then show that any Higgs sheaf $(L,0)\subset g^{*}\operatorname{gr}_{F}\mathbf{M}$ with $\operatorname{rk}(L)=1$ and vanishing Higgs field has $\deg(L)\leq wd_{B,\mathbf{M}}$. Suppose this is not the case, and let $(L,0)$ be a rank one Higgs subsheaf with vanishing Higgs field of maximal degree, $\deg(L)>wd_{B,\mathbf{M}}$. Let $J_{B,\mathbf{M}}=\mathcal{O}_{B}(-D_{B,\mathbf{M}})$. By Proposition 3.7 we have that
$(F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L),\nabla_{can})\to g^{*}C^{-1}_{\tilde{Y}}\operatorname{gr}_{F}\mathbf{M}\xrightarrow{g^{*}\Phi}g^{*}(M,\nabla)$
is an injective morphism of flat sheaves. Let $i\geq 0$ be minimal such that
$F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L)\to g^{*}M/F^{i+1}g^{*}M$
is nonzero. We therefore have an injective map
$F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L)\to\operatorname{gr}^{i}_{F}g^{*}M.$ (3.3.1)
Since $(F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L),\nabla_{can})$ is a flat subsheaf and the Higgs field of $\operatorname{gr}_{F}g^{*}M$ is $\operatorname{gr}_{F}\nabla$, (3.3.1) yields an injective morphism of Higgs sheaves
$(F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L),0)\to g^{*}\operatorname{gr}_{F}\mathbf{M}.$
But
$\deg(F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L))$ $=p(\deg(L)-(m-1)d_{B,\mathbf{M}})$
$\geq p(\deg(L)-wd_{B,\mathbf{M}})$
$\geq\deg(L)-wd_{B,\mathbf{M}}+(p-1)$
$>\deg(L)$
because of the assumption on $p$, and this contradiction finishes the proof. ∎
Let $(V,\theta)$ be a Higgs bundle on a smooth variety $Y/k$. There is a natural endomorphism Higgs bundle $(\operatorname{\mathsf{End}}(V),\operatorname{ad}\theta)$. The Kodaira–Spencer map of $(V,\theta)$ is the Higgs field of $(\operatorname{\mathsf{End}}(V),\operatorname{ad}\theta)$, thought of as a map $\operatorname{ad}\theta:T_{Y}\to\operatorname{\mathsf{End}}(V)$. For $\mathbf{M}\in\operatorname{MF}^{\nabla}_{[a,w+a],1}(Y)$, there is also a natural endomorphism object $\operatorname{\mathsf{End}}(\mathbf{M})=(\operatorname{\mathsf{End}}(M),\operatorname{ad}\nabla,F^{\bullet}\operatorname{\mathsf{End}}(M),\operatorname{ad}\Phi)\in\operatorname{MF}^{\nabla}_{[-w,w],1}(Y)$ in the category of Fontaine–Laffaille modules. The Kodaira–Spencer map of $\mathbf{M}$ is the Kodaira–Spencer map of the associated Higgs sheaf $\operatorname{gr}_{F}\mathbf{M}$.
Proposition 3.10.
Assume Setup 3.8 and let $P=\operatorname{img}(g^{*}\operatorname{gr}_{F}\operatorname{ad}\nabla)$ be the image of the pullback of the Kodaira–Spencer map $g^{*}\operatorname{gr}_{F}\operatorname{ad}\nabla:g^{*}T_{Y}\to g^{*}\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{M})$. Then $P$ does not contain a positive degree subsheaf of rank $r$ provided $p>2rwd_{B,\wedge^{r}\operatorname{\mathsf{End}}(\mathbf{M})}$.
Proof.
For simplicity of notation, set $(V,\theta)=\operatorname{gr}_{F}\mathbf{M}$. Observe that the Kodaira–Spencer map $\operatorname{ad}\theta:(T_{Y},0)\to(\operatorname{\mathsf{End}}(V),\operatorname{ad}\theta)$ is a morphism of Higgs sheaves, as therefore is $\wedge^{r}\operatorname{ad}\theta:(\wedge^{r}T_{Y},0)\to(\wedge^{r}\operatorname{\mathsf{End}}(V),\wedge^{r}\operatorname{ad}\theta)$. Thus, we have morphisms of flat sheaves
$(F^{*}\wedge^{r}T_{Y},\nabla_{can})=C_{\tilde{Y}}^{-1}(\wedge^{r}T_{Y},0)\to C_{\tilde{Y}}^{-1}(\wedge^{r}\operatorname{\mathsf{End}}(V),\wedge^{r}\operatorname{ad}\theta)\xrightarrow{\Phi}(\wedge^{r}\operatorname{\mathsf{End}}(M),\wedge^{r}\operatorname{ad}\nabla)$
and so morphisms of flat sheaves
$g^{*}(F^{*}\wedge^{r}T_{Y},\nabla_{can})\to g^{*}C_{\tilde{Y}}^{-1}\operatorname{gr}_{F}\wedge^{r}\operatorname{\mathsf{End}}(\mathbf{M})\xrightarrow{g^{*}\Phi}g^{*}(\wedge^{r}\operatorname{\mathsf{End}}(M),\wedge^{r}\operatorname{ad}\nabla).$
Observe that the image of the first map is naturally identified with $(F^{*}\wedge^{r}P,\nabla_{\mathrm{can}})$. Indeed, its enough to check this on the level of sheaves. The first map is locally identified with $g^{*}F^{*}\wedge^{r}\operatorname{ad}\theta$, while $(F^{*}\wedge^{r}P,\nabla_{\mathrm{can}})$ is the image of $F^{*}\wedge^{r}g^{*}\operatorname{ad}\theta$, since $F^{*},\wedge^{r}$ preserve injections and surjections of locally free sheaves. The functors $F^{*},g^{*},\wedge^{r}$ all naturally commute, so the claim follows.
If $G\subset P$ is subsheaf with $\operatorname{rk}(G)=r$, then setting $L=\wedge^{r}G$ by the same argument as in the previous proof we get an injective morphism of Higgs sheaves
$(F^{*}L,0)\to g^{*}\operatorname{gr}_{F}\wedge^{r}\operatorname{\mathsf{End}}(\mathbf{M})$
but then by Proposition 3.9 we would have
$2rwd_{B,\wedge^{r}\operatorname{\mathsf{End}}(\mathbf{M})}\geq\deg(F^{*}L)=p\deg(L)$
which given the assumption on $p$ means we must have $\deg(L)=\deg(G)\leq 0$.
Remark 3.11.
The same proof shows for instance that in Setup 3.8, $g^{*}\operatorname{gr}^{0}_{F}M$ does not contain a positive degree subsheaf of rank $r$ provided $p>rwd_{B,\wedge^{r}\mathbf{M}}$.
Proposition 3.10 will allow us to control the liftability of rational components of nodal curves with boundedly many nodes. This will be the key ingredient for the surface case of Proposition 4.4. The normal case requires a bit more care, so we only treat the smooth case now, and postpone the general setup until then.
Definition 3.12.
Given a Higgs bundle $(V,\theta)$ on $Y$, we say the Kodaira–Spencer map is quasi-finite if there is a partition of $Y$ into locally closed smooth subschemes $Y^{i}$ on which the Kodaira–Spencer map $T_{Y^{i}}\to\operatorname{\mathsf{End}}(V)|_{Y^{i}}$ is immersive, meaning it is injective with locally-free cokernel.
Of course, if the Kodaira–Spencer map of $(V,\theta)$ is already immersive, then $(V,\theta)$ has quasi-finite Kodaira–Spencer map, and this will be the relevant case for Shimura varieties.
Corollary 3.13.
Assume Setup 3.8, and further assume that the Kodaira–Spencer map of $\operatorname{gr}_{F}(M,\nabla)$ is quasifinite. Then if $B$ is rational, $g:B\to Y$ is constant provided $p>2wd_{B,\operatorname{\mathsf{End}}(\mathbf{M})}$
Proof.
[…]
Corollary 3.14.
Assume Setup 3.8, and further assume that the Kodaira–Spencer map of $\operatorname{gr}_{F}(M,\nabla)$ is quasifinite. Finally, assume either
  1. 1.
    $\tilde{X}$ is obtained by successively blowing up a smooth curve $\tilde{X}^{\prime}/W_{2}$ at closed points and $p>2w$.
  2. 2.
    $\tilde{X}$ is obtained by successively blowing up a nodal curve $\tilde{X}^{\prime}/W_{2}$ at closed points and $p>4w$.
Then $X\to Y$ factors through $X\to X^{\prime}$.
Proof.
The dual graph of the exceptional curves is a tree and the components that are not contracted form a subtree. The extremal components have $d_{B,\mathbf{M}}\leq 1$ in the first case and $\leq 2$ in the second, so the claim follows form the previous corollary. ∎
4 Criteria for local integral canonical models
The goal of this section is to give a general setting in which local integral canonical models exists, in the sense of Definition 2.14. For simplicity, we first give the smooth case.
Theorem 4.1.
Let $F$ be a mixed characteristic local field with ring of integers $\mathcal{O}$ and residue field $k$ of characteristic $p$. Let $\mathcal{S}/\mathcal{O}$ be a smooth $\mathcal{O}$-scheme equipped with:
  1. 1.
    A log smooth compactification $\bar{\mathcal{S}}/\mathcal{O}$ of $\mathcal{S}$ with natural boundary.
  2. 2.
    A strict $p$-torsion Fontaine-Laffaille module $\mathbf{M}\in\mathrm{MF}^{\nabla}_{[a,w+a],1}(\mathcal{S})$ whose Higgs bundle $\operatorname{gr}_{F}\mathbf{M}$ has quasifinite Kodaira–Spencer map.
Then if $p>2w$, $\mathcal{S}$ is an integral canonical model.
See Definition 2.2 and Definition 3.12 for the definitions of natural boundary and quasifinite Kodaira–Spencer map. In the normal case, we will additionally need a form of stratified resolution.
Definition 4.2.
Let $\mathcal{O}$ be a Dedekind domain and $\mathcal{Y}/\mathcal{O}$ a normal finite type flat $\mathcal{O}$-scheme. A stratified resolution of $\mathcal{Y}$ consists of:
  1. 1.
    An increasing sequence of closed subspaces $\mathcal{Y}^{1}\subset\ldots\subset\mathcal{Y}^{m}=\mathcal{Y}$, each flat over $\mathcal{O}$, such that $\mathcal{Y}^{j}\setminus\mathcal{Y}^{j-1}$ is smooth over $\mathcal{O}$ for each $j$. Let $\mathcal{Y}^{\prime j}\to\mathcal{Y}^{j}$ be the normalization.
  2. 2.
    For each $j$, a resolution $\pi^{\prime j}:\mathcal{S}^{j}\to\mathcal{Y}^{\prime j}$ and a log smooth compactification $\bar{\mathcal{S}}^{j}/\mathcal{O}$ with the property that the composition $\mathcal{S}^{j}\to\mathcal{Y}^{\prime j}\to\mathcal{Y}^{j}$ (which we call $\pi^{j}:\mathcal{S}^{j}\to\mathcal{Y}^{j}$) is an isomorphism over $\mathcal{Y}^{j}\setminus\mathcal{Y}^{j-1}$.
We say the stratified resolution:
  1. 3.
    is uniform if the resolution of the top stratum $\pi^{m}:\mathcal{S}^{m}\to\mathcal{Y}^{m}=\mathcal{Y}$ is uniform.
  2. 4.
    has natural boundary if each $\bar{\mathcal{S}}^{j}$ has natural boundary.
Theorem 4.3.
Let $F$ be a mixed characteristic local field with ring of integers $\mathcal{O}$ and residue field $k$ of characteristic $p$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme equipped with:
  1. 1.
    A uniform stratified resolution $\{\pi^{j}:\mathcal{S}^{j}\to\mathcal{Y}^{j}\}$ with natural boundary.
  2. 2.
    A Higgs bundle $(V,\theta)$ on $\mathcal{Y}_{0}$ with quasifinite Kodaira–Spencer map.
  3. 3.
    For each $j$ a strict $p$-torsion Fontaine-Laffaille module $\mathbf{M}^{j}\in\mathrm{MF}^{\nabla}_{[a^{j},w^{j}+a^{j}],1}(\mathcal{S}^{j})$ and a nonzero torsion-free coherent sheaf $G^{j}$ on $\mathcal{S}^{j}$ such that
    $(G^{j},0)\otimes\operatorname{gr}_{F}\mathbf{M}^{j}$ $\cong(G^{j},0)\otimes\pi^{j*}(V,\theta)$
    $(G^{j},0)\otimes\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{M}^{j})$ $\cong(G^{j},0)\otimes\pi^{j*}(\operatorname{\mathsf{End}}(V),\operatorname{ad}\theta)$
    as Higgs sheaves.
Then if $p>2\max(w^{j})$, $\mathcal{Y}$ is an integral canonical model.
Note that Theorem 4.3 directly implies Theorem 4.1 by taking the trivial stratification, $(V,\theta)=\operatorname{gr}_{F}\mathbf{M}$, and $(G^{m},0)=(\mathcal{O}_{\mathcal{S}_{0}},0)$.
4.1 Extension in codimension $\geq 2$
We now use the results of Section 3.3 to show that, in the presence of a Fontaine–Laffaille module with quasi-finite Kodaira–Spencer map, morphisms from smooth sources which are defined outside of a set of codimension $\geq 2$ extend.
Proposition 4.4.
Assume the setup of Theorem 4.3. Let $\mathcal{X}/\mathcal{O}$ be a smooth $\mathcal{O}$-scheme and $f:\mathcal{X}\setminus\mathcal{Z}\to\mathcal{Y}$ a morphism for $\mathcal{Z}\subset\mathcal{X}$ a subset of codimension $\geq 2$ which is contained in the special fiber. Then there is an extension $f:\mathcal{X}\to\mathcal{Y}$.
Proof.
If $j$ is minimal such that $Y^{j}$ contains the image of the generic fiber $X$, then the image of $X$ meets $Y^{j}\setminus Y^{j-1}$, so there is a dense open set of $X$ which lifts to $S^{j}$. By purity of the branch locus, $\mathcal{X}\setminus\mathcal{Z}\to\mathcal{X}$ induces an isomorphism on fundamental groups, so the map $X\to Y$ is $\mathcal{O}$-admissible. It follows by Lemma 2.3 that there is a finite type flat $\mathcal{X}^{\prime}/\mathcal{O}$, a proper morphism $\mathcal{X}^{\prime}\to\mathcal{X}$ which is an isomorphism on a dense open subset, and a morphism $\mathcal{X}^{\prime}\to\mathcal{S}^{j}$ whose generic fiber lifts the morphism $X\to Y^{j}$.
There is nothing to prove for $\dim\mathcal{X}\leq 1$. For $\dim\mathcal{X}\geq 2$, we argue as in [dejongoort] (see also [moretbailly, faltingschai]) by induction on $\dim\mathcal{X}$ starting with the base case $\dim\mathcal{X}=2$ and using deformation theory for the induction step.
Assume first that $\dim\mathcal{X}=2$. By resolution of singularities for surfaces ([lipman], specifically [stacks-project, Tag 0AHI]), there is a composition of blow-ups at closed points $\mathcal{X}_{n}\to\cdots\to\mathcal{X}_{1}\to\mathcal{X}$ whose composition factors as $\mathcal{X}_{n}\to\mathcal{X}^{\prime}\to\mathcal{X}$. To show that $\mathcal{X}^{\prime}\to\mathcal{S}^{j}\to\mathcal{Y}$ factors through $\mathcal{X}^{\prime}\to\mathcal{X}$ (and hence the existence of the required extension), it suffices (for instance by Lemma 2.16) to show the following:
Claim.
The composition $\mathcal{X}_{n}\to\mathcal{X}^{\prime}\to\mathcal{S}^{j}\to\mathcal{Y}$ contracts all exceptional curves.
Proof.
Note that by the quasifiniteness assumption on the Kodaira–Spencer map of $(V,\theta)$, a component $B$ of the special fiber of $\mathcal{X}_{n}$ is contracted if and only if the Higgs field of $(V,\theta)$ vanishes on the image of $B$ in $\mathcal{Y}_{0}$. Moreover, for any such $B$, the Higgs field of $\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{M}^{j})$ also vanishes on $B$, by assumption (3). The exceptional curves form a tree and the components $B$ which are not contracted form a sub-tree whose extremal components have $d_{B,\operatorname{\mathsf{End}}(\mathbf{M}^{j})}\leq 1$, so by Proposition 3.9 it suffices to show that $g^{*}\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{M}^{j})$ contains a sub-Higgs sheaf of slope larger than $w^{j}$, where $g:B\to\mathcal{S}^{j}_{0}$ is the map on the special fiber restricted to $B$.
[…]
$(G,0)\otimes g^{*}\operatorname{\mathsf{End}}(\mathbf{M}^{j})\cong(G,0)\otimes h^{*}(\operatorname{\mathsf{End}}(V),\operatorname{ad}\theta)\cong(G,0)\otimes(\alpha\circ\gamma)^{*}\operatorname{\mathsf{End}}(\mathbf{M}^{i})$
where $G=g^{*}G^{j}\otimes(\gamma\circ\alpha)^{*}G^{i}$. Thus, $(G,0)\otimes\alpha^{*}(F^{*}T_{B^{\prime}},0)$ is a Higgs subsheaf of $(G,0)\otimes g^{*}\operatorname{\mathsf{End}}(\mathbf{M}^{j})$. Since $B$ is rational, $G$ splits [^3] Possibly after a field extension. as a sum of factors $\mathcal{O}_{B}(d)$, and taking the largest degree factor we find that $((\alpha^{*}F^{*}T_{B^{\prime}})(e),0)$ is a Higgs subsheaf of $g^{*}\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{M}^{j})$ for some $e\geq 0$. Since $\deg(\alpha^{*}F^{*}T_{B^{\prime}})(e)\geq 2p$, the conclusion follows. ∎
We now treat the general case $\dim\mathcal{X}>2$ by induction via deformation theory. It will be sufficient to show that for any $x\in\mathcal{Z}$, $f$ extends on an étale neighborhood of $x$. In particular we may assume $\mathcal{X}$ is affine and $\mathcal{Y}$ is a closed subscheme of a smooth affine $\mathcal{O}$-scheme $\mathcal{V}$. Let $\mathcal{H}\subset\mathcal{X}$ be a divisor containing $x$ such that $\mathcal{H}\cap\mathcal{Z}$ is codimension $\geq 2$ in $\mathcal{H}$. By induction, the map $\mathcal{H}\setminus\mathcal{Z}\to\mathcal{Y}$ extends to $f_{\mathcal{H}}:\mathcal{H}\to\mathcal{Y}$. Let $g_{\mathcal{H}}:\mathcal{H}\to\mathcal{V}$ be the composition with the embedding $\mathcal{Y}\subset\mathcal{V}$. Let $\mathcal{H}_{i}$ be the $i$th order thickening of $\mathcal{H}$ in $\mathcal{X}$. Given a lift $\mathcal{H}_{i}\to\mathcal{Y}$ of $f_{\mathcal{H}}:\mathcal{H}\to\mathcal{Y}$, lifts to a morphism $\mathcal{H}_{i+1}\to\mathcal{V}$ are unobstructed since the obstruction lies in $H^{1}(\mathcal{H},\operatorname{\mathcal{O}}_{\mathcal{H}}(-i\mathcal{H})\otimes g_{\mathcal{H}}^{*}T_{\mathcal{V}})=0$; the set of such lifts is naturally a torsor under $H^{0}(\mathcal{H}_{i},\operatorname{\mathcal{O}}_{\mathcal{H}}(-i\mathcal{H})\otimes g_{\mathcal{H}}^{*}T_{\mathcal{V}})$. Likewise, lifts of $\mathcal{H}_{i}\setminus\mathcal{Z}\to\mathcal{Y}$ to $\mathcal{H}_{i+1}\setminus\mathcal{Z}\to\mathcal{V}$ are a torsor under $H^{0}(\mathcal{H}\setminus\mathcal{Z},\operatorname{\mathcal{O}}_{\mathcal{H}}(-i\mathcal{H})\otimes g_{\mathcal{H}\setminus\mathcal{Z}}^{*}T_{\mathcal{S}_{v}})\cong H^{0}(\mathcal{H},\operatorname{\mathcal{O}}_{\mathcal{X}}(-i\mathcal{H})\otimes g^{*}_{\mathcal{H}}T_{\mathcal{V}})$, since $\operatorname{\mathcal{O}}_{\mathcal{H}}(-i\mathcal{H})\otimes g^{*}_{\mathcal{H}}T_{\mathcal{V}}$ is locally free and $\operatorname{codim}_{\mathcal{H}}\mathcal{Z}\geq 2$. In particular, there is a lift $\mathcal{H}_{i+1}\to\mathcal{V}$ which agrees with the given map $\mathcal{H}_{i+1}\setminus\mathcal{Z}\to\mathcal{Y}\to\mathcal{V}$. Since the ideal of $\mathcal{H}_{i}$ in $\mathcal{H}_{i+1}$ is locally free (over $\mathcal{H}$), this lift factors through a morphism $\mathcal{H}_{i+1}\to\mathcal{Y}$. Letting $\mathcal{X}^{\wedge}_{\mathcal{H}}$ denote the completion of $\mathcal{X}$ along $\mathcal{H}$, it follows that there is a formal map $\mathcal{X}^{\wedge}_{\mathcal{H}}\to\mathcal{Y}$ whose restriction to $(\mathcal{X}\setminus\mathcal{Z})_{\mathcal{H}\setminus\mathcal{Z}}^{\wedge}$ agrees with the completion of $f:\mathcal{X}\setminus\mathcal{Z}\to\mathcal{Y}$ along $\mathcal{H}\setminus\mathcal{Z}$.
Let $\mathcal{W}$ be the closure of the graph of $\mathcal{X}\setminus\mathcal{Z}\to\mathcal{Y}$ in $\mathcal{X}\times\mathcal{Y}$; note that since $\mathcal{X}^{\prime}\to\mathcal{X}$ is proper, it follows that $\mathcal{W}\to\mathcal{X}$ is as well. Since the ideal sheaf of $\mathcal{H}_{i}$ in $\mathcal{H}_{i+1}$ is locally free (over $\mathcal{H}$) for each $i$, the natural formal map $\mathcal{X}_{\mathcal{H}}^{\wedge}\to\mathcal{X}\times\mathcal{Y}$ factors through $\mathcal{W}$ as it does so over $(\mathcal{X}\setminus\mathcal{Z})_{\mathcal{H}\setminus\mathcal{Z}}^{\wedge}$. Thus, the projection $\mathcal{W}\to\mathcal{X}$ admits a section over $\mathcal{X}_{\mathcal{H}}^{\wedge}$. Say the image of $w$ is $x$; then the map $\operatorname{Spec}(\operatorname{\mathcal{O}}_{\mathcal{W},w})\to\operatorname{Spec}(\operatorname{\mathcal{O}}_{\mathcal{X},x})$ admits a formal section. Since $\mathcal{W}\to\mathcal{X}$ is proper and birational, the following lemma implies it is an isomorphism above $x$.
Lemma 4.5.
Let $q:\mathcal{W}\to\mathcal{X}$ be a proper birational morphism of finite-type integral $\mathcal{O}$-schemes, $w$ a closed point of $\mathcal{W}$ and $x=q(w)$. If the induced map $\operatorname{Spec}(\operatorname{\mathcal{O}}_{\mathcal{W},w})\to\operatorname{Spec}(\operatorname{\mathcal{O}}_{\mathcal{X},x})$ admits a formal section then it is an isomorphism at $w$.
Proof.
The claim is étale-local on $\mathcal{X}$. By Artin approximation [artinapprox, Corollary (2.5)] (see also [stacks-project, Tag 0CAU]), after replacing $\mathcal{X}$ with an étale cover there is an algebraic map $\mathcal{X}\to\mathcal{W}$ agreeing with the formal section to first order. But the composition $\mathcal{X}\to\mathcal{W}\to\mathcal{X}$ agrees with the identity to first order, and therefore is an isomorphism. We may assume it is in fact the identity, in which case $\mathcal{W}\to\mathcal{X}$ admits an algebraic section. It follows that the short exact sequence of $\operatorname{\mathcal{O}}_{\mathcal{X},x}$-modules
$0\to\operatorname{\mathcal{O}}_{\mathcal{X},x}\to\operatorname{\mathcal{O}}_{\mathcal{W},w}\to\operatorname{\mathcal{O}}_{\mathcal{W},w}/\operatorname{\mathcal{O}}_{\mathcal{X},x}\to 0$
splits. Since $q$ is birational, $\operatorname{\mathcal{O}}_{\mathcal{W},w}/\operatorname{\mathcal{O}}_{\mathcal{X},x}$ is torsion, hence must be zero since $\operatorname{\mathcal{O}}_{\mathcal{W},w}$ is a domain. ∎
Proof of Theorem 4.3.
By assumption, $\mathcal{Y}$ admits a uniform resolution. Given smooth $\mathcal{X}/\mathcal{O}$ and an $\mathcal{O}$-admissible $f_{F}:X\to Y$, by Corollary 2.4 $f_{F}$ extends in codimension 1 and by Proposition 4.4 the rest of the way.
Remark 4.6.
[…]
4.2 Local models of images of period maps
In this section we give a slightly different formulation of Theorem 4.3 which deduces the existence of the $\mathbf{M}^{j}$ at each level of the resolution from a single global $\mathbf{M}$.
Proposition 4.7.
Let $F$ be a mixed characteristic local field with ring of integers $\mathcal{O}$ and residue field $k$ of characteristic $p$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme equipped with:
  1. 1.
    A uniform stratified resolution $\{\pi^{j}:\mathcal{S}^{j}\to\mathcal{Y}^{j}\}$ with natural boundary.
  2. 2.
    A smooth $\mathcal{P}/\mathcal{O}$ with a proper dominant $\phi:\mathcal{P}\to\mathcal{Y}$ and a Fontaine-Laffaille module $\mathbf{M}\in\mathrm{MF}^{\nabla}_{[a,w+a]}(\mathcal{P})$ with mod $p$ reduction $\mathbf{M}_{0}$ such that:
    1. (a)
      The filtered flat vector bundle $(M_{0},\nabla,F^{\bullet}M_{0})$ underlying $\mathbf{M}_{0}$ is pulled back from $(V,\nabla,F^{\bullet}V)$ on $\mathcal{Y}_{0}$.
    2. (b)
      The Higgs bundle $(\operatorname{gr}_{F}V,\operatorname{gr}_{F}\nabla)$ has quasifinite Kodaira–Spencer map.
    3. (c)
      The local system ${}_{et}\mathbb{V}_{p}$ on $P$ corresponding to $\mathbf{M}$ under the Fontaine–Laffaille correspondence descends to $Y$.
  3. 3.
    We finally suppose for each $j$ there is a smooth $\mathcal{T}^{j}/\mathcal{O}$ and a commutative diagram
    ${\mathcal{T}^{j}}$${\mathcal{P}}$${\mathcal{S}^{j}}$${\mathcal{Y}^{\prime j}}$${\mathcal{Y}}$$\scriptstyle{q^{j}}$$\scriptstyle{t^{j}}$$\scriptstyle{h}$$\scriptstyle{\pi^{j}}$
    where $q^{j}$ is proper dominant and étale over a dense open subset of $\mathcal{S}^{j}$.
Then if $p>2w$, $\mathcal{Y}$ is an integral canonical model.
Before the proof we need the following:
Lemma 4.8.
Let $F$ be a mixed characteristic local field with ring of integers $\mathcal{O}$ and residue field $k$ of characteristic $p$. Let $\mathcal{S},\mathcal{T}/\mathcal{O}$ be smooth.
  1. 1.
    If $f:\mathcal{T}\to\mathcal{S}$ is finite étale of prime to $p$ degree and ${}_{et}\mathbb{V}_{p}$ is a $\mathbb{Z}_{p}$-local system on $S$ such that $f^{*}_{et}\mathbb{V}_{p}$ is crystalline, then ${}_{et}\mathbb{V}_{p}$ is crystalline, and the inclusion ${}_{et}\mathbb{V}_{p}\to f_{*}f^{*}(_{et}\mathbb{V}_{p})$ respects the lift of Frobenius.
  2. 2.
    Suppose given:
    1. (a)
      $f:\mathcal{T}\to\mathcal{S}$ a proper dominant morphism, with base-change $f_{0}:\mathcal{T}_{0}\to\mathcal{S}_{0}$ to the special fiber;
    2. (b)
      $\mathcal{U}_{0}\subset\mathcal{S}_{0}$ a dense open set with preimage $\mathcal{V}_{0}=f_{0}^{-1}(\mathcal{U}_{0})$;
    3. (c)
      strict $p$-torsion Fontaine–Laffaille modules $\mathbf{M}=(M,\nabla,F^{\bullet}M,\Phi_{\mathbf{M}})$ on $\mathcal{T}$ and $\mathbf{N}_{\mathcal{U}}=(N_{\mathcal{U}},\nabla,F^{\bullet}N_{\mathcal{U}},\Phi_{\mathbf{N}_{\mathcal{U}}})$ on $\mathcal{U}$ as well as an extension $(N,\nabla,F^{\bullet}N)$ of the underlying filtered flat vector bundle of $\mathbf{N}_{\mathcal{U}}$ to $\mathcal{S}_{0}$;
    4. (d)
      compatible isomorphisms $f^{*}\mathbf{N}_{\mathcal{U}}\xrightarrow{\cong}\mathbf{M}|_{\mathcal{V}}$ and $f_{0}^{*}(N,\nabla,F^{\bullet}N)\xrightarrow{\cong}(M,\nabla,F^{\bullet}M)$.
    Then there is a $p$-torsion Fontaine–Laffaille module $\mathbf{N}$ extending $\mathbf{N}_{\mathcal{U}}$ and an isomorphism $f_{0}^{*}\mathbf{N}\to\mathbf{M}$ compatible with the above isomorphisms.
Proof.
The category of Fontaine–Laffaille modules is a full subcategory of the category of local systems, so the descent datum of $f^{*}_{et}\mathbb{V}_{p}$ also descends the corresponding Fontaine–Laffaille modules [faltings]. Since ${}_{et}\mathbb{V}_{p}\to f_{*}f^{*}(_{et}\mathbb{V}_{p})$ is a direct factor of local systems, it follows it comes from a direct factor of Fontaine–Laffaille modules as well.
We now prove (2). Observe that the $p$-curvature of $(N,\nabla)$ is nilpotent of the same index as $\mathbf{N}$ (and $\mathbf{M}$). Unwrapping the definitions, we have isomorphisms
$\Phi_{\mathbf{N}_{\mathcal{U}}}:C_{\mathcal{U}}^{-1}\operatorname{gr}_{F}\mathbf{N}|_{\mathcal{U}_{0}}\xrightarrow{\cong}(N,\nabla)|_{\mathcal{U}_{0}}$
$\Phi_{\mathbf{M}}:C_{\mathcal{T}}^{-1}\operatorname{gr}_{F}\mathbf{M}\xrightarrow{\cong}(M,\nabla)$
such that $f_{0}^{*}\Phi_{\mathbf{N}_{\mathcal{U}}}=\Phi_{\mathbf{M}}$ over $\mathcal{V}_{0}$, using the natural identification $f^{*}_{0}C^{-1}_{\mathcal{S}}\operatorname{gr}_{F}\mathbf{N}\cong C^{-1}_{\mathcal{T}}f_{0}^{*}\operatorname{gr}_{F}\mathbf{N}\cong C^{-1}_{\mathcal{T}}\operatorname{gr}_{F}\mathbf{M}$. Thus, we have solid commutative diagrams
[…]
Since $\mathcal{T},\mathcal{S}$ are smooth and $f$ is proper dominant (and therefore the same for $f_{0}$), $\mathcal{O}_{\mathcal{S}_{0}}\to f_{0*}\mathcal{O}_{\mathcal{T}_{0}}$ has torsion-free cokernel. As $N$ and $C^{-1}_{\mathcal{S}}\operatorname{gr}_{F}\mathbf{N}$ are locally free, the vertical arrows in the left square of (4.2.1) have torsion-free cokernel as well, and it follows that $f_{0*}\Phi_{\mathbf{M}}$ restricts to an extension $\Phi_{\mathbf{N}}:C^{-1}_{\mathcal{S}}\operatorname{gr}_{F}\mathbf{N}\xrightarrow{\cong}(N,\nabla)$ of $\Phi_{\mathbf{N}_{\mathcal{U}}}$. ∎
Proof of Proposition 4.7.
By assumption, for each $j$, we have a Fontaine–Laffaille module $t^{j*}\mathbf{M}$ on $\mathcal{T}^{j}$ whose corresponding local system is pulled back from $\mathcal{Y}$. Moreover, the underlying filtered flat vector bundle of $t^{j*}\mathbf{M}_{0}$ is pulled back from $(V,\nabla,F^{\bullet}V)$ on $\mathcal{Y}_{0}$. Applying part 1 of Lemma 4.8 to the open set $\mathcal{U}^{j}$ of $\mathcal{S}^{j}$ over which $q^{j}$ is finite étale, we conclude that there is a Fontaine–Laffaille module $\mathbf{N}_{\mathcal{U}^{j}}$ which pulls back to $(t^{j*}\mathbf{M})|_{(q^{j})^{-1}(\mathcal{U}^{j})}$. Applying part 2, we find that $(\mathbf{N}_{\mathcal{U}^{j}})_{0}$ extends to a Fontaine–Laffaille module $\mathbf{N}^{j}_{0}$ which pulls back to $t^{j*}\mathbf{M}_{0}$. In particular, we have $q^{j*}\operatorname{gr}_{F}\mathbf{N}^{j}_{0}\cong q^{j*}\pi^{j*}(\operatorname{gr}_{F}V,\operatorname{gr}_{F}\nabla)$, so
$(q^{j}_{*}\mathcal{O}_{\mathcal{T}^{j}_{0}},0)\otimes\operatorname{gr}_{F}\mathbf{N}^{j}_{0}\cong(q^{j}_{*}\mathcal{O}_{\mathcal{T}_{0}^{j}},0)\otimes\pi^{j*}(\operatorname{gr}_{F}V,\operatorname{gr}_{F}\nabla)$
as Higgs sheaves. By the same argument, we have
$(q^{j}_{*}\mathcal{O}_{\mathcal{T}^{j}_{0}},0)\otimes\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{N}^{j})\cong(q^{j}_{*}\mathcal{O}_{\mathcal{T}_{0}^{j}},0)\otimes\pi^{j*}(\operatorname{gr}_{F}\operatorname{\mathsf{End}}(V),\operatorname{gr}_{F}\operatorname{ad}\nabla)$
as Higgs sheaves. Thus, all the conditions of Theorem 4.3 are satisfied. ∎
4.3 Making period maps proper
We will need a result allowing us to extend geometric Fontaine–Laffaille modules when the underlying filtered flat vector bundle extends.
Proposition 4.9.
Let $F$ be a mixed charactersitic local field with ring of integers $\mathcal{O}$ and residue field $k$. Let $\mathcal{P}/\mathcal{O}$ be smooth with a partial log smooth compactification $(\mathcal{P}^{\prime},\mathcal{D})/\mathcal{O}$. Let $f:\mathcal{Z}\to\mathcal{P}$ be a smooth projective morphism with geometrically connected fibers of relative dimension $<\textrm{char}(k)$. Let $\mathbf{M}=H^{k}_{DR}(\mathcal{Z}/\mathcal{P})$ be the $p$-torsion Fontaine–Laffaille module with underlying filtered flat vector bundle the De Rham cohomology of the special fiber $(M,\nabla,F^{\bullet}M)=H^{k}_{DR}(\mathcal{Z}_{0}/\mathcal{P}_{0})$. Assume:
  1. 1.
    There is a log smooth curve (possibly not irreducible) $(\mathcal{C}^{\prime},\Delta)/\mathcal{O}$ and a morphism $g^{\prime}:(\mathcal{C}^{\prime},\Delta)\to(\mathcal{P}^{\prime},\mathcal{D})$ with $\Delta=(\mathcal{C}^{\prime}\times_{\mathcal{P}^{\prime}}\mathcal{D})^{\mathrm{red}}$ such that $g^{\prime}(\Delta)$ meets every component of $\mathcal{D}$ and the base-change $\mathcal{Z}_{\mathcal{C}}$ (where $\mathcal{C}=\mathcal{C}^{\prime}\setminus\Delta$) has a log smooth compactification $(\mathcal{Z}^{\prime}_{\mathcal{C}},\mathcal{E})/\mathcal{O}$ and a morphism $f^{\prime}:\mathcal{Z}^{\prime}_{\mathcal{C}}\to\mathcal{C}^{\prime}$ extending the natural one.
  2. 2.
    The filtered flat vector bundle $(M,\nabla,F^{\bullet}M)$ extends to $(M^{\prime},\nabla^{\prime},F^{\bullet}M^{\prime})$ on $\mathcal{P}^{\prime}$.
  3. 3.
    $g^{\prime*}(M^{\prime},\nabla^{\prime},F^{\bullet}M^{\prime})\cong H^{k}_{DR}((\mathcal{Z}^{\prime}_{\mathcal{C},0},\mathcal{E}_{0})/(\mathcal{C}_{0}^{\prime},\Delta_{0}))$ via an isomorphism extending the natural one, where $H^{k}_{DR}((\mathcal{Z}^{\prime}_{\mathcal{C},0},\mathcal{E}_{0})/(\mathcal{C}_{0}^{\prime},\Delta_{0}))$ is the relative log De Rham cohomology.
Then $(M^{\prime},\nabla^{\prime},F^{\bullet}M^{\prime})$ underlies an extension $\mathbf{M}^{\prime}$ of $\mathbf{M}$ to $\mathcal{P}^{\prime}$ as a Fontaine–Laffaille module.
Proof.
The main step is the following straightforward claim. Throughout by “$\tilde{\mathcal{C}}$” we mean the base-change to $W_{2}(k)$.
Claim.
There is an isomorphism
$F^{*}\operatorname{gr}_{F}\mathbb{H}^{k}(\Omega^{\bullet}_{\mathcal{Z}^{\prime}_{\mathcal{C},0}/\mathcal{C}_{0}^{\prime}}(\log\mathcal{E}_{0}))\xrightarrow{\Phi}\mathbb{H}^{k}(\Omega^{\bullet}_{\mathcal{Z}_{\mathcal{C},0}/\mathcal{C}_{0}^{\prime}}(\log\mathcal{E}_{0}))$
extending the usual Frobenius morphism of Deligne–Illusie over $\mathcal{C}_{0}$.
Proof.
By log smoothness, both $\tilde{\mathcal{C}}^{\prime}$ and $\tilde{\mathcal{Z}}^{\prime}_{\mathcal{C}}$tale locally [^4] In fact, the lifts of Frobenius preserving the boundary exist Zariski locally. To simplify notation, if $(\tilde{X},\tilde{D})/W_{2}$ is log smooth and $(X,D)$ the reduction mod $p$, then there is a natural exact sequence $0\to F^{*}T_{X}(-\log D)\to F^{*}T_{X}\to\bigoplus_{i}F^{*}\mathcal{O}_{D_{i}}(D_{i})\to 0$ where the $D_{i}$ are the components of $D$ with defining equations $f_{i}$, and the last map sends $1\otimes\delta$ to $1\otimes\frac{\delta f_{i}}{f_{i}}\mod f_{i}$ on the $i$th factor. Lifts of Frobenius to $\tilde{X}$ are a torsor for $H^{0}(F^{*}T_{X})$, so if $\tilde{F}$ is a local lift of Frobenius to $\tilde{X}$, and $\tilde{f}_{i}$ local defining equations for $\tilde{D}_{i}$, then $\tilde{F}(\tilde{f}_{i})=\tilde{f}^{p}_{i}+pg_{i}$, and Zariski locally we may modify $\tilde{F}$ so that each $\tilde{F}(\tilde{f}_{i})$ is divisible by $\tilde{f}_{i}$. admit compatible lifts of Frobenius fixing each component of the boundary. Thus, as in [deligneillusie], we obtain a morphism of filtered complexes
$(\Omega_{\mathcal{Z}^{\prime}_{\mathcal{C},0}/\mathcal{C}_{0}^{\prime}}^{\bullet}(\log\mathcal{E}_{0}),0)\to(F_{*}\Omega_{\mathcal{Z}^{\prime}_{\mathcal{C},0}/\mathcal{C}_{0}^{\prime}}^{\bullet}(\log\mathcal{E}_{0}),F_{*}d)$
which restricts to the usual quasi-isomorphism over $\mathcal{C}_{0}$ and is easily checked to be a quasi-isomorphism. ∎
To prove the proposition, we must show the Frobenius $\Phi:C^{-1}_{\mathcal{P}}\operatorname{gr}_{F}\mathbf{M}\to\mathbf{M}$ extends to an isomorphism $\Phi^{\prime}:C^{-1}_{\mathcal{P}^{\prime}}(\operatorname{gr}_{F}M^{\prime},\operatorname{gr}_{F}\nabla^{\prime})\to(M^{\prime},\nabla^{\prime})$. Since both sides are locally free, it suffices to show that $\Phi$ extends in codimension one, which may therefore be checked on the pullback $g^{\prime*}(M^{\prime},\nabla^{\prime},F^{\bullet}M^{\prime})$. This is the claim. ∎
5 Existence of integral canonical models
In this section, we deduce the existence of integral canonical models for sufficiently large primes.
5.1 Proof of Theorem 1.3
In this section we show that models of Shimura varieties satisfy the conditions of Theorem 4.1 for large primes, thereby proving Theorem 1.3. Let $S=S_{K}(G,X)$ be a Shimura variety. We let $E$ denote the reflex field. $S_{K}(G,X)$ carries a flat principal $G$-bundle $\mathcal{E}/S$, which admits a canonical model over $E$. A choice of hodge co-character defines a parabolic subgroup $P\subset G$, and gives a parabolic bundle $\mathcal{P}/S$, along with a canonical morphism $\mathcal{P}\rightarrow\mathcal{E}$, where all this data descends to $E_{G}$. We note that $\mathcal{P}$ is not flat for the connection on $\mathcal{E}$.
Let $V$ denote a faithful rational representation of $G$, and $\mathbb{V}\subset V$ denote a lattice. The choice of $\mathbb{V}$ specifies a compact open subgroup of $G(\mathbb{A}_{f})$ by considering the stabilizers of $\mathbb{V}_{\hat{\mathbb{Z}}}$. We choose the level structure $K$ to be a subgroup of this compact open subgroup, such that it is neat, and such that it acts trivially on $\mathbb{V}/3\mathbb{V}$. Corresponding to $V$ there is a family of $\mathbb{Q}_{\ell}$ local systems ${}_{{\mathrm{et}}}V_{\ell}$ with $\mathbb{Z}_{\ell}$ sub-systems ${}_{{\mathrm{et}}}\mathbb{V}_{\ell}$. Moreover, there is a filtered flat vector bundle ${}_{{}_{\mathrm{dR}}}V$ over $S_{K}(G,X)$ – indeed we have that ${}_{{}_{\mathrm{dR}}}V$ is defined by ×$\mathcal{E}\times_{G}V$. The filtration is induced by the fact that ${}_{{}_{\mathrm{dR}}}V$ is also canonically isomorphic to $\mathcal{P}\times_{G_{E}}V_{E}$. Now, let $p$ be a prime and let $v\mid p$ be a prime of $E$. By [DLLZ], there is a canonical isomorphism between $(_{{}_{\mathrm{dR}}}V)_{E,v}$ and ${}_{p-_{\mathrm{dR}}}V$, the filtered flat bundle obtained by applying the $p$-adic Riemann-Hilbert correspondence to ${}_{{\mathrm{et}}}V_{p}|_{S_{K}(G,X)_{E,v}}$.
By choosing $N$ to be a large enough integer, we may assume that $S_{K}(G,X)$ spreads out to a smooth integral model $\mathscr{S}_{K}(G,X)/\mathcal{O}_{E}[1/N]$ such that:
  • $G$ has a reductive model $\mathcal{G}/\mathbb{Z}[1/N]$. There are elements $s_{\alpha}\in\mathbb{V}[1/N]^{\otimes}$ whose point-wise stabilizers define the group $\mathcal{G}$.
  • For every $p\nmid N$, we have $K=\mathcal{G}(\mathbb{Z}_{p})\cdot K^{p}$, where $K^{p}$ is the prime-to-$p$ level structure.
  • $P\subset G_{E}$ is induced by a parabolic subgroup of $\mathcal{G}$ defined over $\mathcal{O}_{E}[1/N]$.
  • The map $\mathcal{P}\rightarrow\mathcal{E}$ is induced by a map $\mathscr{P}\rightarrow\mathscr{E}$ over $\mathcal{S}$, where $\mathscr{E}$ is a flat principal $\mathcal{G}$ bundle and $\mathscr{P}$ is a principal bundle for the parabolic of $\mathcal{G}$ defined over $\mathcal{O}_{E}[1/N]$.
Note that this gives us an integral model ${}_{{}_{\mathrm{dR}}}\mathbb{V}/\mathscr{S}_{K}(G,X)$ for the filtered flat bundle ${}_{{}_{\mathrm{dR}}}V$ (where the flat bundle is defined as ×$\mathscr{E}\times_{\mathcal{G}}\mathbb{V}[\frac{1}{N}]$, and the filtration is induced by $\mathscr{P}\rightarrow\mathscr{E}$).
Lemma 5.1.
After increasing $N$ we have the following.
  • $\mathscr{S}_{K}(G,X)$ admits a log-smooth compactification over $\mathcal{O}_{E}[1/N]$.
  • The local systems ${}_{{\mathrm{et}}}\mathbb{V}_{\ell}$ extend to $\mathscr{S}_{K}(G,X)[1/\ell]$.
  • ${}_{{\mathrm{et}}}\mathbb{V}_{p}$ is crystalline on $\mathscr{S}_{K}(G,X)_{E_{v}}$ for every finite place $v$ of $\mathcal{O}_{E}[1/N]$.
  • For each finite place $v$ of $\mathcal{O}_{E}[1/N]$, let ${}_{{}_{\mathrm{cris}}}\mathbb{V}/\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ denote the Fontaine-Laffaille module corresponding to ${}_{{\mathrm{et}}}\mathbb{V}_{p}$ under the Fontaine-Laffaille correspondence. Then the Kodaira–Spencer map of the Higgs bundle $\operatorname{gr}_{F}(_{{}_{\mathrm{dR}}}\mathbb{V})$ is immersive.
Proof.
The first condition follows by choosing a log smooth compactification over the generic fiber.
For the second condition, we use the following result of [KP, Thm 1.3]:
Theorem 5.2 (Klevdal-Patrikis).
There is an integral model $\mathcal{S}$ of $S_{K}(G,X)$ over $\mathbb{Z}[1/N]$ for some $N$ such that for each $\ell$ the local systems ${}_{{\mathrm{et}}}\mathbb{V}_{\ell}$ all extend to $\mathcal{S}[1/\ell]$.
The third condition is proven by Esnault-Groechenig [EG] for large enough primes, and this in particular gives an integral structure to the filtered flat bundle ${}_{p-_{\mathrm{dR}}}V$. Under the canonical isomorphism ${}_{p-_{\mathrm{dR}}}V\rightarrow_{{}_{\mathrm{dR}}}V|_{S_{K}(G,X)_{E_{v}}}$, the integral structures induced by ${}_{{}_{\mathrm{cris}}}\mathbb{V}$ and ${}_{{}_{\mathrm{dR}}}\mathbb{V}$ are the same by [PST].
Finally, using this isomorphism, the final condition follows because the same is true for the Higgs bundle of ${}_{{}_{\mathrm{dR}}}V$ over the generic point. ∎
It follows from Theorem 4.1 that $\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ is an integral canonical model. It remains to argue that $\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ is the unique model admitting a log smooth compactification. Let $\mathcal{S}$ be another such model. The extension property yields that the identity map over the generic fiber extends to $f:\mathcal{S}\to\mathscr{S}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$, and by purity of the branch locus this map is an open immersion. The inverse image of the boundary of $\bar{\mathcal{S}}$ in the closure $\mathcal{W}\subset\bar{\mathcal{S}}\times\bar{\mathscr{S}}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$ of the graph of $f$ must map to the boundary of $\bar{\mathscr{S}}_{K}(G,X)_{\mathcal{O}_{E_{v}}}$, by the smooth extension property. Thus, the extension is an isomorphism. ∎
5.2 Proof of Theorem 1.9
Suppose $F$ is a number field, $P/F$ a smooth variety, and $f:Z\to P$ a smooth projective family. Let $f:P^{\prime}\to Y$ be the Stein factorization of the (compactification) of the period map associated to $\mathbb{W}_{\mathbb{Z}}:=Rf_{*}\mathbb{Z}_{Z^{\mathrm{an}}}$ as described in 2.7. Then $f$ is defined over $F$. By spreading out over $\mathcal{O}:=\mathcal{O}_{F}[1/n]$ and inverting finitely many primes, we may assume we have smooth $\mathcal{P},\mathcal{Z}/\mathcal{O}$, a log smooth partial compactification $(\mathcal{P}^{\prime},\mathcal{D})/\mathcal{O}$ of $\mathcal{P}$, and $\mathcal{Y}/\mathcal{O}$ normal finite type flat as well as:
  1. 1.
    $f:\mathcal{Z}\to\mathcal{P}$ smooth projective
  2. 2.
    $\phi:\mathcal{P}^{\prime}\to\mathcal{Y}$ proper with $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}\mathcal{O}_{\mathcal{P}^{\prime}}$.
  3. 3.
    $\mathcal{Y}$ admits a uniform stratified resolution $\{\pi^{j}:\mathcal{S}^{j}\to\mathcal{Y}^{j}\}$.
  4. 4.
    For each $j$ there is a smooth $\mathcal{T}^{j}/\mathcal{O}$ and a commutative diagram
    ${\mathcal{T}^{j}}$${\mathcal{P}^{\prime}}$${\mathcal{S}^{j}}$${\mathcal{Y}^{\prime j}}$${\mathcal{Y}}$$\scriptstyle{q^{j}}$$\scriptstyle{t^{j}}$$\scriptstyle{\bar{\phi}}$$\scriptstyle{\pi^{j}}$
    where $q^{j}$ is proper dominant and étale over a dense open subset of $\mathcal{S}^{j}$.
  5. 5.
    There is a log smooth curve (possibly not irreducible) $(\mathcal{C}^{\prime},\Delta)/\mathcal{O}$ and a morphism $g^{\prime}:(\mathcal{C}^{\prime},\Delta)\to(\mathcal{P}^{\prime},\mathcal{D})$ with $\Delta=(\mathcal{C}^{\prime}\times_{\mathcal{P}^{\prime}}\mathcal{D})^{\mathrm{red}}$ such that $g^{\prime}(\Delta)$ meets every component of $\mathcal{D}$ and the base-change $\mathcal{Z}_{\mathcal{C}}$ (where $\mathcal{C}=\mathcal{C}^{\prime}\setminus\mathcal{E}$) has a log smooth compactification $(\mathcal{Z}^{\prime}_{\mathcal{C}},\mathcal{E})/\mathcal{O}$ and a morphism $f^{\prime}:\mathcal{Z}^{\prime}_{\mathcal{C}}\to\mathcal{C}^{\prime}$ extending the natural one.
  6. 6.
    The filtered flat vector bundle $(V,\nabla,F^{\bullet}V)$ on $Y$ extends to $(\mathcal{V},\nabla,F^{\bullet}\mathcal{V})$ on $\mathcal{Y}$, and its pullback to $\mathcal{P}$ is identified with the filtered flat vector bundle underlying $H^{k}_{DR}(\mathcal{Z}/\mathcal{P})$. Moreover, the reduction $(\mathcal{V},\nabla,F^{\bullet}\mathcal{V})_{v}$ has quasifinite Kodaira–Spencer map for every finite place $v$.
  7. 7.
    By Corollary 2.23, the local system $H^{k}_{et}(\mathcal{Z}/\mathcal{P}[1/\ell],\mathbb{Z}_{\ell})$ extends to $\mathcal{P}^{\prime}$ and descends to ${}_{et}\mathbb{V}_{\ell}$ on $\mathcal{Y}[1/\ell]$. By Lemma 2.24 and Lemma 2.7, it follows that the stratified resolution has natural boundary.
This yields part 1(a), and both 2 and 3 follow from 1(b) by Corollary 2.18 and Proposition 2.15 (1(c) is clear from construction). Thus it remains to show the extension property in 1(b), and since by Lemma 2.24 we have the extension in characterstic 0, it suffices to show $\mathcal{Y}_{v}$ is an integral canonical model for each finite place $v$. By Theorem 4.3 we must show for each finite place $v$ of $\mathcal{O}$ that there is a Fontaine–Laffaille module $\mathbf{M}$ on $\mathcal{P}^{\prime}_{v}$ with the required properties. The filtered flat vector bundle underlying the $p$-torsion module of $H^{k}_{DR}(\mathcal{Z}_{v}/\mathcal{P}_{v})$ extends to $\mathcal{P}^{\prime}_{v}$, and so by Proposition 4.9 it extends as a Fontaine–Laffaille module whose underlying filtered flat vector bundle is the pullback of $(\mathcal{V},\nabla,F^{\bullet}\mathcal{V})_{\mathcal{O}_{v}}$ on $\mathcal{Y}_{v}$.∎
6 Application to unramified Borel-extension
In this section, we will apply the extension property for Shimura varieties and images of geometric period maps. To that end, let $F$ be a number field, and let $(Y,\overline{Y},\mathbb{V})$ denote one of the following:
  1. 1.
    $Y\subset\overline{Y}$ is a Shimura variety embedded inside a log smooth toroidal compactification, and $\mathbb{V}$ is ${}_{{\mathrm{et}}}\mathbb{V}_{\ell}$ associated to the adjoint representation of the underlying group $G$ as in Section 5.1. We pick $\ell=3$.
  2. 2.
    $Y\subset\overline{Y}$ is as in Section 5.2. Recall that there is a smooth proper family over a smooth variety $P$, and $Y$ is the image of the period morphism from $P$. For every prime $\ell$, there is an etale $\ell$-adic local system on $P$ induced by the smooth proper family, and this $\ell$-adic local system descends to an $\ell$-adic local system $\mathbb{V}$ on $Y$. As in the Shimura case, we will choose $\ell=3$.
In the non-Shimura setting, we assume the existence of stratified resolution of $\mathcal{Y}/\mathcal{O}_{F}[1/N]$ as in Definition 4.2 where the resolution is also defined over $\mathcal{O}_{F}[1/N]$ where $N$ is some large enough integer. We suppose that $p\nmid N$ is a large enough prime. In particular, we suppose that $p$ is large enough such that the main theorems of Section 5 hold, and such that $p\nmid\#\operatorname{GL}_{n}(\mathbb{Z}/3\mathbb{Z})$, where $n$ is the rank of $\mathbb{V}$.
Definition 6.1.
Let $K$ denote a finite extension of $W(\overline{\mathbb{F}}_{p})[1/p]$ or $\mathbb{F}_{p}((t))$. We say that a point $x\in Y(K)$ has good reduction if its specialization lies in the interior $\mathcal{Y}(\overline{\mathbb{F}}_{p})$. We say that $x$ has bad reduction otherwise.
The content of the following lemma is that a point having good reduction or bad reduction can be tested by considering the fiber of $\mathbb{V}$ at that point.
Lemma 6.2.
Let $K$ be as in Definition 6.1. Let $x\in Y(K)$. Then $x$ has bad reduction if and only if the inertial action of $I_{x}\subset\operatorname{Gal}(\overline{K}/K)$ on $\mathbb{V}_{x}$ is quasi-unipotent of infinite order.
Proof.
By arguments similar to those in [PST] (which we may use, as $\mathbb{V}$ has prime-to-$p$ monodromy given our choice of $p$), we have that a point $x\in Y(K)$ has bad reduction if and only if the inertial action on the fiber at $x$ is quasi-unipotent of infinite order. Indeed, in the Shimura case, this is just [PST, Theorem 4.4] (the result in loc. cit. is for finite extensions of $W(\overline{\operatorname{\mathbb{F}}}_{p})$, but the identical argument goes through in the case of finite extensions of $\overline{\operatorname{\mathbb{F}}}_{p}((t))$). In the geometric case, at the cost of replacing $K$ by a finite extension, there is a point $y\in P(K)$ lifting $x$. The same argument as in [PST, Theorem 4.4] yields that $y$ specializes to the boundary of $P$ if and only if the inertial action is trivial. The claim follows for $x$ as $y$ specializes to the boundary of $Y$ if and only $x$ specializes to the boundary of $P$. ∎
Let $\mathsf{D}$ denote the rigid-analytic closed unit disc (defined over $\mathbb{Q}_{p}$) and let $\mathsf{D}^{\times}=\mathsf{D}\setminus\{0\}$. Recall that a thin annulus is a rigid analytic annulus whose outer and inner radii are the same, and a thick annulus is one whose inner and outer radii are different. We are now ready to state the main theorem of this section – this is a one-dimensional analogue of Borel’s extension theorem in the unramified setting (see the main result of [padicBorel]).
Theorem 6.3.
Let $Y$ be as above, and let $f:\mathsf{D}^{\times}\rightarrow Y^{{\mathrm{an}}}$ denote a rigid-analytic map defined over $W[1/p]$, where $W=W(\overline{\mathbb{F}}_{p})$. Suppose that the image of $f$ contains a point with good reduction. Then, $f$ extends to a rigid-analytic map $\mathsf{D}\rightarrow Y^{{\mathrm{an}}}$.
We first need the following lemma:
Lemma 6.4.
Let $f:\mathsf{D}^{\times}\rightarrow Y^{{\mathrm{an}}}$ be a rigid-analytic map defined over a finite extension of $W[1/p]$, such that the image of $f$ contains a point with good reduction. Then the image of $f$ is entirely contained in the good reduction locus.
Proof.
Note that the local system $f^{-1}\mathbb{V}$ has prime-to-$p$ monodromy. The proof of this lemma is essentially the same as [padicBorel, Theorem 3.3]. As in loc. cit., it suffices to prove the lemma for maps from closed annuli. Proposition 3.6 of loc. cit. proves that an $\ell$-adic local system on a thin annulus must be unramified everywhere if it is at a single point.
To deduce the case of thick annuli from thin annuli, we proceed as in [padicBorel, Section 3.1.2]. The argument applies verbatim to reduce to proving that any map $f:\mathbb{G}_{m}\rightarrow\bar{\mathcal{Y}}_{\mathbb{F}_{p}}$ extends to a map from $\mathbb{P}^{1}$. Note that the local system $\mathbb{V}/\mathcal{Y}_{\mathbb{F}_{p}}$ is point-wise pure. Indeed, in the Shimura case, $\mathbb{V}$ is a direct sum of irreducible local systems with trivial determinant (this follows from the fact that the local system is induced by the adjoint representation of $G$). In the case of geometric period images, it suffices to prove point-wise-purity with $P$ in place of $Y$. But this is true as $\mathbb{V}$ is induced by a geometric family over $P$.
Suppose now that $x\in\mathbb{P}^{1}\setminus\mathbb{G}_{m}$ maps to the boundary. We have that $f^{-1}(\mathbb{V})$ has quasi-unipotent local monodromy of infinite order. We have that $f$ descends to some finite field. In both the cases, the local system originally arises from an irreducible local system over a smooth base, and is therefore point-wise pure over that pure base. As this is a point-by-point condition, it is also point-wise pure over $Y$, and therefore over when pulled back to $\mathbb{G}_{m}$ by $f$. By [DeligneWeil2, Corollaire 3.4.12], the local system has semi-simple geometric monodromy over $\mathbb{G}_{m}$. As the prime-to-$p$ geometric fundamental group of $\mathbb{G}_{m}$ is abelian, the local system pulled back to $\mathbb{G}_{m}$ must have toric geometric monodromy. Therefore, the local monodromy around the boundary of $\mathbb{G}_{m}$ is not quasi-unipotent with infinite order in the first place, and the result follows. ∎
We are now ready to prove Theorem 6.3.
Proof.
By the lemma above, we may assume that the image of $\mathsf{D}^{\times}$ is contained in the good reduction locus of $Y^{{\mathrm{an}}}$. Let $Y^{j}\subset Y$ denote the smallest stratum that contains the image of $\mathsf{D}^{\times}$. We claim that the map $f:\mathsf{D}^{\times}\rightarrow{Y^{j}}^{{\mathrm{an}}}$ lifts to a map $g:\mathsf{D}^{\times}\rightarrow{S^{j}}^{{\mathrm{an}}}$. Indeed, $\mathsf{D}^{\times}\cap{Y^{j-1}}^{{\mathrm{an}}}$ is a analytic closed subset, and so consists of a sequence of points $Q_{i}\in\mathsf{D}^{\times}$ with the property that only finitely many of the $Q_{i}$ lie in any closed sub-annulus of $\mathsf{D}^{\times}$. Pick $D_{i}$ to be a small closed disc centered around $Q_{i}$ such that $D_{i}\cap Y^{j-1}=Q_{i}$. It suffices to show that the map $D_{i}\rightarrow{Y^{j}}^{{\mathrm{an}}}$ lifts to $D_{i}\rightarrow S^{j}$. But this follows from the fact that $S^{j}\rightarrow Y^{j}$ is proper.
Let $\mathsf{A}_{n}$ denote the rigid-analytic annulus with outer radius 1 and inner radius $\frac{1}{p^{n}}$, and let $g_{n}=g|_{\mathsf{A}_{n}}$. Therefore, the map $f_{n}$ is induced by a map of formal schemes $\mathscr{X}_{n}\rightarrow\widehat{\mathcal{S}^{j}}^{p}$, where $\mathscr{X}_{n}$ is a formal scheme with rigid generic fiber $\mathsf{A}_{n}$, and $\widehat{\mathcal{S}^{j}}^{p}$ is the $p$-adic completion of $\mathcal{S}^{j}$. By [padicBorel, Appendix], we may assume that the special fiber of $\mathscr{X}_{n}$ is a union of projective lines and two affine lines, and that the the irreducible components form a tree. As $\mathscr{X}_{n}\mod p^{2}$ is a scheme that maps algebraically to $\mathscr{S}_{K}(G,X)$ over $W/p^{2}$, the Fontaine-Laffaille mod module mod $p$ pulls back to the special fiber of $\mathscr{X}_{n}$. In the Shimura case, Corollary 3.14 yields that every every $\mathbb{P}^{1}$ contracts to a point. In the case of the image of period maps, an argument similar to the claim in the proof of Proposition 4.4 applies to show that the map $\pi^{j}\circ g$ contracts every $\mathbb{P}^{1}$.
Therefore we may assume that the special fiber of $\mathscr{X}_{n}$ is just $\overline{\mathbb{F}}_{p}[x,y]/(xy)$. Further, we have that the tube over the node is the open annulus $\mathsf{A}_{n}^{\circ}\subset\mathsf{A}_{n}$. Therefore, we have that $f(\mathsf{A}_{n}^{\circ})$ is contained in the residue ball over some fixed $z\in\mathcal{Y}^{j}(\overline{F}_{p})$. It follows that the image of the open punctured unit disk $\mathsf{D}^{\circ,\times}\subset\mathsf{D}^{\times}$ is contained in the residue ball over $z\in\mathcal{Y}^{j}(\overline{\mathbb{F}}_{p})$, and thus is bounded. We now apply the rigid-analytic Riemann extension theorem to conclude that $f$ extends to a map $\mathsf{D}\rightarrow Y^{{\mathrm{an}}}$.
7 Tate-semisimplicity and special points
The aim of this section is to first prove that prime-to-$p$ Hecke correspondences extend to the integral model over $\mathcal{O}_{E,v}$ where $v\mid p$ and use this to prove that the Frobenius endomorphism on ${}_{\mathrm{et}}\mathbb{V}_{\ell,x}$ is semisimple for every $x\in\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})$. The main theorem of this section is:
Theorem 7.1.
Let $p\not\mid N$ be a prime, and $x\in\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})$ where $k$ is a finite field of char. $p$. Then ${}_{{\mathrm{et}}}\varphi_{\ell,x}$ acts semisimply on $V_{\ell,x}$.
This is a special case of a folklore conjecture that posits that the Frobenii at closed points are semisimple for any irreducible local system defined over a smooth $\mathbb{F}_{q}$-variety. We will then use this to show that every $\overline{\mathbb{F}}_{p}$-point of $\mathscr{S}_{K}(G,X)$ is special. As the proof of Theorem 7.1 strongly uses Hecke correspondences, we will start this section with a discussion on Hecke correspondences in characteristic zero and their extension to $\mathscr{S}_{K}(G,X)$. We will then reformulate the notion of special points in terms of the Hecke action, and then prove Theorem 7.1 and use that to deduce that every $\overline{\mathbb{F}}_{p}$-point of $\mathscr{S}_{K}(G,X)$ is special.
7.0.1 Extension of Hecke correspondences
Let $\mathbb{V}$ be a lattice in a faithful representation $V$ of $G$, such that $\mathbb{V}\otimes_{\mathbb{Z}}\hat{Z}$ stabilized by $K$. Let $K(n)$ denote the kernel $K\rightarrow\operatorname{Aut}(\mathbb{V}/n\mathbb{V})$, and consider the congruence covers $S_{n}:=S_{K(\ell^{n})}(G,X)$. We focus on the towers $S_{\ell^{n}}$ for a fixed prime $\ell$. Over the generic fiber this is a (possibly disconnected) Galois cover of $S=S_{1}$. By taking the total space of the local systems $\mathbb{V}/\ell^{n}\mathbb{V}$ we obtain smooth integral models $\mathcal{S}_{n}$ with log-smooth compactifications over the base, that are Galois over $\mathscr{S}_{K}(G,X)=\mathcal{S}_{1}$.
We recall some basics about Hecke correspondences. Let $h\in G(\mathbb{Z}[1/\ell])$. Let $\ell^{n}$ be such that $h^{-1}K_{\ell}(\ell^{n})h\subset K_{\ell}$. Then there are well defined maps $\pi_{h}:S_{\ell^{n}}\rightarrow S_{1}$ for large enough $n$. Let $\pi:S_{\ell^{n}}\rightarrow S_{1}$ denote the standard projection.
Then $\pi^{-1}x$ is corresponds to (a subset of) mod $\ell^{n}$-equivalence classes of trivializations of $t:\mathbb{V}_{x}\cong\mathbb{V}$, and $\pi_{h}(x_{1})=\pi_{h}(x_{2})$ for $\pi(x_{1})=\pi(x_{2})=x$ if and only if $t_{x_{2}}t_{x_{1}}^{-1}\in K_{\ell}$.
Theorem 7.2.
For $h\in G(\mathbb{Q}_{\ell})$ the map $\pi_{h}:S_{\ell^{n}}\rightarrow S$ which exists for large enough $n$ extends as an etale map over $\mathbb{Z}[1/\ell N]$
Proof.
This is an immediate consequence of Theorem 1.3.
With this in hand, we have:
Definition 7.3.
The $\ell$-power Hecke correspondence $\tau_{h}$ associated to $h$ on $\mathscr{S}_{K}(G,X)/\mathcal{O}_{E}[1/\ell N]$ is defined as $\tau_{h}(x)=\pi_{h}(\pi^{-1}(x))$.
Let $p\nmid\ell N$ be a prime, let $F$ be a $p$-adic local field with $\mathcal{O}_{F}\subset F$ its ring of integers. Let $y\in\mathscr{S}_{K}(G,X)(\mathcal{O}_{F})$. We then have that $\tau_{h}(y)=\{y_{1}\ldots y_{n}\}$ for some integers $n$, where $y_{i}\in\mathscr{S}_{K}(G,X)(\mathcal{O}_{F_{i}})$ where $F_{i}/F$ are finite unramified extensions. There is a canonical Galois action of $\pi_{1}(\mathcal{O}_{F})$ on the set $\{y_{1}\ldots y_{n}\}$ induced by the action of $\pi_{1}(\mathcal{O}_{F})$ on ${}_{{\mathrm{et}}}\mathbb{V}_{\ell,x}$. Further, we have that the field of definition $F_{i}$ of $y_{i}$ is just the extension corresponding to the stabilizer of $y_{i}$ in $\pi_{1}(\mathcal{O}_{F})$. Finally, the Galois action of $\pi_{1}(\mathcal{O}_{F_{i}})$ on ${}_{{\mathrm{et}}}\mathbb{V}_{\ell,y_{i}}$ is induced by the Galois action of $\pi_{1}(\mathcal{O}_{F})$ on ${}_{{\mathrm{et}}}V_{\ell,x}=_{{\mathrm{et}}}\mathbb{V}_{\ell,x}\otimes\mathbb{Q}_{\ell}$ by considering ${}_{{\mathrm{et}}}\mathbb{V}_{\ell,y_{i}}={t_{i}^{-1}}h_{{\mathrm{et}}}\mathbb{V}_{\ell}\subset_{{\mathrm{et}}}V_{\ell,x}$ (here, $t_{i}$ is the choice of mod $\ell^{n}$-trivialiazation corresponding to the point $y_{i}$). This description also holds with $\mathcal{O}_{F}$ replaced with finite fields, as $\pi_{1}(\mathcal{O}_{F})=\pi_{1}(\mathbb{F}_{q})$ where $\mathbb{F}_{q}$ is the residue field of $\mathcal{O}_{F}$.
This discussion has the following immediate consequence.
Lemma 7.4.
Let $y\in\mathscr{S}_{K}(G,X)(\mathcal{O}_{F})$ (resp. $\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})$). Let $\sigma\in\pi_{1}(\mathcal{O}_{F})$ (resp. $\pi_{1}(\mathbb{F}_{q})$) denote the geometric Frobenius endomorphism, and let ${}_{{\mathrm{et}}}\varphi_{\ell,x}$ denote the automorphism of ${}_{{\mathrm{et}}}\mathbb{V}_{\ell,x}$ induced by $\sigma$. Let $h\in G(\mathbb{Q}_{\ell})$ be some element, and let $\tau_{h}(y)=\{y_{1}\ldots y_{n}\}$. We have $y_{i}$ is in $\mathscr{S}_{K}(G,X)(\mathcal{O}_{F})$ (resp. $\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})$) if $\sigma(t_{i}^{-1}h\mathbb{V}_{\ell})=t_{i}^{-1}h\mathbb{V}_{\ell}$. Further, if we were to identify $\mathbb{V}_{\ell}$ with ${}_{{\mathrm{et}}}\mathbb{V}_{\ell,y}$ using $t_{i}$, we obtain the identification $h\mathbb{V}_{\ell}\cong_{{\mathrm{et}}}\mathbb{V}_{\ell,y_{i}}$, with the Frobenius at $y_{i}$ acting as $h^{-1}_{{\mathrm{et}}}\varphi_{\ell,x}h$.
7.1 Special points
Proposition 7.5.
Let $x\in S_{K}(G,X)(\mathbb{C})$. The following are equivalent:
  1. 1.
    $x$ is a special point.
  2. 2.
    For infinitely many primes $\ell$ at which $G$ is split, there exists a finite set $S:=\{x=x_{1},\ldots,x_{n}\}\subset S_{K}(G,X)(\mathbb{C})$ and a maximal split torus $T\subset G$ defined over $\mathbb{Q}_{\ell}$ such that $\tau_{h}(x)\cap S\neq\emptyset$ for every $h\in T(\mathbb{Q}_{\ell})$.
Proof.
Assume 1. Then there is a $\mathbb{Q}$-torus $T\subset G$ and a cocharacter $h:\mathbf{S}^{1}\rightarrow T_{\mathbb{R}}$ which lands in $X$. By enlarging $T$ assume that $T$ is maximal. Then conjugating by elements of $T$ does not change $h$, and the statement follows.
Now assume 2. Let $y=y_{1},\dots,y_{n}\in G(\mathbb{A})/K$ be lifts of the $x_{i}$. Without loss of generality we may assume that for every $i=1,\dots,n$ there is an element $c_{i}\in G(\mathbb{Q})$ such that $c_{i}\cdot y=y_{i}$. Moreover, we let $z_{i}:=y_{i,\mathbb{R}}\in X.$
Now let $H\subset G_{\mathbb{Q}}$ denote the largest $\mathbb{Q}$-subgroup fixing $z=z_{1}$. Note that $H(\mathbb{R})$ is compact and therefore $H$ is reductive. Let $\ell$ be a sufficiently large prime such that $G,H$ are both split, $K_{\ell}$ and $K_{\ell}\cap H(\mathbb{Q}_{\ell})$ are hyperspecial, the torus $T$ exists as in assumption 2, and all the $y_{i}$ have component $1\cdot K_{\ell}$ at $\ell$. Then $\forall h\in T(\mathbb{Q}_{\ell}),\tau_{h^{-1}}(y)$ consists of the union of $h_{i}\cdot y$ where the multiplication takes place in the $\ell$’th co-ordinate only, and $h_{i}$ ranges over elements of $K_{\ell}h^{-1}K_{\ell}$. Therefore, there is an element $\gamma\in G(\mathbb{Q})$ such that
  • $\gamma_{\ell}\in K_{\ell}hK_{\ell}$
  • $\gamma\cdot z\in\{z_{1},\dots,z_{n}\}$
The second condition is that for some $i,\gamma\in c_{i}\cdot H$, and thus we see that
$\bigcup_{i=1}^{m}c_{i}\cdot H(\mathbb{Q}_{\ell})\twoheadrightarrow K_{\ell}\backslash G(\mathbb{Q}_{\ell})/K_{\ell}.$
Let $Z\subset G$ be a maximal, split $\mathbb{Q}_{\ell}$-torus such that $Z^{\prime}:=Z\cap H$ is also a maximal, split $\mathbb{Q}_{\ell}$ torus of $H$. Let $W,W^{\prime}$ be the Weyl groups of $(G,Z),(H,Z^{\prime})$. By [tits, Pg. 51], we have
$K_{\ell}\backslash G(\mathbb{Q}_{\ell})/K_{\ell}\cong X_{*}(Z)/W$
and similarly
$(K_{\ell}\cap H)\backslash H(\mathbb{Q}_{\ell})/(K_{\ell}\cap H)\cong X_{*}(Z^{\prime})/W^{\prime}.$
Let ${P_{H}}^{+}\subset P^{+}$ be the coroots in a positive Weyl Chamber for $Z^{\prime},Z$ respectively. Let us write $\bigcup_{i=1}^{m}K_{\ell}c_{i}C_{\ell}$ as $\bigcup_{\lambda\in S}K_{\ell}\lambda(\ell)K_{\ell}$ for $S\subset P^{+}$. Then
$G(\mathbb{Q}_{\ell})$ $=\bigcup_{i=1}^{m}K_{\ell}c_{i}\cdot H(\mathbb{Q}_{\ell})K_{\ell}\subset\left(\bigcup_{i=1}^{m}K_{\ell}c_{i}K_{\ell}\right)\cdot K_{\ell}H(\mathbb{Q}_{\ell})K_{\ell}$
$\subset\bigcup_{\lambda\in S}K_{\ell}\lambda(\ell)K_{\ell}\bigcup_{\lambda^{\prime}\in P_{H}^{+}}K_{\ell}\lambda^{\prime}(\ell)K_{\ell}.$
Note that the containments must in fact be equalities. By [tits, Pg. 148] we see that this latter product contains $\mu(\ell)$ for $\mu\in P^{+}$ only if $\mu$ is dominated by some root in $S+P_{H}^{+}$. But $S$ is finite, $P_{H}^{+}$ is contained in a sublattice of $X_{*}(Z)$, so this is only possible if $Z^{\prime}=Z$.
Thus $H$ has the same rank as $G$. Thus $z$ has a stabilizer containing a maximal $\mathbb{Q}$-torus, and therefore $x$ is a special point, as desired.
Let $R$ denote a ring of characteristic $p$. In the case of Shimura varieties of Hodge type, one defines an $R$-valued point to be special if the abelian variety associated to the point has complex multiplication. Guided by Proposition 7.5, we make the following definition.
Definition 7.6.
Let $R$ be a ring of characteristic $p$. We say that $x\in\mathscr{S}_{K}(G,X)(R)$ is special if there exists a finite set $x=x_{1},x_{2},\ldots x_{n}\in\mathscr{S}_{K}(G,X)$, and split maximal tori $T_{\ell}\subset G_{\mathbb{Q}_{\ell}}$ for infinitely many primes $\ell$ such that $\tau_{h}(x)\cap\{x_{1},\ldots,x_{n}\}\neq\emptyset$ for every $h\in T(\mathbb{Q}_{\ell})$.
We will end this section by proving Theorem 7.1 and using it to prove that every $\overline{\mathbb{F}}_{p}$-valued point of $\mathscr{S}_{K}(G,X)$ is special as per our definition.
7.2 Tate semisimplicity
We will now prove Theorem 7.1.
Proof.
We first fix a $G$-equivariant identification ${}_{{\mathrm{et}}}\mathbb{V}_{\ell,x}\cong\mathbb{V}_{\ell}$. Suppose that ${}_{{\mathrm{et}}}\varphi_{\ell,x}$ doesn’t act semisimply – let $(_{{\mathrm{et}}}\varphi_{x,\ell})=\varphi_{\operatorname{ss}}\cdot\varphi_{\operatorname{unip}}$. Note that the Jordan decomposition implies that $\varphi_{\operatorname{ss}},\varphi_{\operatorname{unip}}$ are elements of $G(\mathbb{Q}_{\ell})$. A sufficiently $\ell$-divisible power of $\varphi_{\operatorname{unip}}$ will preserve $\mathbb{V}_{\ell,x}$ and hence by increasing the field $\mathbb{F}_{q}$ we may and do assume $\varphi_{\operatorname{ss}},\varphi_{\operatorname{unip}}$ lies in $K_{\ell}$.
We will find hecke operators in $G(\mathbb{Q}_{\ell})$ which preserve $\mathbb{F}_{q}$-rationality, and therefore which eventually give us that our point is stable under that appropriate operator, which we will use to deduce semisimplicity. In order to preserve $\mathbb{F}_{q}$-rationality, we pick our Hecke operator as follows. Let $H\subset G$ denote the centralizer of $\varphi_{\operatorname{ss}}$ – it contains $\varphi_{x}$ and $\varphi_{\operatorname{unip}}$. Note that $H$ is semisimple by [Conrad:ReductiveSchemes]. Using the Jacobson-Morozov lemma, we can find an element $h\in H(\mathbb{Q}_{\ell})$ such that $h^{-1}\varphi_{\operatorname{unip}}h=\varphi_{\operatorname{unip}}^{\ell^{2}}$. Let $\tau_{h}$ denote the corresponding Hecke correspondence. One of the points of $\tau_{h}(x)$ (which we call $x_{1}$) will correspond to the lattice $h\mathbb{V}_{\ell}\subset V_{\ell}\otimes\mathbb{Q}_{\ell}$, and as this lattice is stable under the action of $\varphi$, the point will also be defined over $\mathbb{F}_{q}$ by Lemma 7.4. The action of $\varphi_{x_{1}}$ is just $\varphi_{x}$ on the lattice $h\mathbb{V}_{\ell}\subset V_{\ell}$, again by Lemma 7.4. Therefore, we have that $\varphi_{x_{1}}=h^{-1}\varphi_{x}h=\varphi_{\operatorname{ss}}\varphi_{\operatorname{unip}}^{\ell^{2}}$ by construction. Therefore, $\varphi_{x_{1},\operatorname{unip}}-1$ is strictly more divisible by $\ell$ then $\varphi_{x,\operatorname{unip}}-1$, and therefore $x\neq x_{1}$.
Iterating this process over and over, we get a sequence of mutually distinct $k$-rational points. This is contradiction arising from our assumption that $\varphi_{\operatorname{unip}}$ is nontrivial.
Theorem 7.1 and its proof has the following immediate consequence.
Corollary 7.7.
Every $x\in\mathscr{S}_{K}(G,X)(\overline{\mathbb{F}}_{p})$ is special.
Proof.
It suffices to prove that the image of $x$ in the adjoint Shimura variety is special, and so we suppose that $G$ is an adjoint group. The point $x$ is defined over some $\mathbb{F}_{q}$. Let $H$ denote the centralizer of ${}_{\ell}\phi_{x}$, where ${}_{\ell}\phi_{x}$ is the $\mathbb{F}_{q}$-Frobenius at $x$ acting on ${}_{{\mathrm{et}},\ell}\mathbb{V}_{x}$. The argument in the proof of Theorem 7.1 yields that $\tau_{h}(x)\cap\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})\neq\emptyset$ for every $h\in H(\mathbb{Q}_{\ell})$. By Theorem 7.1, we have that $H$ must contain a maximal torus of $G$ defined over $\mathbb{Q}_{\ell}$. Therefore, if we were to set $S=\{x=x_{1},\ldots x_{n}\}=\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})$ and $T$ to be some maximal torus of $H$ defined over $\mathbb{Q}_{\ell}$, we would have that $\tau_{h}(x)\cap S\neq\emptyset$ for every $h\in T(\mathbb{Q}_{\ell})$. Therefore, it suffices to prove that there exist infinitely many $\ell$ such that some power of ${}_{\ell}\phi_{x}$ is contained in a split maximal $\mathbb{Q}_{\ell}$-torus of $G$.
By the main theorem of [StefanChristian], the conjugacy class of ${}_{\ell}\phi_{x}$ is $\mathbb{Q}$-rational, and does not depend on $\ell$. Therefore, we may choose $\ell$ such that $G_{\mathbb{Q}_{\ell}}$ is split, and such that the characteristic polynomial of ${}_{\ell}{\phi_{x}}$ on some faithful representation of $G$ splits over $\mathbb{Q}_{\ell}$. By replacing ${}_{\ell}\phi_{x}$ by a power (i.e. replacing $q$ by a power), we may also assume that the roots of this characteristic polynomial do not differ by a non-trivial root of unity. It follows that ${}_{\ell}\phi_{x}$ is contained in split torus. However, in a quasi-split (and therefore split) group, every split torus is contained in a maximal split torus.
8 Complete local rings for large primes
Let notation be as in Section 5.1, and let $p\nmid N$ denote a prime. As we will exclusively work over a $p$-adic field, we let $G$ denote $\mathcal{G}_{\mathbb{Z}_{p}}$. Throughout this section, the level structure at $p$ is $G(\mathbb{Z}_{p})$.
The goal of this section is to fully describe the Fontaine-Laffaille data restricted to complete local rings of $S_{K}(G,X)$ at $\overline{\mathbb{F}}_{p}$-points $x$, and to describe the Fontaine-Laffaille data at $W$-valued points of $S_{K}(G,X)$ in terms of picking a filtration on an appropriate $F$-crystal. For such a point $x$, we let $\mathscr{S}_{K}(G,X)^{x}$ denote the formal neighbourhood of $\mathscr{S}_{K}(G,X)$ at $x$. Let $R^{x}$ denote the complete local ring at $x$, and let ${}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$ denote the restriction of ${}_{{}_{\mathrm{cris}}}\mathbb{V}$ to $\mathscr{S}_{K}(G,X)^{x}$. Similarly, we let ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$ denote the restriction of ${}_{{}_{\mathrm{cris}}}\mathbb{V}$ to any $W$-point of $\mathscr{S}_{K}(G,X)$. Let ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$ denote the induced $F$-crystal at $x\in\mathscr{S}_{K}(G,X)(\mathbb{F})$. We will treat ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$ as a free $W$-module equipped with a semi-linear Frobenius endomorphism. We let $\varphi_{u}$ denote Frobenius on ${}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$, and $\varphi_{\tilde{x}}$ denote Frobenius on ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}$.
8.1 The Fontaine-Laffaille module on complete local rings
8.1.1 F-crystals with $G$-structure and $G$-split filtrations.
We refer to [ShankarZhou, Section 2] for the notion of an $F$-crystal with $G$ structure. Recall that we have tensors $s_{\alpha}\in\mathbb{V}_{\frac{1}{N}}^{\otimes}$. These tensors induce global sections of the $p$-adic etale local system ${}_{{\mathrm{et}}}\mathbb{V}_{p}^{\otimes}$ (defined over $\mathscr{S}_{K}(G,X)/\operatorname{\mathcal{O}}_{E}[\frac{1}{pN}]$) – we will refer to them as etale tensors ${}_{{\mathrm{et}}}s_{\alpha,p}\in_{{\mathrm{et}}}\mathbb{V}_{p}^{\otimes}$.
Via the Riemann-Hilbert correspondence, the tensors $s_{\alpha}$ give rise to flat sections ${}_{{}_{\mathrm{dR}}}s_{\alpha}\in_{{}_{\mathrm{dR}}}V^{\otimes}$. We have the following proposition:
Proposition 8.1.
The tensors ${}_{{}_{\mathrm{dR}}}s_{\alpha}$ descend to integral tensors in ${}_{{}_{\mathrm{dR}}}\mathbb{V}^{\otimes}$ over $\mathcal{O}_{E}[\frac{1}{N}]$.
Proof.
This follows directly from the existence of the integral model $\mathscr{E}/\mathcal{O}_{E}[\frac{1}{N}]$ of the flat principal $\mathcal{G}$-bundle. Then we have the isomorphism
$_{{}_{\mathrm{dR}}}\mathbb{V}^{\otimes}\simeq\mathscr{E}\times_{\mathcal{G}}\mathbb{V}^{\otimes}$ (8.1.1)
over $\mathcal{O}_{E}[\frac{1}{N}]$. Using this isomorphism, the tenors $s_{\alpha}$ induce canonical flat sections of ${}_{{}_{\mathrm{dR}}}\mathbb{V}^{\otimes}$. These are precisely the flat sections ${}_{{}_{\mathrm{dR}}}s_{\alpha}$, and the proposition follows.
Let $v$ denote some prime of $E$ above $p$, and consider the $p$-adic etale local system ${}_{{\mathrm{et}}}\mathbb{V}_{p}/S_{K}(G,X)/E_{v}$. As in [Kisinintegral, Section 1.3.3], we may apply the Fontaine-Laffaille functor to the tensors ${}_{{\mathrm{et}}}s_{\alpha,p}$ to obtain tensors ${}_{{}_{\mathrm{cris}}}s_{\alpha}\in{{}_{{}_{\mathrm{cris}}}\mathbb{V}^{\otimes}}$, which by construction are fixed by Frobenius and contained in ${\mathrm{Fil}}^{0}$.
Recall that the integral structures on ${}_{{}_{\mathrm{dR}}}V/S_{K}(G,X)_{E_{v}}$ induced by ${}_{{}_{\mathrm{cris}}}\mathbb{V}$ and ${}_{{}_{\mathrm{dR}}}\mathbb{V}$ are the same. We have the following proposition:
Proposition 8.2.
Under the identifications above, we have ${}_{{}_{\mathrm{cris}}}s_{\alpha}={{}_{{}_{\mathrm{dR}}}s_{\alpha}}$.
Proof.
As both sets of tensors are flat, it suffices to prove that the two sets of tensors agree at some point – we will show that they agree at a special point. Indeed, at a special point $y$, all the data $\mathbb{V}$ arise from the cohomology of a CM abelian variety. Similar to [Kisinintegral, Section 2.2.2], the equalities follow from the work of Blasius-Wintenberger [Blasius]. ∎
Corollary 8.3.
There is an (non-canonical [^5] The non-canonicity is because any choice of isomorphism can be pre-composed by some element of $G(R^{x})$.) isomorphism of $R^{x}$-modules $\iota^{x}:\mathbb{V}_{p}\otimes R^{x}\rightarrow{{}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}}$ which sends the tensors $s_{\alpha}$ to ${}_{{}_{\mathrm{cris}}}s_{\alpha}$.
Proof.
This follows directly from the ismorphism in Equation (8.1.1) and the fact that $\mathscr{E}|_{{\mathscr{S}_{K}(G,X)}^{x}}$ is trivializable (Note that the connection however is not trivializable as a principal bundle). ∎
Now, let $\tilde{x}\in\mathscr{S}_{K}(G,X)(W)$ denote a $W$-valued lift of $x$. Let ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}$ denote the Fontaine-Laffaille module pulled back to $\tilde{x}$. Corollary 8.3 has the following immediate implication:
Corollary 8.4.
[…]
We will now prove that the second isomorphism in Corollary 8.4 can be chosen to respect filtrations.
Proposition 8.5.
The isomorphism $\iota$ in Corollary 8.4 may be chosen such that ${\mathrm{Fil}}^{\bullet}(_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}})\otimes R^{x}$ maps isomorphically onto ${\mathrm{Fil}}^{\bullet}(_{{}_{\mathrm{cris}}}\mathbb{V}^{x})$.
Proof.
Consider the subgroups of $\operatorname{GL}(_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}})$ and $\operatorname{GL}(_{{}_{\mathrm{cris}}}\mathbb{V}^{x})$ that fix the tensors $s_{\alpha}$ and the filtrations ${\mathrm{Fil}}^{\bullet}$. The existence of the integral model of $\mathscr{P}/\mathscr{S}_{K}(G,X)$ implies that these are parabolic subgroups $P_{W}\subset G_{W}$ and $P_{R^{x}}\subset G_{R^{x}}$ respectively – the identification $\iota_{\tilde{x}}$ and $\iota^{x}$ in Corollaries 8.3 and 8.4 allow us to think of the parabolic subgroups as subgroups of $G_{W}$ and $G_{R^{x}}$. Further, $P_{R^{x}}$ clearly specializes to $P_{W}$ via the point $\tilde{x}$. Consider now the two parabolic subgroups $P_{R^{x}}$ and $P_{W}\otimes R^{x}$ of $F_{R^{x}}$. These two subgroups are equal (and therefore conjugate) at the unique closed point. By [Conrad:ReductiveSchemes, Corollary 5.2.7], the two parabolics are $\mathcal{G}(R^{x})$-conjugate. Conjugating $\iota$ by this element still respects the tensors (which are fixed by elements of $G$) and now respects the filtrations by construction. The proposition follows. ∎
Finally, we have that ${\mathrm{Fil}}^{\bullet}$ is induced by a $G$-split co-character. In other words, we have:
Corollary 8.6.
The filtration on ${\mathrm{Fil}}^{\bullet}(_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}})$ is induced by a $G_{W}$-valued co-character.
Proof.
We will let $\tilde{x}$ also denote the $W[1/p]$-point of $\mathscr{S}_{K}(G,X)$. By [Kisinintegral, Lemma 1.1.4], it suffices to prove that the filtration on ${}_{{}_{\mathrm{dR}}}V_{\tilde{x}}$ is induced by a $G_{W[1/p]}$-valued co-character. Further, we have (by [DLLZ]) that ${}_{{}_{\mathrm{dR}}}V_{\tilde{x}}$ and $D_{{}_{\mathrm{dR}}}(_{{\mathrm{et}}}V_{p,\tilde{x}})$ are isomorphic as filtered vector spaces. The fact that the filtration on $D_{{}_{\mathrm{dR}}}(_{{\mathrm{et}}}V_{p,\tilde{x}})$ is induced by a $G$-valued co-character now follows from [Kisinintegral, Lemma 1.4.5]. ∎
8.1.2 Explicit coordinates
We pick coordinates on $R^{x}$ and write it as $W[[x_{1},\ldots,x_{n}]]$, such that the point $\tilde{x}$ is given by setting all the $x_{i}=0$. We choose a lift of Frobenius $\sigma_{u}$ on $R^{x}$ that extends the usual Frobenius $\sigma$ on $W$ and satisfies $\sigma_{u}(x_{i})=x_{i}^{p}$. Using the isomorphisms $\iota_{\tilde{x}},\iota^{x}$, we may now write $\varphi_{u}=b_{u}\sigma_{u}$ and $\varphi_{\tilde{x}}=b\sigma$, where the fiber of $b_{u}\in G(R^{x}[1/p])$ at $\tilde{x}$ is $b\in G(W[1/p])$. We also fix a choice of $G_{W}$-valued co-character $\mu$ that induces the filtration on ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}$) and on ${}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$ via $\iota$.)
Proposition 8.7.
We have $b_{u}\in G(R^{x})\sigma(\mu(p))$.
Proof.
[…]
We have the immediate corollary:
Corollary 8.8.
We have $b\in G(W)\sigma(\mu(p))$.
8.1.3 The Fontaine-Laffaille module on complete local rings
We will now prove the main result of this section, which is to give a description of the Fontaine-Laffaille data on $S_{K}(G,X)^{x}$ analogous to the case of abelian type Shimura varieties following Kisin [Kisinintegral]. Let notation be as in the previous subsections. Recall that we have chosen an isomorphism ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}\otimes R^{x}\rightarrow_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$ of filered modules compatible with the $G$-structure defined in Subsection 8.1.1. We note that such an isomorphism is unique only up to an element of $P(R^{x})$, where $P$ is the parabolic subgroup of $G$ defined by the filtration. Also recall that we have chosen a $G$-valued co-character (defined over $W$) that splits ${\mathrm{Fil}}^{\bullet}$. Let $U^{\operatorname{opp}}$ denote the opposite unipotent of $\mu$, and let $\widehat{U}^{\operatorname{opp}}$ denote the completion of $U^{\operatorname{opp}}$ at the identity mod $p$.
We now fix $\iota_{\tilde{x}}$. We note that the choice of $\iota$ fixes the choice of $\iota^{x}$.
Theorem 8.9.
The isomorphism of filtered modules $\iota:_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}\otimes R^{x}\rightarrow_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$ can be chosen so that $b_{u}=u\cdot b$, where $u\in U^{\operatorname{opp}}(R^{x})$. Further, the map induced by the point $u:\operatorname{Spf}R^{x}\rightarrow\widehat{U}^{\operatorname{opp}}$ is an isomorphism.
Proof.
We will first prove that we may choose $\iota$ such that $u\in\widehat{U}^{\operatorname{opp}}$, and will then use the versality of the Kodaira Spencer map to show that the tautologocal map $u$ induces is an isomorphism. The first part of the proof will follow [KLSS, Proposition 4.7].
Let $I\subset R^{x}$ denote the kernel of $R^{x}\xrightarrow{\tilde{x}}W$. We have that $b_{u}=u^{\prime}b$, where $u^{\prime}\in G(R^{x})$ with $u^{\prime}\equiv\operatorname{Id}\bmod I$. Therefore, we also have that $u^{\prime}\in U^{\operatorname{opp}}\bmod I$. We may write $u^{\prime}=\lambda u$ with $\lambda\in P$ and $u\in U^{\operatorname{opp}}$, with both $\lambda$ and $u$ reducing to the identity modulo $I$. Replacing $\iota$ by $\lambda\circ\iota$ has the effect of replacing $b_{u}$ by $\lambda^{-1}u^{\prime}b\sigma(\lambda)=u(b\sigma(\lambda)b^{-1})b$. Define $u^{\prime}(1)$ to be $u(b\sigma(\lambda)b^{-1})$. As $\sigma(I)$ is generated by $t_{i}^{p}$, we have that $u^{\prime}(1)\in\widehat{U}^{\operatorname{opp}}\bmod I^{p}$, and we may thus write $u^{\prime}(1)=\lambda(1)u(1)$, with $\lambda(1)\in P$ reducing to the identity modulo $I^{p}$ and $u(1)$ reducing to the identity modulo $I$. Iterating this process, we may change coordinates by an element of $P$ again and replace $b_{u}$ by $u^{\prime}(2)b=u(1)(b\sigma(\lambda(1))b^{-1})b$, whence $u^{\prime}(2)\in U^{\operatorname{opp}}\bmod I^{p^{2}}$. Iterating this process is equivalent to modifying $j$ by the convergent product $\ldots\lambda(2)\cdot\lambda(1)\cdot\lambda$, and has the effect of writing $b_{u}=u\cdot b$ with $u\in U^{\operatorname{opp}}(R^{x})$ as required.
We will now prove that the tautological map $\operatorname{Spf}R^{x}\rightarrow\widehat{U}^{\operatorname{opp}}$ induced by $u$ is an isomorphism. The leading order term of the connection is $du$ (see [Kisinintegral, Proof of Lemma 1.5.2] [^6] Note that in loc. cit., the leading order term is described as $u^{-1}\cdot du$. As $u\equiv\operatorname{Id}$ mod $I$, $du$ and $u^{-1}du$ have the same leading term.). Therefore, the Kodaira spencer map reduces mod $I$ to the element $du_{\tilde{x}}:\operatorname{Hom}(T_{\tilde{x}}R^{x}\rightarrow\operatorname{\mathsf{End}}(\operatorname{Gr}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}))$, where $du_{\tilde{x}}$ is the restriction of $du$ to the tangent space of $T_{\tilde{x}}T^{x}$ of $R^{x}$ at $\tilde{x}$. The image of $du_{\tilde{x}}$ is contained in $\operatorname{Lie}U_{\tilde{x}}^{\operatorname{opp}}$. The versality of the Kodaira-Spencer map implies that $du_{\tilde{x}}$ is an immersion. As the dimension of $\widehat{U}^{\operatorname{opp}}$ is the same as the dimension of $R^{x}$, it follows that $u$ is an isomorphism.
8.2 Versality of Filtrations
The setting and notation will be as in Section 8.1.3. Let $\tilde{x}^{\prime}$ denote a $W$-valued lift of $x$. The $F$-crystals underlying the Fontaine-Laffaille modules ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}$ and ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}^{\prime}}$ are canonically isomorphic, and indeed are canonically isomorphic to the F-crystal ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$. This canonical isomorphism is obtained by considering the horizontal continuation $s_{\nabla}$ of elements $s\in_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}$ to ${}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$ with respect to the connection $\nabla$ on ${}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$. We will use this identification to view ${\mathrm{Fil}}^{\bullet}(_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}})$ and ${\mathrm{Fil}}^{\bullet}(_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}^{\prime}})$ as filtrations on ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$, i.e. different filtrations on the same underlying $F$-crystal. The content of the following proposition is that points $\tilde{x}\in R^{x}(W)$ are in bijection with filtrations of the form $g\cdot{\mathrm{Fil}}^{\bullet}(_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}})$, where $g\in\widehat{U}^{\operatorname{opp}}(W)$. This result will be used crucially in a later section to define the canonical lift of ($\mu$-)ordinary and then prove our CM lifting theorems.
Proposition 8.10.
There is a bijection between $\widehat{U}^{\operatorname{opp}}(W)$ and points $\tilde{x}^{\prime}\in R^{x}(W)$, where the bijection assigns to $g\in\widehat{U}^{\operatorname{opp}}(W)$ the Fontaine-Laffaille module with underlying $F$-crystal ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$ and filtration $g\cdot{\mathrm{Fil}}^{\bullet}(_{{}_{\mathrm{cris}}}\mathbb{V}^{\tilde{x}})$.
We will first need the following result about flat sections on ${}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$. Let $\mathcal{O}$ denote the ring of rigid-analytic functions on the rigid-analytic space given by the tube of $\mathscr{S}_{K}(G,X)$ over $x$. The tube is just a $S_{K}(G,X)$-dimensional open unit ball around $\tilde{x}$. Note that $R^{x}[1/p]\subset\mathcal{O}$ is the subset of bounded functions. Recall that we have the coordinates $x_{i}\in R_{x}$ whose vanishing defines $\tilde{x}$. Let $y_{i}\in\mathcal{O}$ be defined by $y_{i}=\frac{x_{i}}{p}$. We have the inclusion $\mathcal{O}\subset W[\frac{1}{p}]\langle y_{i}\rangle$, given by restricting functions on the open unit ball to the closed ball with radius $\frac{1}{p}$. We let $W\langle y_{i}\rangle$ denote functions on this closed ball that are bounded above by 1. We are now ready to state the following lemma.
Lemma 8.11.
We have the following results:
  1. 1.
    The connection $\nabla$ has a basis of solutions over $\mathcal{O}$. Further, $(_{{}_{\mathrm{cris}}}\mathbb{V}^{x},\nabla)|_{W\langle y_{i}\rangle}$ is isomorphic to $(_{{}_{\mathrm{cris}}}\mathbb{V}^{x}_{\nabla=0}\otimes_{W}W\langle y_{i}\rangle,d)$ Consequently, the horizontal continuation of every $s\in_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}$ is an element of ${}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}\otimes_{R^{x}}W\langle y_{i}\rangle$.
  2. 2.
    Let $s\in_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}$ be some element. Let $\tilde{s}$ denote the element $s\otimes 1\in{{}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}}$ under the isomorphism in Theorem 8.9. Let $s_{\nabla}$ denote the horizontal continuation of $s$ to ${}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}\otimes\mathcal{O}$. There exists a unique element $g\in G(\mathcal{O})$ such that $g\cdot\tilde{s}=s_{\nabla}$ where $g$ is independent of $s$. Further, $g\in G(W\langle y_{i}\rangle)$.
  3. 3.
    $g\equiv u^{-1}\mod(x_{i})^{2}$, where $u$ is the tautological point $u\in\widehat{U}^{\operatorname{opp}}(R^{x})$ as in Theorem 8.9. Further, the coefficients of the series expansion of $g$ in the coordinates $y_{i}$ are divisible by $p^{2}$ for the non-linear and non-constant terms.
Proof.
This lemma is proved by direct calculation. First note that $(_{{}_{\mathrm{cris}}}\mathbb{V}^{x},\nabla)$ has a full set of solutions over $W[1/p][[x_{i}]]$, with flat sections given by:
$s_{\nabla}=\sum_{\vec{w}}\prod_{i}\frac{(-x_{i})^{w_{i}}}{i!}\prod_{i}\nabla(\frac{d}{dx_{i}})^{w_{i}}(\tilde{s}),$ (8.2.1)
where the sum is taken over all tuples $\vec{w}=(w_{i})$ of non-negative integers.
  1. 1.
    That $\nabla$ has a basis of solutions over $\mathcal{O}$ follows from the Frobenius structure (Dwork’s trick). The fact that the flat bundle $(_{{}_{\mathrm{cris}}}\mathbb{V}^{x},\nabla)|_{W\langle y_{i}\rangle}$ is trivial is a general fact about vector bundles with flat connection on $W[[x_{i}]]$ pulled back to $W\langle y_{i}\rangle$ for primes $p>2$. Indeed, the connection matrix has entries valued in $W[[x_{i}]]$, and after substituting $x_{i}=py_{i}$, it is evident that the resulting expressions are contained in $W\langle y_{i}\rangle$, and indeed, we have $s_{\nabla}|_{W\langle y_{i}\rangle}\equiv\tilde{s}|_{W\langle y_{i}\rangle}\bmod p$.
  2. 2.
    By [Kisinintegral, Lemma 1.5.2] and [Kisinmodppoints, E.1] [^7] As the erratum indicates, the conclusion that $\nabla\in\mathrm{Lie}U\otimes\Omega^{1}_{R^{x}}$ is false without extra assumptions (for example, the point $x$ being ordinary), but the weaker claim $\nabla\in\mathrm{Lie}G\otimes\Omega^{1}_{R^{x}}$ is true., it follows that $\nabla\in\mathrm{Lie}G\otimes\Omega^{1}_{R^{x}}$. Therefore, $g$ is indeed a point of $G$, a-priori valued in $W[1/p][[x_{i}]]$. The fact that it is an $\mathcal{O}$-valued point (and indeed, a $W\langle y_{i}\rangle$-valued point ) of $G$ follows from the first part (and the fact that $G(W\langle y_{i}\rangle)=\operatorname{GL}(W\langle y_{i}\rangle)\cap G(W[1/p][[x_{i}]])$).
  3. 3.
    This follows from Equation (8.2.1) and the fact that the leading term of $\nabla$ is $du$.
With this in hand, we will now prove Proposition 8.10:
Proof.
Let $h=g^{-1}$. By Lemma 8.11 (3), we have that $h\equiv u\bmod(y_{i})^{2}\cap(p^{2})$. We have that $\tilde{s}=h\cdot s_{\nabla}$.
We let ${\mathrm{Fil}}^{\bullet}$ denote the filtration ${\mathrm{Fil}}^{\bullet}(_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}})$. The filtration on ${}_{{}_{\mathrm{cris}}}\mathbb{V}^{x}$ is given by $\widetilde{{\mathrm{Fil}}^{\bullet}}={\mathrm{Fil}}^{\bullet}\otimes_{W}R^{x}$. Let ${\mathrm{Fil}}^{\bullet}_{\nabla}$ denote the horizontal continuation of ${\mathrm{Fil}}^{\bullet}$. We then have that $h\cdot{\mathrm{Fil}}^{\bullet}_{\nabla}=\widetilde{{\mathrm{Fil}}^{\bullet}}$. Let $P_{{\mathrm{Fil}}^{\bullet}_{\nabla}}\subset G$ be the parabolic subgroup of $G$ stabilizing the filtration ${\mathrm{Fil}}^{\bullet}_{\nabla}$. We may write $h=u^{\prime}\cdot\lambda^{\prime}$, where $u^{\prime}\in U^{\operatorname{opp}}(W\langle y_{i}\rangle)$ and $\lambda^{\prime}\in P_{{\mathrm{Fil}}^{\bullet}_{\nabla}}(W\langle y_{i}\rangle)$ – we have that $\widetilde{{\mathrm{Fil}}^{\bullet}}=u^{\prime}\cdot{\mathrm{Fil}}^{\bullet}_{\nabla}$. Note that $u^{\prime}\equiv h\mod(y_{i})^{2}\cap(p^{2})$.
Let $U^{\operatorname{opp}}_{1/p}\subset\widehat{U}^{\operatorname{opp},\textrm{rig}}$ denote the closed ball of radius $\frac{1}{p}$ around the identity. The element $u^{\prime}$ induces a rigid-analytic map $\operatorname{Sp}W\langle y_{i}\rangle\rightarrow\widehat{U}^{\operatorname{opp},\textrm{rig}}$, and the conditions on $u^{\prime}$ imply that the image is contained in $U^{\operatorname{opp}}_{1/p}$. Note that unipotent groups are uniquely $p$-divisible, and so the map “dividing by $p$” gives an isomorphism $U^{\operatorname{opp}}_{1/p}\rightarrow U^{\operatorname{opp},\textrm{rig}}$. Here, we abuse notation and also allow $U^{\operatorname{opp}}$ to denote the formal scheme obtained by $p$-adic completing $U^{\operatorname{opp}}$. Therefore, by composing $u^{\prime}$ by the “dividing by $p$” map, we obtain a map $u^{\prime\prime}:\operatorname{Sp}W[1/p]\langle y_{i}\rangle\rightarrow U^{\operatorname{opp},\textrm{rig}}$. Both these rigid spaces admit natural integral models – $\operatorname{Spf}W\langle y_{i}\rangle$ and $U^{\operatorname{opp}}$. The congruence conditions on $u^{\prime}$ imply that $u^{\prime\prime}$ is induced by a map on these integral models. It suffices to prove that $u^{\prime\prime}$ is an isomorphism, which would follow from the claim mod $p$. The congruences established in Lemma 8.11 together with Theorem 8.9 imply that $u^{\prime\prime}$ is an isomorphism modulo $(y_{i})^{2}$, and that the coefficients in the series expansion of $u^{\prime\prime}$ in the coordinates $y_{i}$ are divisible by $p$ for the non-linear and non-constant terms. It is now clear $u^{\prime\prime}$ is an affine linear map mod $p$, and it suffices to check that it is an isomorphism at the tangent space of some closed point. But this follows from the fact that $u^{\prime\prime}$ is an isomorphism mod $(y_{i})^{2}$. The proposition follows from here.
9 The $\mu$-ordinary locus: CM lifts and Tate’s isogeny theorem
In this section, we will define the notion of $\mu$-ordinariness. This notion generalizes the notion of ordinary abelian varieties. We will then prove that the $\mu$-ordinary locus is open-dense, and that every $\mu$-ordinary point admits a lift to a special point. This generalizes the Serre-Tate canonical lift of an ordinary abelian variety (and generalizes work of Moonen for Shimura varieties of PEL type, and work of [ShankarZhou] in the case of Hodge type Shimura varieties).
We keep the notation of the previous section. Let $B\subset G$ be a Borel subgroup defined over $\mathbb{Z}_{p}$, and let $T\subset B\subset G$ be a maximal torus that splits over an etale extension of $\mathbb{Z}_{p}$(that $G/\mathbb{Z}_{p}$ is reductive implies that these objects exist). Let $x\in\mathscr{S}_{K}(G,X)(\mathbb{F})$. Associated to the $F$-crystal ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$ is the Frobenius element $b_{x}\in G(L)$, which is well defined up to $\sigma$-conjugation by $G(W)$. Let $\nu_{x}\in X_{*}(T)$ be the dominant element that is conjugate to $\nu_{b_{x}}$, the Newton cocharacter associated to $b_{x}$ (see work of Kottwitz [Kott1] and [Kott2] for definitions and more details). Note that the conjugacy class of $\nu_{b_{x}}$ depends only on the $\sigma$-conjugacy class of $b_{x}$, and so the definition of $\nu_{x}$ is independent of the choice of representative of the $\sigma$-conjugacy class of $b_{x}$. We note that $\nu_{x}$ is defined over $\mathbb{Z}_{p}$. Let $\mu_{x}\in X_{*}(T)$ denote the dominant co-character with $b_{x}\in G(W)\sigma(\mu_{x})(p)G(W)$. The co-character $\mu_{x}$ might not be defined over $\mathbb{Z}_{p}$, and let $\bar{\mu}_{x}=\frac{1}{[E_{p}:\mathbb{Q}_{p}]}\sum_{\tau\in\operatorname{Gal}(E_{p}/\mathbb{Q}_{p})}\tau(\sigma)$ denote the Galois average of $\mu$, where $E_{p}/\mathbb{Q}_{p}$ is some finite unramified extension splitting $T$. We have that $\nu_{x}\preceq\bar{\mu_{x}}$ (for example, see [Gashi]), where two dominant fractional cocharacters satisfy $\mu^{\prime}\preceq\mu^{\prime\prime}$ if and only if $\mu^{\prime\prime}-\mu^{\prime}$ can be written as a non-negative rational linear combination of positive co-roots. We have the following definition.
Definition 9.1.
  1. 1.
    The point $x$ is said to be $\mu$-ordinary if $\nu_{x}=\bar{\mu}_{x}$.
  2. 2.
    The point $x$ is said to be ordinary if it is $\mu$-ordinary and if $\mu_{x}$ is defined over $\mathbb{Z}_{p}$.
We remark that when $G=T$ is itself a torus, then every point is $\mu$-ordinary. The notion of $\mu$-ordinariness was first introduced by Rapoport in his ICM talk (see also [Wortman], [MoonenSerreTate], and [ShankarZhou]). The $\mu$-ordinary locus will be open-dense if nonempty. This was done in [Wortman] in the case of Shimura varieties of abelian type, when the level structure at $p$ is hyperspercial (see also [Bultel]). We prove this in the case when $p$ is a large enough prime – we remark that our proof works both in the abelian and exceptional cases.
Theorem 9.2.
Let $p$ be a large enough prime and let $v\mid p$ be a prime of $E$. Then the $\mu$-ordinary locus of $\mathscr{S}_{K}(G,X)_{\mathbb{F}_{v}}$ is non-empty.
We will prove this result at the end of this section. The content of [Wortman, Proposition 7.2] is the following result:
Proposition 9.3 (Wortman).
There exists $g\in G(W)$ such that $g^{-1}b\sigma(g)=(\sigma\mu)(p)$.
We now change coordinates on (by $\sigma$-conjugating by an element of $G(W)$) and we may assume that $b=\sigma(\mu)p$. The following proposition defines the canonical lift of a $\mu$-ordinary point.
Proposition 9.4.
Consider the Fontaine-Laffaille module with underlying $F$-crystal ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{x}$, and with the filtration defined by the co-character $\mu_{x}$. This corresponds to a (necessarily) unique point of $S_{K}(G,X)^{x}$.
Proof.
This result follows almost immediately from Proposition 8.7. Indeed, let $\tilde{x}$ denote some $W$-valued lift of $x$, and let $\mu$ denote a co-character inducing the filtration on ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}$. The Fontaine-Laffaille conditions imply that filtrations induced by $\mu_{x}$ and $\mu$ agree on ${}_{{}_{\mathrm{cris}}}\mathbb{V}_{\tilde{x}}\bmod p$.
Consider the parabolic subgroups of $P_{x}\subset G$ and $P_{\mu}\subset G$ induced by $\mu_{x}$ and $\mu$ respectively. That the filtrations agree modulo $p$ implies that $P_{x}=P_{\mu}\bmod p$. By [Conrad:ReductiveSchemes, Corollary 5.2.7], the two parabolics are conjugate over some etale extension of $W$, and therefore over $W$ itself. As the two parabolics agree modulo $p$, they are conjugate by an element $g\in\hat{U}^{\operatorname{opp}}(W)$ (with $\hat{U}^{\operatorname{opp}}$ the opposite unipotent of $\mu$). Therefore, ${\mathrm{Fil}}_{x}=g\cdot{\mathrm{Fil}}_{\mu}$, and so the proposition follows by Proposition 8.7.
Definition 9.5.
Let $x\in S(G,X)(\mathbb{F})$ be a $\mu$-ordinary point. We let $x^{{\mathrm{can}}}$ denote $W$-valued lift constructed in Proposition 9.4, and call it canonical lift of $x$.
In the ordinary case, this construction directly generalizes the Serre-Tate canonical lift of an ordinary abelian variety. More generally, this generalizes the notion of canonical lifts of $\mu$-ordinary points of PEL-type Shimura varieties defined by Moonen in [MoonenSerreTate], and of $\mu$-ordinary points of Hodge-type Shimura varieties defined in [ShankarZhou]. Analogous to those cases, we will shortly prove that $x^{{\mathrm{can}}}$ is actually a CM point. For now, we remark that the canonical lift is functorial in prime-to-$p$ Hecke correspondence. Indeed, prime-to-$p$ Hecke correspondences are finite etale correspondences on the integral model, and therefore every $\tilde{y}\in\tau(\tilde{x})$ is a $W$-valued point of $\mathscr{S}_{K}(G,X)$ whenever $\tilde{x}$ is. Further, the action of prime-to-$p$ Hecke correspondences on $W$-valued points preserves the isomorphism class of the Fontaine-Laffaille modules attached to these points. Therefore, we have the immediate corollary:
Corollary 9.6.
Let $x^{{\mathrm{can}}}$ denote the canonical lift of a $\mu$-ordinary point $x$, and let $\tilde{y}\in\tau(x^{{\mathrm{can}}})$ where $\tau$ is a prime-to-$p$ Hecke correspondence. Let $y$ be the specialization of $\tilde{y}$. Then, $\tilde{y}=y^{{\mathrm{can}}}$. In particular, $\tilde{y}=x^{{\mathrm{can}}}$ if $y=x$.
We can now apply this in conjunction with Proposition 7.5 and Corollary 7.7 to obtain the following result:
Theorem 9.7.
Every $\mu$-ordinary point lifts to a characteristic-zero special point.
We obtain the following corollary.
Corollary 9.8.
Let $x\in\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})$ be $\mu$-ordinary, with $q=p^{n}$, and let ${}_{{}_{\mathrm{cris}}}\varphi_{x}$ be the crystalline (semilinear) Frobenius endomorphism on $\mathbb{D}_{x}$. Then, the linear map ${}_{{}_{\mathrm{cris}}}\varphi_{x}^{n}$ is semisimple.
9.1 Non-emptiness of the $\mu$-ordinary locus
We will now prove that the $\mu$-ordinary locus is non-empty at all large enough primes $v$ of the reflex field $E$, by constructing zero-dimensional Shimura varieties that specialize mod $v$ to $\mu$-ordinary points. It suffices to treat the case of adjoint groups, so we assume throughout that $G$ is adjoint. Let $(T,h)\rightarrow(G,X)$ denote an embedding of Shimura data such that the image of $T$ is a maximal torus of $G$. Let $E_{T}$ denote the reflex field of $(T,h)$ – note that this is just the field of definition of the Hodge co-character $\mu_{h}$ associated to $h$. Let $w$ be a place of $E_{T}$ dividing the place $v$ of $E$. Let $p$ denote the rational prime that $v$ divides. We make the following definition.
Definition 9.9.
Let $T^{\prime}\subset G^{\prime}$ denote a maximal torus inside a reductive group defined over $\mathbb{Z}_{p}$. We say that $T^{\prime}$ is quasi-split if $T^{\prime}$ is contained in a Borel subgroup $B\subset G$ defined over $\mathbb{Z}_{p}$.
We remark that quasi-split tori always exist, they are all conjugate to each other over $\mathbb{Z}_{p}$, and a torus inside a $\mathbb{Z}_{p}$-reductive group is quasi-split if and only if it contains a maximal split torus.
Proposition 9.10.
The zero-dimensional Shimura variety modulo $w$ induces a $\mu$-ordinary of $\mathscr{S}_{K}(G,X)_{\mathbb{F}_{v}}$ if the following two conditions are satisfied:
  1. 1.
    $T$ is quasi-split at $p$.
  2. 2.
    The co-character $\mu_{h}\in X_{*}(T_{E_{T,w}})$ is dominant with respect to a Borel subgroup of $G$ defined over $\mathbb{Z}_{p}$.
Proof.
As remarked earlier, the mod $w$ zero-dimensional Shimura variety is $\mu$-ordinary for the group $T$. Therefore, we have that the Newton co-character is the Galois average of $\mu_{h}$. However, in order to show that we also have $\mu$-ordinariness for the group $G$, we must show the same equality but after replacing both the Newton and the Hodge co-character with their dominant conjugates (dominant with respect to some Borel subgroup defined over $\mathbb{Q}_{p}$). However, $\mu_{h}$ is dominant by definition, and as $B$ is defined over $\mathbb{Q}_{p}$, so is its Galois average. The result follows. ∎
The absolute Galois group of $\mathbb{Q}$ acts on $X_{*}(T)$. We recall Chai and Oort’s definition from [ChaiOortJacobians].
Definition 9.11.
We say that $(T,h)$ is a Weyl special point if the image of the Galois action $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ on the co-character lattice $X_{*}(T)$ contains the Weyl group.
Given this setup, let $E(G)$ denote the finite extension of $\mathbb{Q}$ such that the image of the Galois action $\operatorname{Gal}(\overline{E(G)}/E(G))$ equals the Weyl group. Let $E^{\prime}_{T}$ denote the splitting field of $T$. Note that $E(G)$ contains $E$ and $E^{\prime}_{T}$ contains $E_{T}$.
Proposition 9.12.
Let $(T,h)$ be as above. Suppose that $T$ is quasi-split at $p$, and that $(T,h)$ is a Weyl special point. Then there exists a place $w$ of $E_{T}$ that satisfies the second condition of Proposition 9.10.
Proof.
It suffices to replace $E$, and $E_{T}$ with $E(G)$, and $E^{\prime}_{T}$ respectively. Abusing notation by letting $v$ denote a place of $E(G)$ dividing the original place of $E$, we must find a place $w\mid v$ of $E^{\prime}_{T}$ such that $\mu_{h}\in X_{*}(T_{E^{\prime}_{T,w}})$ satisfies the positivity condition with respect to a $\mathbb{Z}_{p}$-Borel. Given a co-character $\mu$ defined over $E^{\prime}_{T}$ and a place $w\mid v$, we let $\mu^{w}$ denote the base-change of $\mu$ to $E_{T,w}$. We also fix a choice of Borel subgroup over $\mathbb{Z}_{p}$ containing $T$.
We now pick a place $w^{\prime}\mid v$. By construction, the Weyl-group orbit of $\mu_{h}$ is defined over $E^{\prime}_{T}$ and therefore over $E^{\prime}_{T,w^{\prime}}$. As $T$ is split over $E^{\prime}_{T}$, we have that $\mu^{w^{\prime}}_{h}$ and any other element in its Weyl orbit are conjugate over $E^{\prime}_{T,w^{\prime}}$. Let $g$ be the Weyl group element satisfying $g\cdot\mu^{w^{\prime}}_{h}$ is dominant with respect to $B$. Let $\tau\in\operatorname{Gal}(E^{\prime}_{T}/E(G))$ satisfy $\tau(\mu_{h})=g\cdot\mu_{h}$. Then by construction, $\tau(\mu_{h})^{w^{\prime}}\in X_{*}(T_{E^{\prime}_{T,w^{\prime}}})$ is dominant. But this is equivalent to $\mu_{h}^{\tau^{-1}(w^{\prime})}\in X_{*}(T_{E^{\prime}_{T,\tau^{-1}(w^{\prime})}})$ being dominant as $B$ is defined over $\mathbb{Q}_{p}$. Therefore, the proposition follows by picking $w=\tau^{-1}(w^{\prime})$.
We are now ready to prove Theorem 9.2.
Proof.
By Propositions 9.10 and 9.12, it suffices to find a Weyl special point $(T,h)$ where $T$ is quasi-split at $p$. But this follows directly from [ChaiOortJacobians, Proposition 5.11] by choosing $\Sigma_{1}=\{p,\infty\}$, $U_{\infty}$ as in Remark 5.12 of loc. cit., and $U_{p}$ to be the set of set of elements in $\operatorname{Lie}G(\mathbb{Z}_{p})$ which reduce to an element of $\operatorname{Lie}G(\mathbb{Z}_{p})$ whose stabilizer in $G_{\mathbb{F}_{p}}$ is equal to a maximal torus containing a maximal split torus. ∎
9.2 An analogue of Tate’s isogeny theorem for ordinary points
We will now prove prove an analogue of Tate’s isogeny theorem for abelian varieties in the setting of ordinary points.
Theorem 9.13.
Let $x,y\in\mathscr{S}_{K}(G,X)(\mathbb{F}_{q})$ be ordinary points such that the $\ell$-adic Frobenii $\varphi_{x}$ and $\varphi_{y}$ are conjugate in $G(\mathbb{Q}_{\ell})$. Then, the rational Hodge structures associated to the canonical lifts $x^{{\mathrm{can}}}$ and $y^{{\mathrm{can}}}$ are isomorphic.
In the setting of (ordinary) abelian varieties, this result would translate to abelian varieties over finite fields being isogenous if their $\ell$-adic Frobenii are conjugate.
Proof.
Let $(T_{x},\mu_{x})$ and $(T_{y},\mu_{y})$ denote the zero-dimensional Shimura data inducing the special points $x^{{\mathrm{can}}}$ and $y^{{\mathrm{can}}}$. Let $E_{x}$ denote the reflex field of $(T_{x},\mu_{x})$ and let $v$ be the place of $E_{x}$ such that $x=x^{{\mathrm{can}}}\bmod v$. Recall that we have fixed $V$, a $\mathbb{Q}$-representation of $G$, which we will now consider as a representation of $T_{x}$. By [Milnefinitefieldmotives], there exists a surjective morphism of Shimura data $(T,\mu)\rightarrow(T_{x},\mu_{x})$ where $(T,\mu)$ is a zero-dimensional Shimura datum of Hodge type. Let $E$ denote the reflex field of $(T,\mu)$ ($E$ contains $E_{x}$), and let $v^{\prime}\mid v$ be some place of $E$ above $E_{x}$. Let $W$ be the tautological symplectic representation of $T$ associated to a Hodge embedding. Let $z\in S(T,\mu)$ be a point that maps to $\tilde{x}$ and let $A$ denote the CM abelian variety associated to $z$. Note that $T$ is the Mumford-Tate group of $A$. Let $\beta\in\operatorname{\mathsf{End}}(A)$ denote the endomorphism that specializes to the Frobenius endomorphism of $A\bmod v^{\prime}$. We will abuse notation and let $\beta\in T(\mathbb{Q})$ also denote the Betti-realization of the endomorphism $\beta$. Let ${}_{{\mathrm{et}}}\beta_{p}$ and ${}_{{}_{\mathrm{dR}}}\beta$ denote the endomorphisms of ${}_{{\mathrm{et}}}V_{p}$ and ${}_{{}_{\mathrm{dR}}}V$ induced by $\beta$. Restricting the Galois representation to the local Galois group at $v$, we may apply the Crystalline functor $\mathbb{D}_{{}_{\mathrm{cris}}}$ to ${}_{{\mathrm{et}}}\beta$ and obtain an endomorphism of ${}_{{}_{\mathrm{cris}}}V,{\mathrm{Fil}}^{\bullet}$ – this endomorphism is just the power of the crystalline Frobenius operator on ${}_{{}_{\mathrm{cris}}}V,{\mathrm{Fil}}^{\bullet}$, and equals ${}_{{}_{\mathrm{dR}}}\beta$ under the crystalline-deRham comparison isomorphism. All of this follows from the fact that $\beta\in T(\mathbb{Q})$ is induced by an endomorphism of the abelian variety $A$.
We set $q=p^{n}$. Let $\alpha\in T_{x}(\mathbb{Q})$ denote the image of $\beta$. As $V$ is in the Tannakian category generated by $W$ (as $T$-representations), we have that the same compatibility holds for the various realizations of $\alpha$. That is, we have that ${}_{{\mathrm{et}}}\alpha_{\ell}$ acts as Frobenius ${}_{{\mathrm{et}}}\varphi_{x,\ell}$ on ${}_{{\mathrm{et}}}V_{x,\ell}$ for $\ell\neq p$, and the Crystalline functor $\mathbb{D}_{{}_{\mathrm{cris}}}$ applied to ${}_{{\mathrm{et}}}\alpha_{p}$ the $n$th power of the Crystalline Frobenius ${}_{{}_{\mathrm{cris}}}\varphi_{x}$ on ${}_{{}_{\mathrm{cris}}}V_{x}$, and therefore that $(_{{}_{\mathrm{cris}}}\varphi_{x})^{n}=_{{}_{\mathrm{dR}}}\alpha$ under the Crystalline-deRham comparison isomorphism. We will now prove that the group generated by $\alpha$ is Zariski-dense in $T_{x}$. Let $L$ denote any field containing $E_{x}$. Let ${}_{{}_{\mathrm{dR}}}T_{\alpha}\subset\operatorname{GL}(_{{}_{\mathrm{dR}}}V_{\tilde{x},L})$ denote the torus [^8] By replacing $q$ by a sufficiently large power, this group will always be connected and therefore a torus. defined by taking the Zariski-closure of the group generated by ${}_{{}_{\mathrm{dR}}}\alpha$. It suffices to prove that the Hodge co-character factors through ${}_{{}_{\mathrm{dR}}}T_{\alpha}$ (note that this doesn’t depend on the field $L$). We pick $L=W(\mathbb{F}_{q})[1/p]$. By the Crystalline-de Rham comparison (and the compatibility outlined above), we may replace ${}_{{}_{\mathrm{dR}}}\alpha$ and ${}_{{}_{\mathrm{dR}}}V_{x^{{\mathrm{can}}},L},{\mathrm{Fil}}^{\bullet}$ by ${}_{{}_{\mathrm{cris}}}\varphi^{n}$ and ${}_{{}_{\mathrm{cris}}}V_{\tilde{x}},{\mathrm{Fil}}^{\bullet}$. As the point $x$ is ordinary, we have a direct sum decomposition of the crystal ${}_{{}_{\mathrm{cris}}}V_{x},_{{}_{\mathrm{cris}}}\varphi_{x}=\bigoplus_{{}_{\mathrm{cris}}}V_{x}^{i},\oplus_{{}_{\mathrm{cris}}}\varphi^{i}$, where $\frac{{}_{{}_{\mathrm{cris}}}\varphi^{i}}{p^{i}}$ is a bijective $\sigma$-linear Frobenius morphism. By definition, the Hodge filtration on ${}_{{}_{\mathrm{cris}}}V_{x}$ defined by the canonical lifting is ${\mathrm{Fil}}^{j}=\bigoplus_{i\leq j}\ {}_{{}_{\mathrm{cris}}}V_{x}^{i}$. We have that ${}_{{}_{\mathrm{dR}}}\alpha$ is just the $n$th iterate of ${}_{{}_{\mathrm{cris}}}\varphi_{x}$, and the eigenvalues of ${}_{{}_{\mathrm{dR}}}\alpha$ on ${}_{{}_{\mathrm{cris}}}V_{x}^{i}$ have $p$-adic valuation exactly equal to $ni$. In order to prove that the Hodge co-character factors through $T_{\alpha}$ it suffices to prove the following general claim:
Claim.
Let $K$ be a finite extension of $\mathbb{Q}_{p}$, let $T^{\prime}$ be a and let $V^{\prime}$ be a faithful algebraic representation of $T^{\prime}$ defined over $K$. Let $\mu:\operatorname{\mathbb{G}}_{m}\rightarrow T^{\prime}$ be a co-character that induces weight decomposition $V^{\prime}=\bigoplus V^{\prime,i}$. Let $\alpha^{\prime}\in T^{\prime}(K)$ be an element that acts on $V^{\prime,i}$ with eigenvalues having valuation $ni$. Then, $\mu$ factors through $T^{\prime}_{\alpha^{\prime}}$ the Zariski closure of the group generated by $\alpha^{\prime}$.
Proof.
By replacing $K$ by a finite extension if need be, we may assume that $T^{\prime}$ is a split torus. Without loss of generality, we may also assume that the rank of $T^{\prime}$ equals the dimension of $V^{\prime}$. Let $\chi$ denote any character of $T^{\prime}$ such that $\chi(\alpha^{\prime})=1$. It suffices to prove that $\chi\circ\mu$ is the trivial character on $\operatorname{\mathbb{G}}_{m}$. We may write $T^{\prime}=\prod\prod T^{\prime,i}$ according to the decomposition $V^{\prime}=\bigoplus V^{\prime,i}$. Picking coordinates on $V^{\prime,i}$, we may write $T^{\prime,i}=\prod\operatorname{\mathbb{G}}_{m}$, with the diagonal $\operatorname{\mathbb{G}}_{m}$ acting by scalars on $V^{\prime,i}$. In these coordinates, $\alpha^{\prime}$ will decompose as $(\alpha^{\prime,i})$ with the entries of $\alpha^{\prime,i}$ having $p$-adic valuation $ni$. Let $\chi$ be a character on $T^{\prime}$. Then we may write $\chi$ as $\prod_{i}\chi_{i}$, with $\chi_{i}=(a_{i,1},\ldots a_{i,N_{i}})$, where $\chi_{i}((x_{j}))=\prod x_{j}^{a_{i,j}}$ for $(x_{j})\in T^{\prime,i}=\prod\operatorname{\mathbb{G}}_{m}$. A necessary condition for $\chi(\alpha^{\prime})=1$ is that the $p$-adic valuation of $\chi(\alpha^{\prime})$ is 0, i.e. $\sum_{i}ni(a_{i,1}+\ldots a_{i,N_{i}})=0$. On the other hand, the co-character $\mu$ is given by $\mu=\prod\mu_{i}$, with $\mu_{i}(z)=(z^{i},z^{i},\ldots z^{i})$. Therefore, $\chi\circ\mu$ is trivial if and only if $\sum_{i}i(a_{i,1}+\ldots a_{i,N_{i}})=0$. Therefore, we’ve shown that if $\chi(\alpha)=1$, then $\chi\circ\mu$ is trivial, and therefore $\mu$ does factor through $T_{\alpha^{\prime}}$ as required.
Therefore, we have that $T_{\alpha}=T_{x}$ as required. We have the analogous statement for the Frobenius element associated to $y$ and the Torus $T_{y}$. By assumption, the Frobenii satisfy the same characteristic polynomial on $V$. Therefore, there exists a $\mathbb{Q}$-endomorphism $X\in\operatorname{GL}(V)$ with $X^{-1}\alpha X=\gamma$. As the groups generated by $\alpha$ and $\gamma$ are respectively Zariski dense in $T_{x}$ and $T_{y}$ respectively, we have that $X^{-1}T_{x}X=T_{y}$. Finally, the Hodge co-characters are conjugate to each other because the CM-types are uniquely determined by the elements $\alpha$ and $\gamma$. The proposition follows. ∎